### v1

# A topological constraint on clean Lagrangian intersections from $\mathbb{Q}$-valued augmentations

Abstract
For a knot $K$ in $\mathbb{R}^{3}$, consider its conormal bundle $L_{K}$ in $T^{*}\mathbb{R}^{3}$. We prove that if $K$ is the connected sum $K^{\prime}\#3_{1}$ of an arbitrary knot $K^{\prime}$ and the trefoil knot $3_{1}$ (or its mirror), then there is no compactly supported Hamiltonian diffeomorphism $\varphi$ on $T^{*}\mathbb{R}^{3}$ such that $\varphi(L_{K})$ and the zero section $\mathbb{R}^{3}$ intersect cleanly along the unknot in $\mathbb{R}^{3}$. For the proof, we utilize the knot DGA, whose homology was proved to be isomorphic to the knot contact homology by Ekholm, Etnyre, Ng and Sullivan, with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$. We discover a certain algebraic constraint imposed on any augmentation of the knot DGA of $K=K^{\prime}\#3_{1}$ to a field $\mathbb{k}$. This constraint is non-trivial for $\mathbb{k}=\mathbb{Q}$ and plays a key role in the proof, but trivial if $\mathbb{k}$ is quadratically closed or if an augmentation factors through the knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$ obtained by setting $U=1$.
1 Introduction
Background. Let us consider the cotangent bundle $T^{*}\mathbb{R}^{3}$ of $\mathbb{R}^{3}$ with a canonical symplectic structure. We identify the image of the zero section with $\mathbb{R}^{3}$. Given any knot $K$ in $\mathbb{R}^{3}$, its conormal bundle is defined by
$L_{K}\coloneqq\{(q,p)\in T^{*}\mathbb{R}^{3}\mid q\in K,\ \langle p,v\rangle=0\text{ for every }v\in T_{q}K\}.$
It is a Lagrangian submanifold of $T^{*}\mathbb{R}^{3}$ and intersects $\mathbb{R}^{3}$ cleanly along the knot $K$. Let $\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$ denote the set of Hamiltonian diffeomorphisms on $T^{*}\mathbb{R}^{3}$ associated to compactly supported Hamiltonians on $T^{*}\mathbb{R}^{3}$. Using these notations, we ask a question which was observed by the author in [17].
Question 1.1.
Is there $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K})$ and $\mathbb{R}^{3}$ have a clean intersection along a knot in $\mathbb{R}^{3}$ whose isotopy class is different from $K$?
We remark that it is essential that $\varphi$ is a Hamiltonian diffeomorphism. Indeed, for any pair of knots $(K_{0},K_{1})$ in $\mathbb{R}^{3}$, there exists a $C^{\infty}$ isotopy $(\varphi^{t})_{t\in[0,1]}$ on $T^{*}\mathbb{R}^{3}$ with compact support such that $\varphi^{1}(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along $K_{1}$. See [17, Proposition 4.1].
As a special case, the answer to Question 1.1 when $K$ is the unknot is ‘no’. This can be deduced from the results by Ganatra-Pomerleano [10, Proposition 6.29], Smith-Wemyss [21, Proposition 4.10] or Asplund-Li [1, Theorem 1.3] about clean Lagrangian intersections in the plumbings of two copies of the cotangent bundle of $S^{3}$ along the unknots. These results are concerned with Smith’s question on the ‘persistence (or rigidity) of unknottedness of Lagrangian intersections’ [10, Question 1.5]. For its relation to Question 1.1, see [17, Section 1].
In a general case, under the assumption that $\varphi(L_{K})$ for $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ and $\mathbb{R}^{3}$ intersect cleanly along a knot in $\mathbb{R}^{3}$, the author proved [17, Theorem 1.1] by using the (framed) knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$, which is a differential graded algebra (DGA) defined by Ng [14]. [17, Theorem 1.1] is stated by using the augmentation variety in $(\mathbb{C}^{*})^{2}$ derived from the set of augmentations of the knot DGA to $\mathbb{C}$. This theorem leads us to the same answer to Question 1.1 as above when $K$ is the unknot [17, Theorem 1.3] and we can also deduce some constraints on the knot type of $\varphi(L_{K})\cap\mathbb{R}^{3}$ when $K$ has certain non-trivial knot types [17, Theorem 1.4, 1.5]. However, we cannot deduce any constraint from [17, Theorem 1.1] on the knot type of $K$ if we suppose that $\varphi(L_{K})\cap\mathbb{R}^{3}$ is the unknot in $\mathbb{R}^{3}$. For the reason, see Remark 1.5 below.
Main results. Let $3_{1}$ denote the right-handed trefoil knot in $\mathbb{R}^{3}$ and $\overline{3_{1}}$ denote its mirror. The goal of this paper is to prove the next theorem which cannot be deduced from [17, Theorem 1.1].
Theorem 1.1 (Theorem 3.10).
Let $K$ be the connected sum of an arbitrary knot in $\mathbb{R}^{3}$ with $3_{1}$ or $\overline{3_{1}}$. Then, there is no $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K})$ intersects $\mathbb{R}^{3}$ cleanly along the unknot in $\mathbb{R}^{3}$.
For the proof, the geometric arguments in the present paper are similar to those in [17], which are outlined as follows: Let $K_{0}$ and $K_{1}$ be knots in $\mathbb{R}^{3}$. For $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$, suppose that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ intersect cleanly along $K_{1}$. We use the results by Ekholm-Etnyre-Ng-Sullivan [7, 4] which show that the Chekanov-Eliashberg DGA of the unit conormal bundle of $K_{i}$ is stable tame isomorphic to the knot DGA of $K_{i}$ for $i=0,1$. The knot DGA, introduced by Ng [12, 13], is defined combinatorially from a braid representation of a knot. We then apply a DGA map associated to an exact Lagrangian cobordism between the unit conormal bundles of $K_{0}$ and $K_{1}$. This is a part of the symplectic field theory introduced by Eliashberg-Givental-Hofer [9]. Consequently, we obtain a DGA map from the knot DGA of $K_{0}$ to that of $K_{1}$. (Precisely, we need a base change on the coefficient ring of a DGA.)
The feature of the present paper is that we utilize the knot DGA, which we denote here by $(\mathcal{A}^{\mathrm{knot}}_{*}(K),\partial_{K})$ for an oriented knot $K$, with coefficients in the ring
$\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]\cong\mathbb{Z}[H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K})].$
Here, $S^{*}\mathbb{R}^{3}$ denotes the unit cotangent bundle of $\mathbb{R}^{3}$ and $\Lambda_{K}$ denotes the unit conormal bundle of $K$. For details, we refer to [15, 4, 16]. We note that in [17], referring to [14, 7], the knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]\cong\mathbb{Z}[H_{1}(\Lambda_{K})]$ was used.
To prove Theorem 1.1, we show the following algebraic property of the knot DGA under an assumption on the isotopy class of $K$. Let $K^{\prime}\#3_{1}$ denote the connected sum of a knot $K^{\prime}$ in $\mathbb{R}^{3}$ and the trefoil knot $3_{1}$.
Lemma 1.2 (Lemma 3.9).
Let $\mathbb{k}$ be a field and $K^{\prime}$ be a knot. Then, for any augmentation $\varepsilon\colon\mathcal{A}_{*}^{\mathrm{knot}}(K^{\prime}\#3_{1})\to\mathbb{k}$ of the knot DGA of $K^{\prime}\#3_{1}$ to $\mathbb{k}$,
$D_{\varepsilon}\coloneqq(\varepsilon(\mu)+1-\varepsilon(\mu)^{-1}\varepsilon(U))^{2}-4(\varepsilon(\mu)-1)$
satisfies $D_{\varepsilon}\in\{k^{2}\mid k\in\mathbb{k}\}$.
Here, $\varepsilon(\mu)$ and $\varepsilon(U)$ are abbreviations of $\varepsilon(\mu\cdot\mathbb{1})$ and $\varepsilon(U\cdot\mathbb{1})$ for the unit $\mathbb{1}$ of the knot DGA. In fact, $D_{\varepsilon}$ is the discriminant of a quadratic equation in $\mathbb{k}$ determined by $\varepsilon$ (see the proof of Lemma 3.9). Note that if $\varepsilon$ factors through the knot DGA over $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$ obtained by setting $U=1$, then $\varepsilon(U)=1$ and we have an identity
$(\varepsilon(\mu)+1-\varepsilon(\mu)^{-1})^{2}-4(\varepsilon(\mu)-1)=(\varepsilon(\mu)-1+\varepsilon(\mu)^{-1})^{2},$
so the constraint on $D_{\varepsilon}$ is trivial.
For the proof of Theorem 1.1, we use the constraint from Lemma 1.2 in the case of $\mathbb{k}=\mathbb{Q}$, that is, $D_{\varepsilon}$ is the square of some rational number for any augmentation $\varepsilon$ of the knot DGA of $K^{\prime}\#3_{1}$ to $\mathbb{Q}$. (In fact, for this purpose, we may choose other fields such as $\mathbb{Q}[\sqrt{-1}]$ and $\mathbb{F}_{5}$. See Remark 3.11.) It is crucial that $\mathbb{Q}$ is not quadratically (and algebraically) closed, and we remark that the constraint on $D_{\varepsilon}$ is trivial for any $\varepsilon$ if $\mathbb{k}=\mathbb{C}$.
Remark 1.3.
In order to study the knot DGA of a knot $K$ (or the Chekanov-Eliashberg DGA of $\Lambda_{K}$), the set of augmentations to $\mathbb{k}$ has been used with $\mathbb{k}=\mathbb{C}$ or a finite field $\mathbb{F}_{p}$ for a prime number $p$. For examples of applications, see Ng [14, Section 5.2 and 5.3] and Cornwell [2] about the augmentation polynomial when $\mathbb{k}=\mathbb{C}$ and [14, Section 5.1] about the augmentation number when $\mathbb{k}=\mathbb{F}_{p}$. On the other hand, the rational numbers $D_{\varepsilon}$ from Lemma 1.2 when $\mathbb{k}=\mathbb{Q}$ are derived in a way different from those previous researches.
Lastly, we provide a theorem which is a refinement of [17, Theorem 1.1]. For an oriented knot $K$ in $\mathbb{R}^{3}$, let $V_{K}$ denote the augmentation variety defined by [16, Definition 5.1] as a subset of $(\mathbb{C}^{*})^{3}$. Let $(x,y,Z)$ be the coordinate of $\mathbb{C}^{3}$.
Theorem 1.4 (Corollary 3.5).
For oriented knots $K_{0}$ and $K_{1}$ in $\mathbb{R}^{3}$, suppose that there exists $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ intersects $\mathbb{R}^{3}$ cleanly along $K_{1}$. Then, there exist integers $b,c$ and $d$ for which we have an injective map
$V_{K_{1}}\hookrightarrow V_{K_{0}}\colon(x,y,Z)\mapsto(xy^{b}Z^{c},yZ^{d},Z).$
Remark 1.5.
Identifying $\{(x,y,Z)\in(\mathbb{C}^{*})^{3}\mid Z=1\}$ with $(\mathbb{C}^{*})^{2}$, the augmentation variety considered in [17] coincides with $V_{K}\cap\{Z=1\}\subset(\mathbb{C}^{*})^{2}$. It follows from Theorem 1.4 that there exists an injective map
$V_{K_{1}}\cap\{Z=1\}\hookrightarrow V_{K_{0}}\cap\{Z=1\}\colon(x,y)\mapsto(xy^{b},y)$
for some $b\in\mathbb{Z}$. This can be seen as a slight refinement of [17, Theorem 1.1], but it gives no constraint on $K_{0}$ if we suppose that $K_{1}$ is the unknot. The reason is that $V_{K_{1}}\cap\{Z=1\}=\{x=1\}\cup\{y=1\}$ if $K_{1}$ is the unknot (see [16, Example 3.13]) and $\{x=1\}\cup\{y=1\}$ is a subset of $V_{K_{0}}\cap\{Z=1\}$ for any knot $K_{0}$ (see [14, Proposition 5.6] and [16, Exercise 3.17]).
Notations. In this paper, for any manifold $X$, we take $\mathbb{Z}$ as the coefficient ring of the singular homology group of $X$ and denote it by $H_{*}(X)$, unless otherwise specified. Likewise, the relative homology group over $\mathbb{Z}$ is denoted by $H_{*}(X,A)$ for any submanifold $A$ of $X$.
Organization of paper. In Section 2, for a contact manifold $Y$ and its Legendrian submanifold $\Lambda$ satisfying certain conditions, we introduce the Chekanov-Eliashberg DGA of $(Y,\Lambda)$ with coefficients in $\mathbb{Z}[H_{2}(Y,\Lambda)]$. We then define a DGA map associated to an exact Lagrangian cobordism in ×$\mathbb{R}\times Y$. In Section 2.5, we will see how to apply this DGA map to a setting with a clean Lagrangian intersection as in Question 1.1. In Section 3, we combine the argument in Section 2.5 with the results of [4, 16] about the knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$. Theorem 1.4 is proved in Section 3.1. Theorem 1.1 is proved by using Lemma 1.2 (Lemma 3.9) in Section 3.2. In Appendix A, by using Floer cohomology group over $\mathbb{F}_{2}[\lambda^{\pm}]$, we prove Proposition 2.7 which gives a homological constraint on $1$-dimensional clean intersections. This enhances some results in [17]. Nevertheless, the present proof is much simpler than that in [17].
Acknowledgements. The author would like to thank Kei Irie for making valuable comments. He would also like to thank Fumikazu Nagasato and Kouki Sato for fruitful discussion before this work and for giving some insight from knot theory. This work was supported by JSPS KAKENHI Grant Number JP23KJ1238.
2 Symplectic field theory for exact Lagrangian cobordisms
We fix a $2n$-dimensional simply connected manifold $P$ with a Liouville $1$-form $\lambda$ such that $P$ is the completion of a compact Liouville domain $(\bar{P},\left.\lambda\right|_{\bar{P}})$ satisfying $c_{1}(T\bar{P})=0$. Let us consider
$\begin{array}[]{ll}Y\coloneqq P\times\mathbb{R},&\alpha\coloneqq dz-\lambda,\end{array}$
where $z$ is the coordinate of $\mathbb{R}$. Then, $\alpha$ is a contact form on $Y$ and the Reeb vector field on $Y$ with respect to $\alpha$ agrees with $\partial_{z}$. Let $\pi_{P}\colon Y\to P$ denote the projection to $P$.
The arguments from Section 2.1 to Section 2.4 are completely parallel to those in [17, Section 3], except the difference of the coefficient rings. In Section 2.5, we will discuss a particular case related to clean Lagrangian intersections.
2.1 Preliminaries
We fix a complex structure $j_{0}$ on ×$\mathbb{R}\times[0,1]$ such that $(j_{0})_{(s,t)}(\partial_{s})=\partial_{t}$, where $(s,t)$ is the coordinate of ×$\mathbb{R}\times[0,1]$. Let $D\coloneqq\{z\in\mathbb{C}\mid|z|\leq 1\}$. For any $m\in\mathbb{Z}_{\geq 0}$, we fix $p_{0},p_{1},\dots,p_{m}\in\partial D$ ordered counterclockwise and define $D_{m+1}\coloneqq D\setminus\{p_{0},p_{1},\dots,p_{m}\}$. Let $\mathcal{C}_{m+1}$ denote the space of conformal structures on $D_{m+1}$ which can be extended smoothly on $D$. Given $\kappa\in\mathcal{C}_{m+1}$, we let $j_{\kappa}$ denote a complex structure on $D_{m+1}$ which associates $\kappa$. Near each boundary puncture of the Riemann surface $(D_{m+1},j_{\kappa})$, we fix a strip coordinate by a biholomorphic map
$\psi_{0}$ $\colon[0,\infty)\times[0,1]\to D_{m+1},$
$\psi_{k}$ $\colon(-\infty,0]\times[0,1]\to D_{m+1}\text{ for }k=1,\dots,m,$
with respect to $j_{0}$ and $j_{\kappa}$ such that $\lim_{s\to\infty}\psi_{0}(s,t)=p_{0}$ and $\lim_{s\to-\infty}\psi_{k}(s,t)=p_{k}$ uniformly on $t\in[0,1]$.
Let $\overline{D_{m+1}}$ be a compactification of $D_{m+1}$ obtained by gluing $[0,\infty]\times[0,1]$ via $\psi_{0}$ and $[-\infty,0]\times[0,1]$ via $\psi_{k}$ for every $k=1,\dots,m$. In a natural way, $\overline{D_{m+1}}$ has a structure of a compact oriented surface with boundary and corner.
Let $\Lambda$ be a compact Legendrian submanifold of $(Y,\alpha)$. Later, we will impose the following conditions on $\Lambda$:
  1. 1.
    $\Lambda$ is a connected and oriented manifold and admits a spin structure.
  2. 2.
    For the immersion $\left.\pi_{P}\right|_{\Lambda}\colon\Lambda\to P$, the self-intersection of $\pi_{P}(\Lambda)$ consists only of transversal double points.
  3. 3.
    For the immersed Lagrangian submanifold $\pi_{P}(\Lambda)$ of $P$, its Maslov class vanishes.
The second condition is satisfied for generic $\Lambda$. We denote the set of Reeb chords of $\Lambda$ by
${\mathcal{R}(\Lambda)\coloneqq\{a\colon[0,T]\to Y\mid T>0,\ \frac{da}{dt}(t)=(\partial_{z})_{a(t)}\text{ for every }t\in[0,T]\text{ and }a(0),a(T)\in\Lambda\}.}$
Hereafter, for any Reeb chord $(a\in[0,T]\to Y)\in\mathcal{R}(\Lambda)$, we write $T=\int a^{*}\alpha$ as $T_{a}$. For each $a\in\mathcal{R}(\Lambda)$, we fix a continuous map $w_{a}\colon\overline{D_{1}}\to Y$ such that $w_{a}(\partial D_{1})\subset\Lambda$ and
$\lim_{s\to\infty}w_{a}\circ\psi_{0}(s,t)=a(T_{a}t)$
for every $t\in[0,1]$. When the second condition above is satisfied, the Conley-Zehnder index $\mu(a)\in\mathbb{Z}$ of $a$ is defined by using the capping path $\left.w_{a}\right|_{\partial D_{1}}$ for $a$ as described in [17, Section 3.1]. We then define
$|a|\coloneqq\mu(a)-1.$
2.2 Moduli space of pseudo-holomorphic curves with boundary in an exact Lagrangian cobordism
Let $\Lambda_{+}$ and $\Lambda_{-}$ be two compact Legendrian submanifolds of $(Y,\alpha)$. Recall that a submanifold $L$ of ×$\mathbb{R}\times Y$ is called an exact Lagrangian cobordism from $\Lambda_{-}$ to $\Lambda_{+}$ if the following are satisfied:
  • There exist $r_{+},r_{-}\in\mathbb{R}$ with $r_{-}<r_{+}$ such that
    $\begin{array}[]{ll}L\cap([r_{+},\infty)\times Y)=[r_{+},\infty)\times\Lambda_{+},&L\cap((-\infty,r_{-}]\times Y)=(-\infty,r_{-}]\times\Lambda_{-},\end{array}$
    and $L\cap([r_{-},r_{+}]\times Y)$ is compact.
  • There exists a $C^{\infty}$ function $f\colon L\to\mathbb{R}$ such that $df=\left.(e^{r}\alpha)\right|_{L}$ and
    $\left.f\right|_{(-\infty,r_{-}]\times\Lambda_{-}}\colon(-\infty,r_{-}]\times\Lambda_{-}\to\mathbb{R}$
    is a constant function. Here, $r$ is the $\mathbb{R}$-coordinate of ×$\mathbb{R}\times Y$.
In the rest of this paper, we only consider the case where $\Lambda_{-}$ is connected. Therefore, we do not need to care about the last condition that $\left.f\right|_{(-\infty,r_{-}]\times\Lambda_{-}}$ is a constant function since
$\left.df\right|_{(-\infty,r_{-}]\times\Lambda_{-}}=\left.(e^{r}\alpha)\right|_{(-\infty,r_{-}]\times\Lambda_{-}}=0,$
which implies that $\left.f\right|_{(-\infty,r_{-}]\times\Lambda_{-}}$ is locally constant.
Associated to $L$, we use the following notations for maps between pairs of spaces:
$\begin{split}i_{+}&\colon(Y,\Lambda_{+})\to(\mathbb{R}\times Y,L)\colon x\mapsto(r_{+},x),\\ i_{-}&\colon(Y,\Lambda_{-})\to(\mathbb{R}\times Y,L)\colon x\mapsto(r_{-},x),\end{split}$ (1)
and
$R_{L}\colon(\mathbb{R}\times Y,L)\to([r_{-},r_{+}]\times Y,L\cap([r_{-},r_{+}]\times Y))$
is a strong deformation retract such that $R_{L}(r,x)=(r_{+},x)$ if $r\geq r_{+}$ and $R_{L}(r,x)=(r_{-},x)$ if $r\leq r_{-}$. In addition, for every $s\in\mathbb{R}$, we define a map
$\tau_{s}\colon\mathbb{R}\times Y\to\mathbb{R}\times Y\colon(r,x)\mapsto(r+s,x).$
Let $J$ be an almost complex structure on $P$ such that $d\lambda(\cdot,J\cdot)$ is a Riemannian metric on $P$. We associate to $J$ an almost complex structure $\widetilde{J}$ on ×$\mathbb{R}\times Y$ determined by $\tau_{s}^{*}\widetilde{J}=\widetilde{J}$ for every $s\in\mathbb{R}$, $\widetilde{J}(\operatorname{Ker}\alpha)=\operatorname{Ker}\alpha$$(\pi_{P})_{*}(\widetilde{J}(v))=J((\pi_{P})_{*}(v))$ for every $v\in\operatorname{Ker}\alpha$ and $\widetilde{J}(\partial_{z})=\partial_{r}$.
For any $a\in\mathcal{R}(\Lambda_{+})$, $b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda_{-})$ and $C\in H_{2}(\mathbb{R}\times Y,L)$, we define
^$\hat{\mathcal{M}}_{L,C}(a;b_{1},\dots,b_{m})$
to be the space of pairs $(u,\kappa)$ of $\kappa\in\mathcal{C}_{m+1}$ and a $C^{\infty}$ map $u\colon D_{m+1}\to P$ such that:
  • $du+\widetilde{J}\circ du\circ j_{\kappa}=0$.
  • $u(\partial D_{m+1})\subset L$.
  • There exist $s_{0},s_{1},\dots,s_{m}\in\mathbb{R}$ such that
    $\lim_{s\to\infty}\tau_{-s}\circ u\circ\psi_{0}$ $=(s_{0},a(T_{a}(1-t))),$
    $\lim_{s\to-\infty}\tau_{s}\circ u\circ\psi_{k}$ $=(s_{k},b_{k}(T_{b_{k}}(1-t)))\text{ for }k=1,\dots,m,$
    $C^{\infty}$-uniformly on $[0,1]$.
  • Let $\bar{u}\colon\overline{D_{m+1}}\to L$ denote the continuous extension of the map $R_{L}\circ u\colon D_{m+1}\to L$. Then, the 2-chain
    $\bar{u}+(i_{+}\circ w_{a})-((i_{-}\circ w_{b_{1}})+\dots+(i_{-}\circ w_{b_{m}}))$
    represents the homology class $C$. (More precisely, each term is regarded as a singular 2-chain after taking suitable triangulations of $\overline{D_{m+1}}$ and $\overline{D_{1}}$.)
When $m\leq 1$, $\mathcal{C}_{m+1}$ consists of a single element and the group $\mathcal{A}_{m+1}$ of conformal automorphisms on $D_{m+1}$ acts on ^$\hat{\mathcal{M}}_{L,C}(a;b_{1},\dots,b_{m})$ by $(u,\kappa)\cdot\varphi\coloneqq(u\circ\varphi,\kappa)$ for every $\varphi\in\mathcal{A}_{m+1}$. As in [17, (10) in Section 3.2], let us define a moduli space to be
$\mathcal{M}_{L,C}(a;b_{1},\dots,b_{m})\coloneqq\begin{cases}\hat{\mathcal{M}}_{L,C}(a;b_{1},\dots,b_{m})/\mathcal{A}_{m+1}&\text{ if }m=0,1,\\ \hat{\mathcal{M}}_{L,C}(a;b_{1},\dots,b_{m})&\text{ if }m\geq 2.\end{cases}$
In the case where $\Lambda_{+}=\Lambda_{-}=\Lambda$ and $L=\mathbb{R}\times\Lambda$, for any $a,b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda)$ and $A\in H_{2}(Y,\Lambda)$, an $\mathbb{R}$-action on ×$\mathcal{M}_{\mathbb{R}\times\Lambda,(i_{+})_{*}(A)}(a;b_{1},\dots,b_{m})$ is defined by $s\cdot(u,\kappa)\coloneqq(\tau_{s}\circ u,\kappa)$ for every $s\in\mathbb{R}$. Let us denote the quotient space by
¯$\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})\coloneqq\mathbb{R}\backslash\mathcal{M}_{\mathbb{R}\times\Lambda,(i_{+})_{*}(A)}(a;b_{1},\dots,b_{m}).$
2.3 Chekanov-Eliashberg DGA over $\mathbb{Z}[H_{2}(Y,\Lambda)]$
We consider a compact Legendrian submanifold $\Lambda$ of $(Y,\alpha)$ satisfying the three conditions in Section 2.1. We also assume that $\Lambda$ is admissible in the sense of the definition in [6, Section 2.3]. Let us take a spin structure $\mathfrak{s}$ on $\Lambda$.
We take an almost complex structure $J$ on $P$ as in [6, Proposition 2.3] such that the following hold for every $a,b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda)$ and $A\in H_{2}(Y,\Lambda)$: The moduli space ¯$\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ is cut out transversely and it is a manifold of dimension $|a|-(|b_{1}|+\dots+|b_{m}|)-1$. Moreover, when $|b_{1}|+\dots+|b_{m}|=|a|-1$, $\coprod_{A\in H_{2}(Y,\Lambda)}\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ is a compact 00-dimensional manifold. (For the relation between $J$-holomorphic disks in $P$ defined as in [6, Section 2.3] and $\widetilde{J}$-holomorphic maps in ×$\mathbb{R}\times Y$ defined as in Section 2.2, see [3, Theorem 2.1] and [17, Appendix A.4].)
In addition, ¯$\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ has an orientation depending on $\mathfrak{s}$ and a choice of an orientation of the determinant line of the capping operator for each Reeb chord of $\Lambda$. See [17, Definition A.4] for details which follow the conventions in [5] related to orientations.
Let $\mathcal{A}_{*}(\Lambda)$ be the graded $\mathbb{Z}[H_{2}(Y,\Lambda)]$-algebra freely generated by $\mathcal{R}(\Lambda)$ whose graded is given by $|\cdot|\colon\mathcal{R}(\Lambda)\to\mathbb{Z}$. A derivative over $\mathbb{Z}[H_{2}(Y,\Lambda)]$ of degree $-1$
$\partial_{J,\mathfrak{s}}\colon\mathcal{A}_{*}(\Lambda)\to\mathcal{A}_{*-1}(\Lambda)$
is defined for every $a\in\mathcal{R}(\Lambda)$ by
$\partial_{J,\mathfrak{s}}(a)\coloneqq\sum_{|b_{1}|+\dots+|b_{m}|=|a|-1}\sum_{A\in H_{2}(Y,\Lambda)}(-1)^{(n-1)(|a|+1)}\left(\#_{\operatorname{sign}}\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})\right)e^{A}\cdot b_{1}\cdots b_{m}$
and extended by the Leibniz rule. Here, #$\#_{\operatorname{sign}}$ is the number of the points counted with the signs induced by the orientation.
We refer to [5, Theorem 4.1, Theorem 4.12] which takes $\mathbb{Z}[H_{1}(\Lambda)]$ as the coefficient ring of the Chekanov-Eliashberg DGA. We modify the arguments in their proofs to the present case where we take $\mathbb{Z}[H_{2}(Y,\Lambda)]$ as the coefficient ring. Then, we have the following theorem.
Theorem 2.1.
$\partial_{J,\mathfrak{s}}$ is a differential, that is, $\partial_{J,\mathfrak{s}}\circ\partial_{J,\mathfrak{s}}=0$. Moreover, the stable tame isomorphism class over $\mathbb{Z}[H_{2}(Y,\Lambda)]$ of the DGA $(\mathcal{A}_{*}(\Lambda),\partial_{J,\mathfrak{s}})$ is independent of the choice of $J$ and invariant under Legendrian isotopies of $\Lambda$ preserving spin structures.
Let us discuss the effect by changing the spin structure on $\Lambda$. For any two spin structures $\mathfrak{s}_{0}$ and $\mathfrak{s}_{1}$ on $\Lambda$, a cohomology class $d(\mathfrak{s}_{0},\mathfrak{s}_{1})\in H^{1}(\Lambda;\mathbb{Z}/2)$ is defined as in [5, Subsection 4.4.1]. Let $\partial\colon H_{2}(Y,\Lambda)\to H_{1}(\Lambda)$ denote the boundary map. For any $\beta\in H_{1}(\Lambda)$, if we change the choice of a spin structure from $\mathfrak{s}$ to $\mathfrak{s}^{\prime}$, the orientation on the moduli space
$\coprod_{|b_{1}|+\dots+|b_{m}|=|a|-1}\coprod_{A\in H_{2}(Y,\Lambda)\colon\partial A=\beta}\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$
is changed by the sign $(-1)^{\langle d(\mathfrak{s},\mathfrak{s}^{\prime}),\beta\rangle}$. See the proof of [5, Theorem 4.29]. We define a ring isomorphism
$\psi(\mathfrak{s},\mathfrak{s}^{\prime})\colon\mathbb{Z}[H_{2}(Y,\Lambda)]\to\mathbb{Z}[H_{2}(Y,\Lambda)]\colon e^{A}\mapsto(-1)^{\langle d(\mathfrak{s},\mathfrak{s}),\partial A\rangle}e^{A}.$ (2)
Then, the next proposition easily follows as in the proof of [5, Theorem 4.30].
Proposition 2.2.
We define a graded ring isomorphism $I(\mathfrak{s},\mathfrak{s}^{\prime})\colon\mathcal{A}_{*}(\Lambda)\to\mathcal{A}_{*}(\Lambda)$ by
$I(\mathfrak{s},\mathfrak{s}^{\prime})(k\cdot a)=\psi(\mathfrak{s},\mathfrak{s}^{\prime})(k)\cdot a$
for every $k\in\mathbb{Z}[H_{2}(Y,\Lambda)]$ and $a\in\mathcal{R}(\Lambda)$. Then $\partial_{J,\mathfrak{s}^{\prime}}\circ I(\mathfrak{s},\mathfrak{s}^{\prime})=I(\mathfrak{s},\mathfrak{s}^{\prime})\circ\partial_{J,\mathfrak{s}}$.
2.4 DGA map associated to an exact Lagrangian cobordism
We consider an exact Lagrangian cobordism $L$ from $\Lambda_{-}$ to $\Lambda_{+}$ as in Section 2.2 such that the Maslov class of $L$ vanishes and both $\Lambda_{+}$ and $\Lambda_{-}$ are connected. We assume that $L$ is oriented and has a spin structure $\mathfrak{s}_{L}$. Then, on a boundary $\Lambda_{+}$ (resp. $\Lambda_{-}$) of $L$, an orientation and a spin structure $\mathfrak{s}_{+}$ (resp. $\mathfrak{s}_{-}$) are induced. See [17, Remark 3.6] for their precise descriptions.
Associated to $(\Lambda_{+},\mathfrak{s}_{+})$ and $(\Lambda_{-},\mathfrak{s}_{-})$, the Chekanov-Eliashberg DGAs
$\begin{array}[]{cc}(\mathcal{A}_{*}(\Lambda_{+}),\partial_{J,\mathfrak{s}_{+}}),&(\mathcal{A}_{*}(\Lambda_{-}),\partial_{J,\mathfrak{s}_{-}})\end{array}$
are defined as in Section 2.3. The maps $i_{+}$ and $i_{-}$ of (1) induce ring homomorphisms
$\begin{array}[]{ll}(i_{+})_{*}\colon\mathbb{Z}[H_{2}(Y,\Lambda_{+})]\to\mathbb{Z}[H_{2}(\mathbb{R}\times Y,L)],&(i_{-})_{*}\colon\mathbb{Z}[H_{2}(Y,\Lambda_{-})]\to\mathbb{Z}[H_{2}(\mathbb{R}\times Y,L)].\end{array}$
Via $(i_{-})_{*}$, we obtain a DGA over $\mathbb{Z}[H_{2}(\mathbb{R}\times Y,L)]$
$(\mathcal{A}_{*}(\Lambda_{-})\otimes_{\mathbb{Z}[H_{2}(Y,\Lambda_{-})]}\mathbb{Z}[H_{2}(\mathbb{R}\times Y,L)],\partial_{J,\mathfrak{s}_{-}}\otimes\operatorname{id}_{\mathbb{Z}[H_{2}(\mathbb{R}\times Y,L)]})$
and we regard it as a DGA over $\mathbb{Z}[H_{2}(Y,\Lambda_{+})]$ via $(i_{+})_{*}$.
By [8, Lemma 3.8], the moduli space $\mathcal{M}_{L,C}(a;b_{1},\dots,b_{m})$ is cut out transversely for generic $J$ and it is a manifold of dimension $|a|-(|b_{1}|+\dots+|b_{m}|)$ for any $a\in\mathcal{R}(\Lambda_{+})$, $b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda_{-})$ and $C\in H_{2}(\mathbb{R}\times Y,L)$. By the Gromov compactness, if $|b_{1}|+\dots+|b_{m}|=|a|$, $\coprod_{C\in H_{2}(\mathbb{R}\times Y,L)}\mathcal{M}_{L,C}(a;b_{1},\dots,b_{m})$ is a compact 00-dimensional manifold. In addition, it has an orientation depending on $\mathfrak{s}_{L}$ and a choice of an orientation of the determinant line of the capping operator for each Reeb chord of $\Lambda_{+}$ and $\Lambda_{-}$. See [17, Section A.3].
We refer to the proof of [17, Theorem 3.7] which takes $\mathbb{Z}[H_{1}(\Lambda_{\pm})]$ as the coefficient rings. We modify the arguments in [17, Appendix A.6] to the present case where we take $\mathbb{Z}[H_{2}(Y,\Lambda_{\pm})]$ as the coefficient rings. (As referred in [17, Appendix A], see also [11, Theorem 2.5] about the DGA map over $\mathbb{Z}$, but note that some conventions in [11] related to orientations are different from those in [5, 17].) Then, we obtain the following theorem.
Theorem 2.3.
We define a graded $\mathbb{Z}[H_{2}(Y,\Lambda_{+})]$-algebra map
$\Phi_{L}\colon\mathcal{A}_{*}(\Lambda_{+})\to\mathcal{A}_{*}(\Lambda_{-})\otimes_{\mathbb{Z}[H_{2}(Y,\Lambda_{-})]}\mathbb{Z}[H_{2}(\mathbb{R}\times Y,L)]$
such that for every $a\in\mathcal{R}(\Lambda_{+})$,
$\Phi_{L}(a)=\sum_{|b_{1}|+\dots+|b_{m}|=|a|}\sum_{C\in H_{2}(\mathbb{R}\times Y,L)}(-1)^{(n-1)(|a|+1)+m}\left(\#_{\operatorname{sign}}\mathcal{M}_{L,C}(a;b_{1},\dots,b_{m})\right)b_{1}\cdots b_{m}\otimes e^{C}.$
Then, $\Phi_{L}$ is a DGA map, that is, $\Phi_{L}\circ\partial_{J,\mathfrak{s}_{+}}=(\partial_{J,\mathfrak{s}_{-}}\otimes\operatorname{id}_{\mathbb{Z}[H_{2}(\mathbb{R}\times Y,L)]})\circ\Phi_{L}$.
2.5 Application to clean Lagrangian intersections
2.5.1 Chekanov-Eliashberg DGA of the unit conormal bundle of a knot in $\mathbb{R}^{3}$
Let $(q_{1},q_{2},q_{3})$ be the coordinate on $\mathbb{R}^{3}$. The manifold $\mathbb{R}^{3}$ is equipped with a standard Riemannian metric $\langle\cdot,\cdot\rangle\coloneqq\sum_{i=1}^{3}dq_{i}\otimes dq_{i}$ and an orientation such that $dq_{1}\wedge dq_{2}\wedge dq_{3}$ is a positive volume form. A coordinate $((q_{1},q_{2},q_{3}),(p_{1},p_{2},p_{3}))$ on $T^{*}\mathbb{R}^{3}$ is defined via the diffeomorphism
$\mathbb{R}^{3}\times\mathbb{R}^{3}\to T^{*}\mathbb{R}^{3}\colon((q_{1},q_{2},q_{3}),(p_{1},p_{2},p_{3}))\mapsto\sum_{i=1}^{3}p_{i}(dq_{i})_{(q_{1},q_{2},q_{3})}.$
For any $r>0$, we denote $D^{*}_{r}\mathbb{R}^{3}\coloneqq\{(q,p)\in T^{*}\mathbb{R}^{3}\mid|p|\leq r\}$. The canonical Liouville form on $T^{*}\mathbb{R}^{3}$ is given by $\lambda_{\mathbb{R}^{3}}=\sum_{i=1}^{3}p_{i}dq_{i}$.
We consider the unit cotangent bundle of $\mathbb{R}^{3}$
$S^{*}\mathbb{R}^{3}\coloneqq\{(q,p)\in T^{*}\mathbb{R}^{3}\mid|p|=1\}.$
It is endowed with a contact form $\alpha_{\mathbb{R}^{3}}\coloneqq\left.\lambda_{\mathbb{R}^{3}}\right|_{S^{*}\mathbb{R}^{3}}$. Note that there exists a diffeomorphism
$S^{*}\mathbb{R}^{3}\to T^{*}S^{2}\times\mathbb{R}\colon(q,p)\mapsto((p,q-\langle q,p\rangle p),\langle q,p\rangle)$
for the unit sphere $S^{2}\subset\mathbb{R}^{3}$ such that $\alpha_{\mathbb{R}^{3}}$ on $S^{*}\mathbb{R}^{3}$ agrees with the pullback of the $1$-form $dz-\lambda_{S^{2}}$ on $T^{*}S^{2}\times\mathbb{R}$, where $\lambda_{S^{2}}$ is the canonical Liouville form on $T^{*}S^{2}$. Therefore, by taking $(T^{*}S^{2},\lambda_{S^{2}})$ as $(P,\lambda)$, we can define the Chekanov-Eliashberg DGA for Legendrian submanifolds of $S^{*}\mathbb{R}^{3}$ as in Section 2.3.
We focus on a special kind of Legendrian submanifolds. For any knot $K$ in $\mathbb{R}^{3}$, recall the conormal bundle $L_{K}$ defined in Section 1. Then, the unit conormal bundle of $K$ is defined by
$\Lambda_{K}\coloneqq L_{K}\cap S^{*}\mathbb{R}^{3}=\{(q,p)\in S^{*}\mathbb{R}^{3}\mid q\in K,\ \langle p,v\rangle=0\text{ for every }v\in T_{q}K\}.$
It is a Legendrian submanifold of $S^{*}\mathbb{R}^{3}$.
Let us prepare several notations. We identify a disk bundle of $L_{K}$ with a tubular neighborhood of $K$ in $\mathbb{R}^{3}$ via the map $L_{K}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to\mathbb{R}^{3}\colon(q,p)\mapsto q+p$ for sufficiently small $\varepsilon>0$. Then, an orientation on $L_{K}$ is determined from the orientation on $\mathbb{R}^{3}$. We also fix a spin structure $\mathfrak{s}_{L_{K}}$ given by a trivial $\mathrm{Spin}(3)$-bundle on $L_{K}$. As the positive end of $L_{K}$, an orientation and a spin structure $\mathfrak{s}_{K}$ on $\Lambda_{K}$ are determined as in [17, Remark 3.6].
We fix an orientation on $K$. Let $l\colon K\to\Lambda_{K}$ be a section which gives the zero framing of $K$ and $m\colon S^{1}\to\Lambda_{K}$ be the meridian of $K$. The conventions are as follows: $S^{1}$ is oriented as the boundary of the unit disk $D$, and $(\nu,dl(v_{l}),dm(v_{m}))$ is a positive basis of $T_{q}\mathbb{R}^{3}$ at $q\in l(K)\cap m(S^{1})$, where $\nu\in T_{q}(L_{K}\cap D^{*}_{1}\mathbb{R}^{3})$ points outward and $v_{l}$ and $v_{m}$ are positive vectors tangent to $K$ and $S^{1}$ respectively. Then, the homology classes $[l]$ and $[m]$ give a basis of $H_{1}(\Lambda_{K})$.
We choose a non-vanishing vector field $V$ on $K$. Then, $\Lambda_{K}$ is bounded by a compact orientable manifold
$\{(q,p)\in S^{*}\mathbb{R}^{3}\mid q\in K,\ p(V_{q})\geq 0\}.$
Therefore, the homology class $[\Lambda_{K}]\in H_{2}(S^{*}\mathbb{R}^{3})$ vanishes. For the pair $(S^{*}\mathbb{R}^{3},\Lambda_{K})$, we have a short exact sequence
[…]
We choose a basis $\{\widehat{l},\widehat{m},\widehat{s}\}$ of $H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K})$ such that $\partial(\widehat{l})=[l]$, $\partial(\widehat{m})=[m]$ and $\widehat{s}$ is represented by the fiber $S^{*}_{q}\mathbb{R}^{3}$ of any $q\in\mathbb{R}^{3}$. For this choice, we refer to the remark after the proof of Theorem 1.2 in [4, Section 3.7], which means that we use a knot isotopic to $K$ and transversal to the contact structure $\operatorname{Ker}(dq_{3}-q_{1}dq_{2}+q_{2}dq_{1})$ on $\mathbb{R}^{3}$.
Remark 2.4.
From the above short exact sequence, the lifts of $[l]$ and $[m]$ to $H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K})$ can only be well-defined modulo $\mathbb{Z}\cdot\widehat{s}$. For the proof of the main results in Section 3, it is not important how to specify the lifts $\widehat{l}$ and $\widehat{m}$ (see Theorem 3.2).
We denote $\lambda\coloneqq e^{\widehat{l}}$, $\mu\coloneqq e^{\widehat{m}}$ and $U\coloneqq e^{\widehat{s}}$, then
$\mathbb{Z}[H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K})]=\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}].$
From the arguments in Section 2.3, the Chekanov-Eliashberg DGA $(\mathcal{A}_{*}(\Lambda_{K}),\partial_{J,\mathfrak{s}_{K}})$ for $(\Lambda_{K},\mathfrak{s}_{K})$ is defined with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$.
2.5.2 Exact Lagrangian cobordism arising from clean intersection
Given a $C^{\infty}$ Hamiltonian $H\colon T^{*}\mathbb{R}^{3}\times[0,1]\to\mathbb{R}$ with compact support, we define $(\varphi^{t}_{H})_{t\in[0,1]}$ to be the flow of a time-dependent vector field $(X_{H}^{t})_{t\in[0,1]}$ on $T^{*}\mathbb{R}^{3}$ characterized by $d\lambda_{\mathbb{R}^{3}}(\cdot,X^{t}_{H})=d(H(\cdot,t))$ for every $t\in[0,1]$. We denote
$\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})\coloneqq\{\varphi=\varphi^{1}_{H}\mid H\colon T^{*}\mathbb{R}^{3}\times[0,1]\to\mathbb{R}\text{ is a compactly supported }C^{\infty}\text{ Hamiltonian}\}.$
The image of the zero section of $T^{*}\mathbb{R}^{3}$ is identified with $\mathbb{R}^{3}$.
Let us recall the definition of a clean intersection of submanifolds.
Definition 2.5.
Let $L$ and $L^{\prime}$ be submanifolds of $T^{*}\mathbb{R}^{3}$. We say that $L$ and $L^{\prime}$ have a clean intersection along $L\cap L^{\prime}$ if the following hold:
  • $L\cap L^{\prime}$ is a submanifold of both $L$ and $L^{\prime}$.
  • For every $x\in L\cap L^{\prime}$, $T_{x}(L\cap L^{\prime})=T_{x}L\cap T_{x}L^{\prime}$.
Let us make a setup to deal with Question 1.1. Let $K_{0}$ and $K_{1}$ be oriented knots in $\mathbb{R}^{3}$ and tak $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$. Suppose that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along $K_{1}$. From [17, Lemma 2.3], we may assume that there exists $\varepsilon>0$ such that
$\varphi(L_{K_{0}})\cap D_{\varepsilon}^{*}\mathbb{R}^{3}=L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}$
by moving $\varphi(L_{K_{0}})$ by a Hamiltonian isotopy supported in a neighborhood of $\varphi(L_{K_{0}})\cap\mathbb{R}^{3}=K_{1}$. We consider a diffeomorphism
$F\colon\mathbb{R}\times S^{*}\mathbb{R}^{3}\to T^{*}\mathbb{R}^{3}\setminus\mathbb{R}^{3}\colon(r,(q,p))\mapsto(q,e^{r}p),$
for which $F^{*}(\lambda_{\mathbb{R}^{3}})=e^{r}\alpha_{\mathbb{R}^{3}}$. Furthermore,
$L_{\varphi}\coloneqq F^{-1}(\varphi(L_{K_{0}})\setminus K_{1})$
is an exact Lagrangian cobordism from $\Lambda_{K_{1}}$ to $\Lambda_{K_{0}}$. The inclusion maps (1) are written as
$i_{+}\colon(S^{*}\mathbb{R}^{3},\Lambda_{K_{0}})\to(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi})\colon(q,p)\mapsto(r_{+},(q,p)),$
$i_{-}\colon(S^{*}\mathbb{R}^{3},\Lambda_{K_{1}})\to(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi})\colon(q,p)\mapsto(r_{-},(q,p)),$
for some $r_{+},r_{-}\in\mathbb{R}$ with $r_{+}>r_{-}$.
The Lagrangian filling $\varphi(L_{K_{0}})$ of $\Lambda_{K_{0}}$ in $T^{*}\mathbb{R}^{3}$ is obtained by gluing the Lagrangian cobordism $L_{\varphi}$ in ×$\mathbb{R}\times S^{*}\mathbb{R}^{3}$ and the Lagrangian filling $L_{K_{1}}$ in $T^{*}\mathbb{R}^{3}$ along $\Lambda_{K_{1}}$. Therefore, as a part of the Mayer-Vietoris exact sequence, we have a short exact sequence
[…]
Note that $H_{2}(T^{*}\mathbb{R}^{3},L_{K_{1}})\cong H_{1}(L_{K_{1}})=\mathbb{Z}\cdot[K_{1}]$ and $H_{2}(T^{*}\mathbb{R}^{3},\varphi(L_{K_{0}}))\cong H_{1}(L_{K_{0}})=\mathbb{Z}\cdot[K_{0}]$. The left map is a surjection since $(i_{+})_{*}(\widehat{l})\in H_{2}(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi})$ is mapped to a generator of $H_{2}(T^{*}\mathbb{R}^{3},\varphi(L_{K_{0}}))$. It follows that $H_{2}(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi})$ is a free abelian group of rank $3$ with a basis $\{\widehat{l}_{0},\widehat{m}_{0},\widehat{s}_{0}\}$ defined by
$\widehat{l}_{0}$ $\coloneqq(i_{+})_{*}(\widehat{l})\text{ for }\widehat{l}\in H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K_{0}}),$
$\widehat{m}_{0}$ $\coloneqq(i_{-})_{*}(\widehat{m})\text{ for }\widehat{m}\in H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K_{1}}),$
$\widehat{s}_{0}$ [$\coloneqq[\{0\}\times S^{*}_{q}\mathbb{R}^{3}]\text{ for any }q\in\mathbb{R}^{3}.$
We abbreviate
$\begin{array}[]{ccc}R_{+}\coloneqq\mathbb{Z}[H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K_{0}})],&R_{-}\coloneqq\mathbb{Z}[H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K_{1}})],&R_{0}\coloneqq\mathbb{Z}[H_{2}(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi})].\end{array}$
The maps $i_{+}$ and $i_{-}$ induce ring homomorphisms $(i_{+})_{*}\colon R_{+}\to R_{0}$ and $(i_{-})_{*}\colon R_{-}\to R_{0}$. Let us denote
$\begin{array}[]{ccc}\lambda_{0}\coloneqq e^{\widehat{l}_{0}}=(i_{+})_{*}(\lambda),&\mu_{0}\coloneqq e^{\widehat{m}_{0}}=(i_{-})_{*}(\mu),&U_{0}\coloneqq e^{\widehat{s}_{0}}=(i_{+})_{*}(U)=(i_{-})_{*}(U).\end{array}$
Then $R_{0}=\mathbb{Z}[(\lambda_{0})^{\pm},(\mu_{0})^{\pm},(U_{0})^{\pm}]$.
Lemma 2.6.
$(i_{+})_{*}\colon R_{+}\to R_{0}$ and $(i_{-})_{*}\colon R_{-}\to R_{0}$ are ring isomorphisms. Moreover, There exist $a\in\{\pm 1\}$ and $b,c,d\in\mathbb{Z}$ such that the isomorphism $(i_{-})_{*}^{-1}\circ(i_{+})_{*}$ is given by
$(i_{-})_{*}^{-1}\circ(i_{+})_{*}\colon R_{+}\to R_{-}\colon\begin{cases}\lambda\mapsto\lambda^{a}\mu^{b}U^{c},\\ \mu\mapsto\mu^{a}U^{d},\\ U\mapsto U.\end{cases}$ (3)
To prove this lemma, we use the following result.
Proposition 2.7.
The open embedding $\varphi^{-1}\colon L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to L_{K_{0}}$ preserves the orientations of $L_{K_{1}}$ and $L_{K_{0}}$. Moreover, $\varphi_{*}([K_{0}])=a\cdot[K_{0}]$ in $H_{1}(\varphi(L_{K_{0}}))$ for $a\in\{1,-1\}$.
The proof is left to Appendix A. Note that the first assertion about orientations was shown in [17, Lemma 4.4] and a version of the second assertion with coefficients in $\mathbb{F}_{2}$ was shown by [17, Proposition 2.8, Lemma 4.5] using Floer cohomology group over $\mathbb{F}_{2}$ with the triangle product. It should be possible to prove Proposition 2.7 in a way parallel to [17, Section 2.3] by lifting the coefficient ring from $\mathbb{F}_{2}$ to $\mathbb{Z}$. However, we will give a simpler proof in Appendix A by using Floer cohomology with coefficients in $\mathbb{F}_{2}[H_{1}(L_{K_{0}})]\cong\mathbb{F}_{2}[\lambda^{\pm}]$.
Proof of Lemma 2.6.
It is clear that $(i_{\pm})_{*}\colon R_{\pm}\to R_{0}$ maps $U$ to $U_{0}$. The circle $\varphi(K_{0})$ is oriented via $\varphi\colon K_{0}\to\varphi(K_{0})$. From Proposition 2.7, we take $a\in\{\pm 1\}$ such that $[K_{1}]=a[\varphi(K_{0})]$. Consider the map induced by $i_{\pm}$ on the first homology groups
$\begin{array}[]{ll}(i_{+})_{*}\colon H_{1}(\Lambda_{K_{0}})\to H_{1}(L_{\varphi}),&(i_{-})_{*}\colon H_{1}(\Lambda_{K_{1}})\to H_{1}(L_{\varphi}).\end{array}$
A basis of $H_{1}(L_{\varphi})$ is given by $\{l_{0},m_{0}\}$ for $l_{0}\coloneqq(i_{+})_{*}([l])$ and $m_{0}\coloneqq(i_{-})_{*}([m])$. It suffices to show that for some $b\in\mathbb{Z}$,
$\begin{array}[]{cc}(i_{-})_{*}([l])=a\cdot l_{0}+b\cdot m_{0},&(i_{+})_{*}([m])=a\cdot m_{0}.\end{array}$ (5)
By the map $H_{1}(L_{\varphi})\to H_{1}(\varphi(L_{K_{0}}))$ induced by $\left.F\right|_{L_{\varphi}}\colon L_{\varphi}\to\varphi(L_{K_{0}})$, $l_{0}$ is mapped to $[\varphi(K_{0})]$ and $m_{0}$ is mapped to 00. Since $(\left.F\right|_{L_{\varphi}}\circ i_{-})_{*}([l])=[K_{1}]=a[\varphi(K_{0})]$ in $H_{1}(\varphi(L_{K_{0}}))$, the coefficient of $l_{0}$ in $(i_{-})_{*}([l])$ is $a$. This shows the first equation of (5) for some $b\in\mathbb{Z}$.
Next, take an embedded disk $D\subset\varphi(L_{K_{0}})$ which bounds the circle $(F\circ i_{+}\circ m)(S^{1})$. $D$ is oriented so that
$[\partial D]=(\left.F\right|_{L_{\varphi}})_{*}((i_{+})_{*}([m]))\text{ in }H_{1}(\varphi(L_{K_{0}})\setminus K_{1})$
for $\left.F\right|_{L_{\varphi}}\colon L_{\varphi}\to\varphi(L_{K_{0}})\setminus K_{1}$. We may assume that $D$ intersects $K_{1}$ transversely. Then, for sufficiently small $\varepsilon>0$, $D\cap(L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3})=\coprod_{i\in I}D^{\prime}_{i}$, where $D^{\prime}_{i}\subset D$ is a small disk and $I$ is a finite set. Then, using $a\in\{\pm 1\}$ determined by $[K_{1}]=a[\varphi(K_{0})]$, we have
$\sum_{i\in I}[\partial D_{i}]=a\cdot(\left.F\right|_{L_{\varphi}}\circ i_{-})_{*}([m])=a\cdot(\left.F\right|_{L_{\varphi}})_{*}(m_{0})\text{ in }H_{1}(\varphi(L_{K_{0}})\setminus K_{1}).$
For the first equality, note that the inclusion map $L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to\varphi(L_{K_{0}})$ preserves orientations by Proposition 2.7 when $\varphi(L_{K_{0}})$ is oriented via the map $\varphi\colon L_{K_{0}}\to\varphi(L_{K_{0}})$. Let $\Sigma$ be the closure of $F^{-1}(D\setminus\coprod_{i\in I}D^{\prime}_{i})$ in $L_{\varphi}$. Then,
$\begin{array}[]{cc}0=[\partial\Sigma]=(\left.F\right|_{L_{\varphi}})_{*}^{-1}([\partial D])-\sum_{i\in I}(\left.F\right|_{L_{\varphi}})^{-1}_{*}([\partial D_{i}])=(i_{+})_{*}([m])-a\cdot m_{0}&\text{in }H_{1}(L_{\varphi})\end{array}$
This shows the second equation of (5). ∎
The exact Lagrangian cobordism $L_{\varphi}$ from $\Lambda_{K_{1}}$ to $\Lambda_{K_{0}}$ induces a DGA map as below.
Proposition 2.8.
There exists a DGA map over $R_{+}$
$\Phi_{L_{\varphi}}\colon(\mathcal{A}_{*}(\Lambda_{K_{0}}),\partial_{J,\mathfrak{s}_{K_{0}}})\to(\mathcal{A}_{*}(\Lambda_{K_{1}}),\partial_{J,\mathfrak{s}_{K_{1}}}),$
where the $R_{-}$-algebra $\mathcal{A}_{*}(\Lambda_{K_{1}})$ is changed to an $R_{+}$-algebra via the isomorphism (3).
Proof.
$L_{K_{0}}$ is equipped with the spin structure $\mathfrak{s}_{L_{K_{0}}}$ given by a trivial $\mathrm{Spin}(3)$-bundle. It is restricted on $L_{K_{0}}\setminus\varphi^{-1}(K_{1})$, then via the diffeomorphism $F^{-1}\circ\varphi\colon L_{K_{0}}\setminus\varphi^{-1}(K_{1})\to L_{\varphi}$, we obtain a spin structure $s_{L_{\varphi}}$ on the exact Lagrangian cobordism $L_{\varphi}$. This induces $\mathfrak{s}_{K_{0}}$ on the positive end $\Lambda_{K_{0}}$ and $\mathfrak{s}_{K_{1}}$ on the negative end $\Lambda_{K_{0}}$. (Here, we need to care about the underlying orientations of spin manifolds. See the arguments in [17, Subsection 4.2.3].) Then, the existence of a DGA map $\Phi_{L_{\varphi}}$ follows from Theorem 2.3 and the description of the base change from $R_{-}$ to $R_{+}$ follows from Lemma 2.6. ∎
Lastly, let us introduce a spin structure $\mathfrak{s}^{0}_{K}$ on $\Lambda_{K}$ as in [7, Section 6.2]. When $T\Lambda_{K}$ is trivialized as an oriented vector bundle, then $\mathfrak{s}^{0}_{K}$ is given by the trivial $\mathrm{Spin}(2)$-bundle on $\Lambda_{K}$. It determines the orientations of moduli spaces which are used to define the differential of the DGAs in [4, Section 2.3.4] and [7, Section 3.6]. The ring isomorphism $\psi(\mathfrak{s}^{0}_{K},\mathfrak{s}_{K})$ from (2) is given by
$\psi(\mathfrak{s}^{0}_{K},\mathfrak{s}_{K})\colon\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]\to\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]\colon\lambda\mapsto\lambda,\ \mu\mapsto-\mu,\ U\mapsto U.$ (6)
3 Proof of main results
In this section, we abbreviate the Laurent polynomial ring $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$ by $R$.
3.1 General case
We will combine Proposition 2.8 with the results in [4, 15, 16] about the knot DGA. Some conventions about the knot DGA in these papers are slightly different from each other, as summarized in [16, Appendix A]. In this section, we mainly refer to [16] for the definition of the knot DGA.
Let $K$ be an oriented knot in $\mathbb{R}^{3}$ given by the closure of a braid $B$. Then, a DGA over $R$, called the knot DGA of $K$, is constructed from $B$ as in [16, Definition 3.11]. Its stable tame isomorphism class as a DGA over $R$ depends only on the isotopy class of $K$ as an oriented knot. See [15, Theorem 1.4] and [16, Theorem 3.1]. Let us denote this DGA by $(\mathcal{A}^{\mathrm{knot}}_{*}(K),\partial_{K})$ and its homology $\operatorname{Ker}\partial_{K}/\operatorname{Im}\partial_{K}$ by $HC_{*}(K)$. We will review a description of the 00-th degree part $HC_{0}(K)$ in Section 3.2.
Theorem 3.1 (Theorem 1.2 of [4], Theorem 3.14 and Appendix A of [16] ).
$(\mathcal{A}_{*}(\Lambda_{K}),\partial_{J,\mathfrak{s}^{0}_{K}})$ is stable tame isomorphic over $R$ to $(\mathcal{A}^{\mathrm{knot}}_{*}(K),\partial_{K})$ whose $R$-algebra structure is changed via a ring isomorphism
$\psi^{\prime}\colon R\to R\colon\lambda\mapsto-\lambda,\ \mu\mapsto-\mu,\ U\mapsto U.$
The base change by $\psi^{\prime}$ is due to the difference of conventions between [4] and [16]. See [16, Appendix A]. It follows from this theorem that there are unital graded ring homomorphisms
$\begin{array}[]{cc}I_{K}\colon\mathcal{A}_{*}(\Lambda_{K})\to\mathcal{A}^{\mathrm{knot}}_{*}(K),&\bar{I}_{K}\colon\mathcal{A}^{\mathrm{knot}}_{*}(K)\to\mathcal{A}_{*}(\Lambda_{K}),\end{array}$ (8)
intertwining $\partial_{K}$ and $\partial_{J,\mathfrak{s}^{0}_{K}}$ such that
$\begin{array}[]{cc}I_{K}(r\cdot x)=\psi^{\prime}(r)\cdot I_{K}(x),&\bar{I}_{K}(r\cdot y)=(\psi^{\prime})^{-1}(r)\cdot\bar{I}_{K}(y),\end{array}$
for every $r\in R$, $x\in\mathcal{A}_{*}(\Lambda_{K})$ and $y\in\mathcal{A}^{\mathrm{knot}}_{*}(K)$.
We continue to consider the setup in Subsection 2.5.2. Let $K_{0}$ and $K_{1}$ be oriented knots in $\mathbb{R}^{3}$. For $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$, suppose that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along $K_{1}$. Then we have the following theorem analogous to [17, Proposition 4.6].
Theorem 3.2.
There exist integers $a,b,c$ and $d$ with $a\in\{\pm 1\}$ and a unital graded ring homomorphism
$\Psi_{\varphi}\colon\mathcal{A}^{\mathrm{knot}}_{*}(K_{0})\to\mathcal{A}^{\mathrm{knot}}_{*}(K_{1})$
such that $\Psi_{\varphi}\circ\partial_{K_{0}}=\partial_{K_{1}}\circ\Psi_{\varphi}$ and
$\Psi_{\varphi}(r\cdot x)=\psi(r)\cdot\Psi_{\varphi}(x)$ (9)
for every $x\in\mathcal{A}^{\mathrm{knot}}_{*}(K_{0})$ and $r\in R$, where $\psi$ a ring isomorphism defined by
$\psi\colon R\to R\colon\lambda\mapsto\lambda^{a}\mu^{b}U^{c},\ \mu\mapsto\mu^{a}U^{d},\ U\mapsto U.$
Proof.
A unital graded ring homomorphism $\Psi_{\varphi}$ satisfying $\Psi_{\varphi}\circ\partial_{K_{0}}=\partial_{K_{1}}\circ\Psi_{\varphi}$ is defined as the composition of the following morphisms of differential graded rings:
[…]
Here, $\bar{I}_{K_{0}}$ and $I_{K_{1}}$ are the morphisms (8) from Theorem 3.1. $I(\mathfrak{s}^{0}_{K_{i}},\mathfrak{s}_{K_{i}})$ for $i\in\{0,1\}$ are the isomorphisms in Proposition 2.2. The vertical morphism $\Phi_{L_{\varphi}}$ is given by Proposition 2.8. From the relation between each morphism and the $R$-algebra structure on each DGA, the relation (9) holds if we define $\psi$ by
[…]
where $\psi^{\prime}$ is defined in Theorem 3.1 and $\psi(\mathfrak{s}^{0}_{K_{i}},\mathfrak{s}_{K_{i}})$ for $i\in\{0,1\}$ are given by (6). We compute that for $i\in\{0,1\}$
$\psi(\mathfrak{s}^{0}_{K_{i}},\mathfrak{s}_{K_{i}})\circ(\psi^{\prime})^{-1}\colon\lambda\mapsto-\lambda,\ \mu\mapsto\mu,\ U\mapsto U,$
and $\psi$ is described as in the theorem for the integers $a,b,c$ and $d$ which determine (3). ∎
We recall the notion of an augmentation of a DGA. See, for instance, [16, Definition 2.6].
Definition 3.3.
Let $\mathbb{k}$ be a field and $(\mathcal{A}_{*},\partial)$ be a DGA over $R$. A ring homomorphism $\varepsilon\colon\mathcal{A}_{*}\to\mathbb{k}$ is called an augmentation of $(\mathcal{A}_{*},\partial)$ to $\mathbb{k}$ if $\varepsilon(\mathbb{1})=1$, $\varepsilon(x)=0$ for any $x\in\mathcal{A}_{p}$ with $p\neq 0$, and $\varepsilon\circ\partial\colon\mathcal{A}_{1}\to\mathbb{k}$ is the zero map. Here, $\mathbb{1}$ is the unit of $\mathcal{A}_{*}$.
A knot invariant is defined from the set of augmentations of the knot DGA to $\mathbb{C}$.
Definition 3.4 (Definition 5.1 of [16]).
For an oriented knot $K$ in $\mathbb{R}^{3}$, its augmentation variety is a subset of $(\mathbb{C}^{*})^{3}$ defined by
$V_{K}\coloneqq\left\{(x,y,Z)\in(\mathbb{C}^{*})^{3}\ \middle|\begin{array}[]{l}\text{There exists an augmentation }\varepsilon\text{ of }(\mathcal{A}^{\mathrm{knot}}_{*}(K),\partial_{K})\text{ to }\mathbb{C}\\ \text{such that }\varepsilon(\lambda)=x,\ \varepsilon(\mu)=y\text{ and }\varepsilon(U)=Z\end{array}\right\}.$
Here, $\varepsilon(r\cdot\mathbb{1})$ for any $r\in R$ is abbreviated by $\varepsilon(r)$.
From Theorem 3.2, we obtain a refinement of [17, Theorem 1.1].
Corollary 3.5.
There exist integers $b^{\prime},c^{\prime}$ and $d^{\prime}$ for which we have an injective map
$V_{K_{1}}\hookrightarrow V_{K_{0}}\colon(x,y,Z)\mapsto(xy^{b^{\prime}}Z^{c^{\prime}},yZ^{d^{\prime}},Z).$
Proof.
By Theorem 3.2, for any augmentation $\varepsilon$ of $(\mathcal{A}_{*}^{\mathrm{knot}}(K_{1}),\partial_{K_{1}})$ to $\mathbb{C}$ with $(\varepsilon(\lambda),\varepsilon(\mu),\varepsilon(U))=(x,y,Z)$, the composite map $\varepsilon\circ\Psi_{\varphi}\colon\mathcal{A}_{*}^{\mathrm{knot}}(K_{0})\to\mathbb{C}$ is an augmentation of $(\mathcal{A}_{*}^{\mathrm{knot}}(K_{0}),\partial_{K_{0}})$ to $\mathbb{C}$ such that
$(\varepsilon\circ\Psi_{\varphi}(\lambda),\varepsilon\circ\Psi_{\varphi}(\mu),\varepsilon\circ\Psi_{\varphi}(U))=(x^{a}y^{b}Z^{c},y^{a}Z^{d},Z).$
From the definition of $V_{K_{i}}$ for $i\in\{0,1\}$, this implies that we have a map
$V_{K_{1}}\to V_{K_{0}}\colon(x,y,Z)\mapsto(x^{a}y^{b}Z^{c},y^{a}Z^{d},Z).$
We use the following fact by [15, Proposition 4.3]: Up to chain isomorphisms of DGA over $R$, $(\mathcal{A}_{*}^{\mathrm{knot}}(K_{0}),\partial_{K_{0}})$ is unchanged by the substitution $\lambda\mapsto\lambda^{-1}U$, $\mu\mapsto\mu^{-1}U$ and $U\mapsto U$. Hence, we have a bijection
$V_{K_{0}}\to V_{K_{0}}\colon(x,y,Z)\mapsto(x^{-1}Z,y^{-1}Z,Z).$
If $a=1$, we take $b^{\prime}\coloneqq b$, $c^{\prime}\coloneqq c$ and $d^{\prime}\coloneqq d$. If $a=-1$, we take $b^{\prime}\coloneqq-b$, $c^{\prime}\coloneqq-c+1$ and $d^{\prime}\coloneqq-d+1$. Then, the assertion of the corollary holds. ∎
3.2 Proof of Theorem 1.1
First, let us review how to compute $HC_{0}(K)$ for an oriented knot $K$. We refer to [16, Section 3]. Note that since $\mathcal{A}^{\mathrm{knot}}_{p}(K)=0$ for $p<0$, $HC_{0}(K)=\mathcal{A}^{\mathrm{knot}}_{0}(K)/\partial_{K}(\mathcal{A}^{\mathrm{knot}}_{1}(K))$.
For any $n\in\mathbb{Z}_{\geq 1}$, let $B_{n}$ denote the group of braids on $n$ strands, which is generated by $\{\sigma_{1},\dots,\sigma_{n-1}\}$ with relations $\sigma_{i}\sigma_{i+1}\sigma_{i}=\sigma_{i+1}\sigma_{i}\sigma_{i+1}$ for $21\leq i\leq n-$ 2 and $\sigma_{i}\sigma_{j}=\sigma_{j}\sigma_{i}$ for $|i-j|\geq 2$. Let $\mathcal{A}_{n}$ be the $R$-algebra freely generated by the set $\{a_{i,j}\mid i,j\in\{1,\dots,n\},\ i\neq j\}$. Then, a group homomorphism $B_{n}\to\mathrm{Aut}(\mathcal{A}_{n})\colon B\mapsto\phi_{B}$ is defined as in [16, Definition 3.3]. It is determined by $\phi_{\sigma_{k}}\colon\mathcal{A}_{n}\to\mathcal{A}_{n}$ for each generator $\sigma_{k}$, and $\phi_{\sigma_{k}}$ is given by
$\phi_{\sigma_{k}}\colon\begin{cases}a_{i,j}\mapsto a_{i,j}&\text{if }i,j\neq k,k+1,\\ a_{k+1,i}\mapsto a_{k,i}&\text{ if }i\neq k,k+1,\\ a_{i,k+1}\mapsto a_{i,k}&\text{ if }i\neq k,k+1,\\ a_{k,k+1}\mapsto-a_{k+1,k},&\\ a_{k+1,k}\mapsto-a_{k,k+1},&\\ a_{k,i}\mapsto a_{k+1,i}-a_{k+1,k}a_{k,i}&\text{ if }i\neq k,k+1,\\ a_{i,k}\mapsto a_{i,k+1}-a_{i,k}a_{k,k+1}&\text{ if }i\neq k,k+1.\end{cases}$ (10)
See also [7, Section 2.1]. We remark that $\phi_{B\cdot B^{\prime}}=\phi_{B}\circ\phi_{B^{\prime}}$ for any $B,B^{\prime}\in B_{n}$.
The automorphism $\phi_{B}\in\mathrm{Aut}(\mathcal{A}_{n})$ for $B\in B_{n}$ is extended to an automorphism on $\mathcal{A}_{n+1}$ as follows: We consider the group homomorphism $B_{n+1}\to\mathrm{Aut}(\mathcal{A}_{n+1})$ defined as above and an injective map $B_{n}\to B_{n+1}$ defined by adding the extra $(n+1)$-th strand to any braid on $n$ strands (see [16, Remark 3.5]). Then, for any $B\in B_{n}$, $\phi_{B}\in\mathrm{Aut}(\mathcal{A}_{n+1})$ is defined via the composite map $B_{n}\hookrightarrow B_{n+1}\to\mathrm{Aut}(\mathcal{A}_{n+1})$. In addition, let us label the $(n+1)$-st strand by * and denote $a_{i,n+1}$ by $a_{i,*}$ and $a_{n+1,i}$ by $a_{*,i}$ for every $i\in\{1,\dots,n\}$.
We define $n\times n$ matrices $\mathbb{A}$ and $\widehat{\mathbb{A}}$ by
$\begin{array}[]{ll}\mathbb{A}_{i,j}\coloneqq\begin{cases}a_{i,j}&\text{ if }i<j,\\ 1-\mu&\text{ if }i=j,\\ -\mu a_{i,j}&\text{ if }i>j,\end{cases}&\widehat{\mathbb{A}}_{i,j}\coloneqq\begin{cases}Ua_{i,j}&\text{ if }i<j,\\ U-\mu&\text{ if }i=j,\\ -\mu a_{i,j}&\text{ if }i>j.\end{cases}\end{array}$
Fix $B\in B_{n}$ which has the writhe $w$ (the sum of the exponents in the braid word). Let $\bm{\Lambda}$ be an $n\times n$ diagonal matrix with $\bm{\Lambda}_{1,1}=\lambda\mu^{w}U^{-(w-n+1)/2}$ and $\bm{\Lambda}_{i,i}=1$ for every $i\in\{2,\dots,n\}$. Let $\bm{\Phi}^{L}_{B}$ be an $n\times n$ matrix whose entries $(\bm{\Phi}^{L}_{B})_{i,j}\in\mathcal{A}_{n}$ for $i,j\in\{1,\dots,n\}$ are determined by
$\phi_{B}(a_{i,*})=\sum_{j=1}^{n}(\bm{\Phi}^{L}_{B})_{i,j}a_{j,*}.$
Let $K$ be the oriented knot in $\mathbb{R}^{3}$ given by the closure of the braid $B$. By [16, Theorem 4.1, Remark 4.2], $HC_{0}(K)$ is the quotient of $\mathcal{A}_{n}$ by the two-sided ideal generated by all entries of the two matrices
$\begin{array}[]{cc}\mathbb{A}-\bm{\Lambda}\cdot\phi_{B}(\mathbb{A})\cdot\bm{\Lambda}^{-1},&\widehat{\mathbb{A}}-\bm{\Lambda}\cdot\bm{\Phi}^{L}_{B}\cdot\mathbb{A},\end{array}$ (12)
where $(\phi_{B}(\mathbb{A}))_{i,j}\coloneqq\phi_{B}(\mathbb{A}_{i,j})$ for $i,j\in\{1,\dots,n\}$. The isomorphism class of $HC_{0}(K)$ as a unital $R$-algebra depends only on the isotopy class of $K$. See [16, Theorem 3.15] and [15, Theorem 1.4 (2)].
We consider a commutative $R$-algebra $HC_{0}^{\mathrm{ab}}(K)$ defined as the quotient of $HC_{0}(K)$ by the relations $\xi\eta=\eta\xi$ for all $\xi,\eta\in HC_{0}(K)$. Using the $R$-algebra $R[a_{i,j}\mid 1\leq i,j\leq n,\ i\neq j]$ of polynomials and a natural map $\pi_{n}\colon\mathcal{A}_{n}\to R[a_{i,j}\mid 1\leq i,j\leq n,\ i\neq j]\colon a_{i,j}\mapsto a_{i,j}$, we have
$HC_{0}^{\mathrm{ab}}(K)=R[a_{i,j}\mid 1\leq i,j\leq n,\ i\neq j]/\mathcal{I}_{B},$
where $\mathcal{I}_{B}$ is the ideal generated by the images under $\pi_{n}$ of all entries of the two matrices (12).
[…]
Figure. 1: The braid $B^{\#}\in B_{n+2}$. The box with ‘$B$’ in the center represents the braid $B$ on $n$ strands.
Associated to $B$, consider a braid $B^{\#}$ on $(n+2)$ strands defined by
$B^{\#}\coloneqq\sigma_{n}\cdot(\sigma_{n+1})^{3}\cdot B\in B_{n+2}.$
See Figure 1. The closure of $B^{\#}$ gives a knot which is the connected sum $K\#3_{1}$ of $K$ and the right-handed trefoil knot $3_{1}$.
In general, $HC^{\mathrm{ab}}_{0}(K\#3_{1})$ is a complicated $R$-algebra. However, it has the following property.
Lemma 3.6.
We define an $R$-coefficient polynomial $F\in R[X]$ by
$F(X)\coloneqq X^{2}+(\mu+1-\mu^{-1}U)X+(\mu-1).$
Then, there exists a unital $R$-algebra map
$J\colon R[X]/(F)\to HC^{\mathrm{ab}}_{0}(K\#3_{1}).$
Proof.
We compute $\phi_{B^{\#}}(a_{n+2,*})$ by (10). Since $\phi_{\sigma_{k}}(a_{n+2,*})=a_{n+2,*}$ for every $k\in\{1,\dots,n-1\}$, $\phi_{B}(a_{n+2,*})=a_{n+2,*}$. Furthermore,
${a_{n+2,*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{B}}$${a_{n+2,*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{\sigma_{n+1}}}$${a_{n+1,*}}$$\scriptstyle{\phi_{\sigma_{n+1}}}$${a_{n+2,*}-a_{n+2,n+1}a_{n+1,*}}$$\scriptstyle{\phi_{\sigma_{n+1}}}$${a_{n+1,*}+a_{n+1,n+2}(a_{n+2,*}-a_{n+2,n+1}a_{n+1,*})}$$\scriptstyle{\phi_{\sigma_{n}}}$${a_{n,*}+a_{n,n+2}(a_{n+2,*}-a_{n+2,n}a_{n,*}).}$
Let us write $P=a_{n,n+2}$ and $Q=a_{n+2,n}$. Then, the bottom $(n+2)$-th row of $\bm{\Phi}^{L}_{B^{\#}}$ is
$\begin{pmatrix}0&\cdots&0&1-PQ&0&P\end{pmatrix}.$
It follows that the $(n+2,n)$-th entry of $\widehat{\mathbb{A}}-\bm{\Lambda}\cdot\bm{\Phi}^{L}_{B^{\#}}\cdot\mathbb{A}$ is
$\widehat{\mathbb{A}}_{n+2,n}-((1-PQ)\mathbb{A}_{n,n}+P\mathbb{A}_{n+2,n})$ $=-\mu Q-((1-PQ)(1-\mu)+P(-\mu Q))$
$=PQ-\mu Q-(1-\mu),$
and the $(n+2,n+2)$-th entry of $\widehat{\mathbb{A}}-\bm{\Lambda}\cdot\bm{\Phi}^{L}_{B^{\#}}\cdot\mathbb{A}$ is
$\widehat{\mathbb{A}}_{n+2,n+2}-((1-PQ)\mathbb{A}_{n,n+2}+P\mathbb{A}_{n+2,n+2})$ $=(U-\mu)-((1-PQ)P+P(1-\mu))$
$=PQP-(2-\mu)P+(U-\mu).$
Hence $P^{2}Q^{2}-(2-\mu)PQ+(U-\mu)Q\in\mathcal{I}_{B^{\#}}$ and by substituting $PQ$ with $\mu Q+(1-\mu)$, we can find an element of $\mathcal{I}_{B^{\#}}$
$(\mu Q+(1-\mu))^{2}-(2-\mu)(\mu Q+(1-\mu))+(U-\mu)Q,$
which is equal to
$\mu^{2}Q^{2}+(2\mu(1-\mu)-(2-\mu)\mu+(U-\mu))Q+(1-\mu)^{2}-(2-\mu)(1-\mu)$
=\ = $\mu^{2}Q^{2}+(-\mu^{2}-\mu+U)Q+(\mu-1).$
Therefore, if we define $F(X)=X^{2}+(\mu+1-\mu^{-1}U)X+(\mu-1)$, we have a well-defined unital $R$-algebra map
$J\colon R[X]/(F)\to HC^{\mathrm{ab}}_{0}(K\#3_{1})$
which sends $X$ to $-\mu Q=-\mu a_{n+2,n}$. ∎
Let us check this property in the simplest case.
Example 3.7.
From a braid representation $\sigma_{1}^{3}\in B_{2}$ of $3_{1}$, it is not difficult to compute $HC_{0}(3_{1})$ as outlined in [16, Exercise 3.10]. $HC_{0}(3_{1})$ is isomorphic to a commutative $R$-algebra $R[Y]/(G_{1},G_{2})$, where
$\begin{array}[]{cc}G_{1}(Y)\coloneqq UY^{2}-\mu UY+\lambda\mu^{3}(1-\mu),&G_{2}(Y)\coloneqq UY^{2}+\lambda\mu^{2}Y+\lambda\mu^{2}(\mu-U).\end{array}$
See [16, Exercise 5.5]. We have identities
$YG_{1}+\mu G_{2}$ $=UY^{3}+\lambda\mu^{3}(2-\mu)Y+\lambda\mu^{3}(\mu-U),$
$UY(YG_{1}+\mu G_{2})$ $=(UY^{2})^{2}+\lambda\mu^{3}(2-\mu)(UY^{2})+\lambda\mu^{3}(\mu-U)UY,$
$UY^{2}$ $=G_{1}+(\mu UY-\lambda\mu^{3}(1-\mu)),$
so the ideal $(G_{1},G_{2})$ has an element
$(\mu UY-\lambda\mu^{3}(1-\mu))^{2}+\lambda\mu^{3}(2-\mu)(\mu UY-\lambda\mu^{3}(1-\mu))+\lambda\mu^{3}(\mu-U)UY$
=\ = $(\mu UY)^{2}+(\lambda\mu^{4}+\lambda\mu^{3}-\lambda\mu^{2}U)(\mu UY)+(\lambda\mu^{3})^{2}(\mu-1).$
This means that a unital $R$-algebra map
$R[X]/(F)\to HC^{\mathrm{ab}}_{0}(3_{1})\cong R[Y]/(G_{1},G_{2})\colon X\mapsto\lambda^{-1}\mu^{-2}UY=(\lambda\mu^{3})^{-1}(\mu UY)$
is well-defined for the polynomial $F$ in Lemma 3.6, since
$(\lambda\mu^{3})^{2}F(\lambda^{-1}\mu^{-2}UY)=(\mu UY)^{2}+(\lambda\mu^{3})(\mu+1-\mu^{-1}U)(\mu UY)+(\lambda\mu^{3})^{2}(\mu-1)\in(G_{1},G_{2}).$
Remark 3.8.
If we reduce the coefficient ring $R$ to $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$ by setting $U=1$, Lemma 3.6 implies the existence of a unital $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra map
$\mathbb{Z}[\lambda^{\pm},\mu^{\pm}][X]/(\left.F\right|_{U=1})\to\left.HC^{\mathrm{ab}}_{0}(K\#3_{1})\right|_{U=1}$
induced by $J$. (Here, the notation $\left.HC_{0}(K\#3_{1})\right|_{U=1}$ for an $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra is the same one as in [16, Section 4].) However, this assertion is trivial. Indeed,
$\left.F\right|_{U=1}(X)=X^{2}+(\mu+1-\mu^{-1})X+(\mu-1)=(X+\mu)(X+1-\mu^{-1})$
and thus, for any unital $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra $A$, there exists a unital $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra map $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}][X]/(\left.F\right|_{U=1})\to A\colon X\mapsto-\mu$.
Lemma 3.6 implies the following constraint on augmentations.
Lemma 3.9.
Let $\mathbb{k}$ be a field. For any augmentation $\varepsilon$ of $(\mathcal{A}^{\mathrm{knot}}_{*}(K\#3_{1}),\partial_{K\#3_{1}})$ to $\mathbb{k}$,
$D_{\varepsilon}\coloneqq(\varepsilon(\mu)+1-\varepsilon(\mu)^{-1}\varepsilon(U))^{2}-4(\varepsilon(\mu)-1)$
satisfies $D_{\varepsilon}\in\{k^{2}\mid k\in\mathbb{k}\}$.
Proof.
Since $HC_{0}(K\#3_{1})=\mathcal{A}^{\mathrm{knot}}_{0}(K\#3_{1})/\partial_{K\#3_{1}}(\mathcal{A}^{\mathrm{knot}}_{1}(K\#3_{1}))$, any augmentation $\varepsilon$ to $\mathbb{k}$ of $(\mathcal{A}^{\mathrm{knot}}_{*}(K\#3_{1}),\partial_{K\#3_{1}})$ induces a ring homomorphism
$\bar{\varepsilon}\colon HC^{\mathrm{ab}}_{0}(K\#3_{1})\to\mathbb{k}$
such that $\bar{\varepsilon}(r)=\varepsilon(r)$ for every $r\in R$. Composing with $J$ in Lemma 3.6, we obtain a map
$\bar{\varepsilon}\circ J\colon R[X]/(F)\to\mathbb{k}.$
Let us denote $\zeta=\bar{\varepsilon}\circ J(X)$. Since $\bar{\varepsilon}\circ J(F)=0$ and $J$ is a unital $R$-algebra map, we obtain an equation in $\mathbb{k}$
$\zeta^{2}+(\varepsilon(\mu)+1-\varepsilon(\mu)^{-1}\varepsilon(U))\zeta+(\varepsilon(\mu)-1)=0.$
Then, $D_{\varepsilon}$ in the lemma is the discriminant of this quadratic equation for $\zeta$. In order that this equation has a solution in $\mathbb{k}$, it is necessary that $D_{\varepsilon}$ is the square of some element of $\mathbb{k}$. ∎
Combining Theorem 3.2, Lemma 3.9 and a computation of the knot DGA of the unknot, we obtain a constraint on the isotopy class of clean Lagrangian intersections.
Theorem 3.10.
Let $K_{0}$ be the connected sum of an arbitrary knot $K$ in $\mathbb{R}^{3}$ with $3_{1}$ or its mirror $\overline{3_{1}}$. Then, there is no $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along the unknot in $\mathbb{R}^{3}$.
Proof.
Assume that there exists $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection $K_{1}=\varphi(L_{K_{0}})\cap\mathbb{R}^{3}$ and $K_{1}$ is the unknot in $\mathbb{R}^{3}$. In the case of $K_{0}=K\#3_{1}$, we can deduce a contradiction as follows: From the trivial braid on a single strand, we can compute as in [16, Example 3.13] that $\mathcal{A}^{\mathrm{knot}}_{0}(K_{1})=R$ and $\partial_{K_{1}}(\mathcal{A}_{1}^{\mathrm{knot}}(K_{1}))$ is the ideal generated by
$U-\lambda-\mu+\lambda\mu.$
Let us choose an augmentation $\varepsilon_{1}\colon\mathcal{A}^{\mathrm{knot}}_{*}(K_{1})\to\mathbb{Q}$ determined by
$\begin{array}[]{ccc}\varepsilon_{1}(\lambda)=3,&\varepsilon_{1}(\mu)=2,&\varepsilon_{1}(U)=-1.\end{array}$
Composing with the map $\Psi_{\varphi}$ in Theorem 3.2, we obtain $\varepsilon\coloneqq\varepsilon_{1}\circ\Psi_{\varphi}\colon\mathcal{A}^{\mathrm{knot}}_{*}(K_{0})\to\mathbb{Q}$. It is an augmentation of the knot DGA of $K_{0}=K\#3_{1}$ to $\mathbb{Q}$ such that
$\begin{array}[]{ccc}\varepsilon(\lambda)=3^{a}2^{b}(-1)^{c},&\varepsilon(\mu)=2^{a}(-1)^{d},&\varepsilon(U)=-1.\end{array}$
We apply Lemma 3.9 for $\mathbb{k}=\mathbb{Q}$. Then,
$D_{\varepsilon}=(\varepsilon(\mu)+1+\varepsilon(\mu)^{-1})^{2}-4(\varepsilon(\mu)-1)$
must be the square of some rational number. However, we can calculate as follows:
$D_{\varepsilon}=\begin{cases}(\frac{7}{2})^{2}-4\cdot 1=\frac{33}{4}&\text{ if}\varepsilon(\mu)=2,\\ (-\frac{3}{2})^{2}-4\cdot(-3)=\frac{57}{4}&\text{ if }\varepsilon(\mu)=-2,\\ (\frac{7}{2})^{2}-4\cdot(-\frac{1}{2})=\frac{57}{4}&\text{ if }\varepsilon(\mu)=\frac{1}{2},\\ (-\frac{3}{2})^{2}-4\cdot(-\frac{3}{2})=\frac{33}{4}&\text{ if }\varepsilon(\mu)=-\frac{1}{2},\end{cases}$
and this is a contradiction.
In the case of $K_{0}=K\#\overline{3_{1}}$, we can deduce a contradiction as follows: Take a Hamiltonian $H\colon T^{*}\mathbb{R}^{3}\times[0,1]\to\mathbb{R}$ such that $\varphi=\varphi^{1}_{H}$. Consider an involution
$\iota\colon T^{*}\mathbb{R}^{3}\to T^{*}\mathbb{R}^{3}\colon(q,p)=((q_{1},q_{2},q_{3}),(p_{1},p_{2},p_{3}))\mapsto((q_{1},q_{2},-q_{3}),(p_{1},p_{2},-p_{3}))$
which preserves $\lambda_{\mathbb{R}^{3}}$. Then, $\iota\circ\varphi^{t}_{H}\circ\iota=\varphi^{t}_{\overline{H}}$ for $\overline{H}\coloneqq H\circ(\iota\times\operatorname{id}_{[0,1]})$. Moreover, for the knots
$\overline{K_{i}}\coloneqq\{(q_{1},q_{2},-q_{3})\in\mathbb{R}^{3}\mid(q_{1},q_{2},q_{3})\in K_{i}\}\text{ for }i=0,1,$
$\varphi^{1}_{\overline{H}}(L_{\overline{K_{0}}})$ and $\mathbb{R}^{3}=\iota(\mathbb{R}^{3})$ intersect cleanly along $\overline{K_{1}}$. Since $\overline{K_{1}}$ is the unknot and $\overline{K_{0}}$ is the connected sum of the mirror of $K$ and $3_{1}$, this contradicts to the result in the previous case. ∎
Remark 3.11.
For the proof of the theorem, we may choose other field $\mathbb{k}$ and other augmentation $\varepsilon_{1}$ of the knot DGA of $K\#3_{1}$ to $\mathbb{k}$ such that $\varepsilon(U)\neq 1$. Let us see two examples.
  1. 1.
    Choose $\mathbb{k}=\mathbb{Q}[\sqrt{-1}]$ and $\varepsilon_{1}$ determined by
    $\begin{array}[]{ccc}\varepsilon_{1}(\lambda)=2-\sqrt{-1},&\varepsilon_{1}(\mu)=2,&\varepsilon_{1}(U)=\sqrt{-1},\end{array}$
    in $\mathbb{Q}[\sqrt{-1}]$. For $\varepsilon\coloneqq\varepsilon_{1}\circ\Psi_{\varphi}$, let us write $D_{\varepsilon}=D_{y}$ if $\varepsilon(\mu)=y$. Then,
    $D_{y}=(y+1-\sqrt{-1}y^{-1})^{2}-4(y-1)=y^{2}-2y-2\sqrt{-1}y^{-1}-y^{-2}+(5-2\sqrt{-1}),$
    and $y$ varies in the set {±$\{\pm 2,\pm 2\sqrt{-1},\pm\frac{1}{2},\pm\frac{\sqrt{-1}}{2}\}$. Noting that $D_{\sqrt{-1}y^{-1}}=D_{y}$, we can compute that $D_{y}$ takes its value in the set
    ${S\coloneqq\left\{\frac{19}{4}-3\sqrt{-1},\frac{51}{4}-\sqrt{-1},\frac{1}{4}-6\sqrt{-1},\frac{9}{4}+2\sqrt{-1}\right\}.}$
    In general, suppose that $D=a+\sqrt{-1}b$ for $a,b\in\mathbb{Q}$ satisfies $D\in\{k^{2}\mid k\in\mathbb{Q}[\sqrt{-1}]\}$. Then, $a^{2}+b^{2}=D\cdot\bar{D}\in\{r^{2}\mid r\in\mathbb{Q}\}$, where $\bar{D}$ is the complex conjugate of $D$. However, for any $D=a+\sqrt{-1}b\in S$, we can check that $a^{2}+b^{2}\notin\{r^{2}\mid r\in\mathbb{Q}\}$. Therefore, $D_{y}\notin\{k^{2}\mid k\in\mathbb{Q}[\sqrt{-1}]\}$ for any $y\in\{\pm 2,\pm 2\sqrt{-1},\pm\frac{1}{2},\pm\frac{\sqrt{-1}}{2}\}$.
  2. 2.
    Choose $\mathbb{k}=\mathbb{F}_{5}\cong\mathbb{Z}/5$ and $\varepsilon_{1}$ determined by
    $\begin{array}[]{ccc}\varepsilon_{1}(\lambda)=3,&\varepsilon_{1}(\mu)=2,&\varepsilon_{1}(U)=-1,\end{array}$
    in $\mathbb{F}_{5}$, similar to the one in the proof of Theorem 3.10. In general, if $D\in\mathbb{Z}/5$ satisfies $D\equiv k^{2}\mod 5$ for some $k\in\mathbb{Z}$, then $D\equiv 0,1$ or $4\mod 5$. However, for $\varepsilon\coloneqq\varepsilon_{1}\circ\Psi_{\varphi}$,
    $D_{\varepsilon}\equiv 33\cdot 4\equiv 2\text{ or }D_{\varepsilon}\equiv 57\cdot 4\equiv 3\mod 5.$
    Therefore, $D_{\varepsilon}\notin\{k^{2}\mid k\in\mathbb{F}_{5}\}$.
Remark 3.12.
To the author’s knowledge, it seems difficult to show Theorem 3.10, even when $K_{0}=3_{1}$, by only using the knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$. Indeed, for the trefoil knot $3_{1}$ and the unknot $O$,
$\left.HC_{0}(3_{1})\right|_{U=1}$ $\cong\mathbb{Z}[\lambda^{\pm},\mu^{\pm}][Y]/(Y^{2}-\mu Y+\lambda\mu^{3}(1-\mu),\ Y^{2}+\lambda\mu^{2}Y+\lambda\mu^{2}(\mu-1)),$
$\left.HC_{0}(O)\right|_{U=1}$ $\cong\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]/((1-\lambda)(1-\mu)),$
and we can find a unital $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra map
$\left.HC_{0}(3_{1})\right|_{U=1}\to\left.HC_{0}(O)\right|_{U=1}\colon Y\mapsto\mu-\mu^{2}.$
Therefore, we cannot deduce a contradiction from the existence of a unital $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra map which is induced on the 00-the degree part of the homology from the DGA map as in Theorem 3.2.
A Proof of Proposition 2.7
Let $K_{0}$ be an oriented knot in $\mathbb{R}^{3}$ and consider its conormal bundle $L_{K_{0}}$. The Lagrangian submanifolds $\mathbb{R}^{3}$ and $L_{K_{0}}$ of $T^{*}\mathbb{R}^{3}$ are oriented as in Subsection 2.5.1.
For any $\psi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$, we introduce a $\mathbb{Z}/2$-graded Floer cochain complex
$(CF^{*}(\psi(\mathbb{R}^{3}),L_{K_{0}};H),d_{J})$
with coefficients in the Laurent polynomial ring $\mathbb{F}_{2}[\lambda^{\pm}]$. It is a slight enhancement of the Floer cochain complex over $\mathbb{F}_{2}$ constructed as in [20, Section 8].
We choose a Hamiltonian $H\colon T^{*}\mathbb{R}^{3}\times[0,1]\to\mathbb{R}$ with compact support such that the Lagrangian submanifold $P_{H}\coloneqq(\varphi^{1}_{H}\circ\psi)(\mathbb{R}^{3})$ intersects $L_{K_{0}}$ transversely. $P_{H}$ is oriented so that the diffeomorphism $\varphi^{1}_{H}\circ\psi\colon\mathbb{R}^{3}\to P_{H}$ preserves orientations. From the orientations of $P_{H}$ and $L_{K_{0}}$, a $\mathbb{Z}/2$-grading $|x|\in\mathbb{Z}/2$ is defined for every $x\in P_{H}\cap L_{K_{0}}$. See [19, Example 2.7, Section 2.d]. Note that the grading is shifted by $1$ if we use the opposite orientation of $P_{H}$. We also choose an almost complex structure $J$ such that $d\lambda_{\mathbb{R}^{3}}(\cdot,J\cdot)$ is a Riemannian metric on $T^{*}\mathbb{R}^{3}$ and $J(\partial_{p_{i}})=\partial_{q_{i}}$ for $i\in\{1,2,3\}$ outside a compact subset of $T^{*}\mathbb{R}^{3}$.
[…]
Figure. 2: The shaded region is the image of $u$. $c_{x}$ and $c_{y}$ are paths in $L_{K_{0}}$ and $u(\mathbb{R}\times\{1\})$ is contained in $L_{K_{0}}$. $u(\mathbb{R}\times\{0\})$ is contained in $\psi(\mathbb{R}^{3})$.
Fix a base point $z_{0}\in L_{K_{0}}$. For any $x\in P_{H}\cap L_{K_{0}}$, we choose a path $c_{x}\colon[0,1]\to L_{K_{0}}$ such that $c_{x}(0)=z_{0}$ and $c_{x}(1)=x$. Let $c_{x}^{-1}$ denote the path $[0,1]\to L_{K_{0}}\colon t\mapsto c_{x}(1-t)$. For every $x,y\in P_{H}\cap L_{K_{0}}$ and a homology class $A\in H_{1}(L_{K_{0}})$, we define $\mathcal{M}_{(H,J),A}(x,y)$ to be the space of maps $u\colon\mathbb{R}\times[0,1]\to T^{*}\mathbb{R}^{3}$ satisfying:
  • su($\partial_{s}u(s,t)+J_{u(s,t)}(\partial_{t}u(s,t)-(X^{t}_{H})_{u(s,t)})=0$ for every $(s,t)\in\mathbb{R}\times[0,1]$.
  • $u(\mathbb{R}\times\{0\})\subset\psi(\mathbb{R}^{3})$ and $u(\mathbb{R}\times\{1\})\subset L_{K_{0}}$.
  • $\lim_{s\to-\infty}u(s,t)=\varphi^{t}_{H}((\varphi^{1}_{H})^{-1}(x))$ and $\lim_{s\to\infty}u(s,t)=\varphi_{H}^{t}((\varphi^{1}_{H})^{-1}(y))$ $C^{\infty}$-uniformly on $t\in[0,1]$.
  • $A\in H_{1}(L_{K_{0}})$ is represented by a loop based at $z_{0}$ obtained by concatenating the three paths in $L_{K_{0}}$: $c_{x}$, the path $[-\infty,\infty]\to L_{K_{0}}\colon s\mapsto u(s,1)$ from $x$ to $y$, and $c_{y}^{-1}$.
Figure 2 depicts $u(\mathbb{R}\times[0,1])\subset T^{*}\mathbb{R}^{3}$ and the three paths in $L_{K_{0}}$. The group $\mathbb{R}$ acts on $\mathcal{M}_{(H,J),A}(x,y)$ by $(u\cdot\sigma)(s,t)=u(s+\sigma,t)$ for every $\sigma\in\mathbb{R}$.
To define the $\mathbb{Z}/2$-graded Floer cochain complex, we refer to [19, Section 2.f]. We take the pair $(H,J)$ to be regular so that the moduli space $\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ is cut out transversely. The dimension of $\coprod_{A\in H_{1}(L_{K_{0}})}\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ is equal to $|x|-|y|-1$ modulo $2$. Moreover, by the Gromov compactness, the union of 00-dimensional components of $\coprod_{A\in H_{1}(L_{K_{0}})}\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ is a finite set. We define $n_{(H,J),A}(x,y)\in\mathbb{F}_{2}$ to be the modulo $2$ count of the number of the 00-dimensional components of $\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$.
We define $CF^{*}(\psi(\mathbb{R}^{3}),L_{K_{0}};H)$ to be the $\mathbb{Z}/2$-graded $\mathbb{F}_{2}[\lambda^{\pm}]$-module freely generated by the set $P_{H}\cap L_{K_{0}}$ whose grading is given by $|\cdot|\colon P_{H}\cap L_{K_{0}}\to\mathbb{Z}/2$. The $\mathbb{F}_{2}[\lambda^{\pm}]$-linear map $d_{J}$ of degree $1$ is defined by
$d_{J}(y)\coloneqq\sum_{m=-\infty}^{\infty}\sum_{|x|\equiv|y|+1\mod 2}\left(n_{(H,J),m[K_{0}]}(x,y)\right)\lambda^{m}\cdot x$
for every $y\in P_{H}\cap L_{K_{0}}$. By a standard argument using the Gromov compactification of $\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ for each $A\in H_{1}(L_{K_{0}})$ when its dimension is $1$, we can prove that $d_{J}\circ d_{J}=0$.
From the cochain complex, we obtain a cohomology group
$HF^{*}(\psi(\mathbb{R}^{3}),L_{K_{0}};(H,J))\coloneqq\operatorname{Ker}d_{J}/\operatorname{Im}d_{J}.$
Its isomorphism class as a $\mathbb{Z}/2$-graded $\mathbb{F}_{2}[\lambda^{\pm}]$-module is independent of the choice of the paths $\{c_{x}\mid x\in P_{H}\cap L_{K_{0}}\}$. To see this claim, we choose another path $c^{\prime}_{x}$ instead of $c_{x}$ for every $x\in P_{H}\cap L_{K_{0}}$ and define $d^{\prime}_{J}$ by using $\{c^{\prime}_{x}\mid x\in P_{H}\cap L_{K_{0}}\}$ in the same way as $d_{J}$. Take $m_{x}\in\mathbb{Z}$ such that $m_{x}[K_{0}]\in H_{1}(L_{K_{0}})$ is represented by a loop obtained by concatenating $c^{\prime}_{x}$ and $c_{x}^{-1}$. Then we can define an $\mathbb{F}_{2}$-linear isomorphism
$\begin{array}[]{cc}\Psi\colon CF^{*}(\psi(\mathbb{R}^{3}),L_{K_{0}};H)\to CF^{*}(\psi(\mathbb{R}^{3}),L_{K_{0}};H)\colon x\mapsto\lambda^{m_{x}}x&\text{for }x\in P_{H}\cap L_{K_{0}},\end{array}$
for which we have $d_{J}\circ\Psi=\Psi\circ d^{\prime}_{J}$. This proves the claim.
Moreover, by a standard argument using a homotopy from $\psi$ to $\operatorname{id}_{T^{*}\mathbb{R}^{3}}$ in $\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$, we can show the invariance of the Floer cohomology under Hamiltonian isotopies, which we state as follows.
Theorem A.1.
If we choose a pair $(H_{0},J_{0})$ as above to define $(CF^{*}(\mathbb{R}^{3},L_{K_{0}};H_{0}),d_{J_{0}})$ (i.e. the case $\psi=\operatorname{id}_{T^{*}\mathbb{R}^{3}}$), then there exists an isomorphism of $\mathbb{Z}/2$-graded $\mathbb{F}_{2}[\lambda^{\pm}]$-modules
$HF^{*}(\psi(\mathbb{R}^{3}),L_{K_{0}};(H,J))\cong HF^{*}(\mathbb{R}^{3},L_{K_{0}};(H_{0},J_{0})).$
Now, we prove Proposition 2.7 about a clean Lagrangian intersection along a circle.
Proof of Proposition 2.7.
For $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ in the statement of Proposition 2.7, let us set $\psi\coloneqq\varphi^{-1}$. Then, we can change the setup as follows: $\psi(\mathbb{R}^{3})$ and $L_{K_{0}}$ have a clean intersection along $\psi(K_{1})$, where $K_{1}$ is an oriented knot in $\mathbb{R}^{3}$. We may assume that for sufficiently small $\varepsilon>0$, $\psi(L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3})\subset L_{K_{0}}$. We fix an orientation of $\psi(K_{1})$ via $\psi\colon K_{1}\to\psi(K_{1})$ and define $a\in\mathbb{Z}$ by $[\psi(K_{1})]=a[K_{0}]\in H_{1}(L_{K_{0}})$. Let $(-1)^{\sigma}$ denote the sign of orientation of the open embedding $\psi\colon L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to L_{K_{0}}$. Then, the claim of the proposition is equivalent to that $a\in\{1,-1\}$ and $\sigma\equiv 0\mod 2$.
We choose a Mose function $f\colon\psi(K_{1})\to\mathbb{R}$ which has only two critical points $x_{0},x_{1}\in\psi(K_{1})$ such that $f$ takes its minimum at $x_{0}$ and its maximum at $x_{1}$. We also choose a Riemannian metric $g$ on $\psi(K_{1})$. By using $g$, we obtain the gradient vector field $V_{f}$ for $f$ by $g(V_{f},\cdot)=df$. Then, we consider the set $\mathcal{T}_{(f,g)}(x_{0},x_{1})$ of trajectories $\gamma\colon\mathbb{R}\to\psi(K_{1})$ such that:
  • $\frac{d\gamma}{ds}(s)=(V_{f})_{\gamma(s)}$ for every $s\in\mathbb{R}$.
  • $\lim_{s\to-\infty}\gamma(s)=x_{0}$ and $\lim_{s\to\infty}\gamma(s)=x_{1}$.
Since $\psi(\mathbb{R}^{3})$ and $L_{K_{0}}$ intersects cleanly along $\psi(K_{1})$ and $\psi(K_{1})$ is connected, we can apply [18, Theorem 3.4.11, Lemma 3.4.12] to get the following results: From the pair $(f,g)$ with $||f||_{C^{1}}$ sufficiently small, we can construct a Hamiltonian $H_{f}\colon T^{*}\mathbb{R}^{3}\to\mathbb{R}$ supported in a neighborhood of $\psi(K_{1})$ and an almost complex structure $J_{g}$ on $T^{*}\mathbb{R}^{3}$ such that the pair $(H_{f},J_{g})$ is regular and the following hold:
  • $P_{H_{f}}\cap L_{K_{0}}=\{x_{0},x_{1}\}$.
  • There is a bijection
    $\Phi\colon\coprod_{A\in H_{1}(L_{K_{0}})}\mathcal{M}_{(H_{f},J_{g}),A}(x_{0},x_{1})\to\mathcal{T}_{(f,g)}(x_{0},x_{1})\colon u\mapsto u(\cdot,1).$
It is clear that $\mathcal{T}_{(f,g)}(x_{0},x_{1})$ has only two elements. Let us label them as $\gamma_{1}$ and $\gamma_{2}$ so that $\frac{d\gamma_{2}}{ds}(s)$ is a positive vector with respect to the orientation of $\psi(K_{1})$. We tale $k\in\mathbb{Z}$ such that $\Phi^{-1}(\gamma_{1})$ belongs to $\mathcal{M}_{(H_{f},J_{g}),k[K_{0}]}(x_{0},x_{1})$. ($k$ depends on the choice of the paths $c_{x_{0}}$ and $c_{x_{1}}$.) Then, $\Phi^{-1}(\gamma_{2})$ belongs to $\mathcal{M}_{(H_{f},J_{g}),A}(x_{0},x_{1})$ for the homology class
$A=k[K_{0}]+[\psi(K_{1})]=(k+a)[K_{0}]$
since $[\psi(K_{1})]\in H_{1}(L_{K_{0}})$ is represented by a loop obtained by concatenating the two paths: $[-\infty,\infty]\to L_{K_{0}}\colon s\mapsto\gamma_{1}(-s)$ and $[-\infty,\infty]\to L_{K_{0}}\colon s\mapsto\gamma_{2}(s)$. From this observation, the differential $d_{J_{g}}\colon\mathbb{F}_{2}[\lambda^{\pm}]\cdot x_{1}\to\mathbb{F}_{2}[\lambda^{\pm}]\cdot x_{0}$ is given by
$d_{J_{g}}(x_{1})=(\lambda^{k}+\lambda^{k+a})x_{0}.$
Therefore, $HF^{*}(\psi(\mathbb{R}^{3}),L_{K_{0}};(H_{f},J_{g}))$ is computed as follow (we omit writing $(H_{f},J_{g})$):
$\begin{array}[]{lll}HF^{0+\sigma_{\psi}}(\psi(\mathbb{R}^{3}),L_{K_{0}})=0,&HF^{1+\sigma_{\psi}}(\psi(\mathbb{R}^{3}),L_{K_{0}})\cong\mathbb{F}_{2}[\lambda^{\pm}]/(1+\lambda^{a})&\text{ if }a\neq 0,\\ HF^{0+\sigma_{\psi}}(\psi(\mathbb{R}^{3}),L_{K_{0}})\cong\mathbb{F}_{2}[\lambda^{\pm}],&HF^{1+\sigma_{\psi}}(\psi(\mathbb{R}^{3}),L_{K_{0}})\cong\mathbb{F}_{2}[\lambda^{\pm}]&\text{ if }a=0.\end{array}$
Here, $\sigma_{\psi}\in\mathbb{Z}/2$ is the shift of degrees. Note that $\sigma_{\psi}$ depends only on the orientation of $L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}$ induced by the open embedding $\psi\colon L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to L_{K_{0}}$ and the fixed orientation of $L_{K_{0}}$. As $\mathbb{F}_{2}$-vector spaces,
$\begin{array}[]{lll}\dim_{\mathbb{F}_{2}}HF^{0+\sigma_{\psi}}(\psi(\mathbb{R}^{3}),L_{K_{0}})=0,&\dim_{\mathbb{F}_{2}}HF^{1+\sigma_{\psi}}(\psi(\mathbb{R}^{3}),L_{K_{0}})=|a|&\text{ if }a\neq 0,\\ \dim_{\mathbb{F}_{2}}HF^{0+\sigma_{\psi}}(\psi(\mathbb{R}^{3}),L_{K_{0}})=\infty,&\dim_{\mathbb{F}_{2}}HF^{1+\sigma_{\psi}}(\psi(\mathbb{R}^{3}),L_{K_{0}})=\infty&\text{ if }a=0.\end{array}$ (15)
If we consider a special case where $\psi$ is the identity map $\operatorname{id}=\operatorname{id}_{T^{*}\mathbb{R}^{3}}$, then $K_{1}=\psi(K_{1})=K_{0}$ and we have $a=\pm 1$. Therefore,
$\begin{array}[]{ll}\dim_{\mathbb{F}_{2}}HF^{0+\sigma_{\operatorname{id}}}(\mathbb{R}^{3},L_{K_{0}})=0,&\dim_{\mathbb{F}_{2}}HF^{1+\sigma_{\operatorname{id}}}(\mathbb{R}^{3},L_{K_{0}})=1.\end{array}$ (17)
From the definition of the orientation sign $(-1)^{\sigma}$, the shift of degrees $\sigma_{\psi}$ is equal to $\sigma_{\operatorname{id}}+\sigma$ modulo $2$. From (15), (17) and Theorem A.1, it follows that $a\in\{1,-1\}$ and $\sigma\equiv 0\mod 2$. ∎

### v2

@@
-A topological constraint on clean Lagrangian intersections from $\mathbb{Q}$-valued augmentations
+Topological constraints on clean Lagrangian intersections from $\mathbb{Q}$-valued augmentations
 Abstract
-For a knot $K$ in $\mathbb{R}^{3}$, consider its conormal bundle $L_{K}$ in $T^{*}\mathbb{R}^{3}$.
-We prove that if $K$ is the connected sum $K^{\prime}\#3_{1}$ of an arbitrary knot $K^{\prime}$ and the trefoil knot $3_{1}$ (or its mirror), then there is no compactly supported Hamiltonian diffeomorphism $\varphi$ on $T^{*}\mathbb{R}^{3}$ such that $\varphi(L_{K})$ and the zero section $\mathbb{R}^{3}$ intersect cleanly along the unknot in $\mathbb{R}^{3}$.
-For the proof, we utilize the knot DGA, whose homology was proved to be isomorphic to the knot contact homology by Ekholm, Etnyre, Ng and Sullivan, with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$.
-We discover a certain algebraic constraint imposed on any augmentation of the knot DGA of $K=K^{\prime}\#3_{1}$ to a field $\mathbb{k}$.
-This constraint is non-trivial for $\mathbb{k}=\mathbb{Q}$ and plays a key role in the proof, but trivial if $\mathbb{k}$ is quadratically closed or if an augmentation factors through the knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$ obtained by setting $U=1$.
+Let $K$ be the connected sum in $\mathbb{R}^{3}$ of an arbitrary knot and the $(2,q)$-torus knot for $q\neq\pm 1$, and consider its conormal bundle $L_{K}$ in $T^{*}\mathbb{R}^{3}$.
+We prove that there is no compactly supported Hamiltonian diffeomorphism $\varphi$ on $T^{*}\mathbb{R}^{3}$ such that $\varphi(L_{K})$ and the zero section $\mathbb{R}^{3}$ intersect cleanly along the unknot in $\mathbb{R}^{3}$.
+For the proof, we examine the knot DGA, whose homology was proved to be isomorphic to the knot contact homology by Ekholm, Etnyre, Ng and Sullivan, with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$.
+We find an algebraic constraint imposed on every $\mathbb{Q}$-valued augmentation of the knot DGA of $K$ when $q\geq 3$.
+This constraint involves rational roots of polynomials in $\mathbb{Q}[T]$, and we apply Hilbert’s irreducibility theorem to complete the proof.
+Using a similar idea, we also prove an analogous result replacing the $(2,q)$-torus knot with the figure-eight knot.
 1 Introduction
@@
 Let $\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$ denote the set of Hamiltonian diffeomorphisms on $T^{*}\mathbb{R}^{3}$ associated to compactly supported Hamiltonians on $T^{*}\mathbb{R}^{3}$.
-Using these notations, we ask a question which was observed by the author in [17].
-Question 1.1.
+Using these notations, we ask the following question, which was observed by the author in [16].
+Question 1.1 (Question 1.1 of [16]).
 Is there $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K})$ and $\mathbb{R}^{3}$ have a clean intersection along a knot in $\mathbb{R}^{3}$ whose isotopy class is different from $K$?
@@
 Indeed, for any pair of knots $(K_{0},K_{1})$ in $\mathbb{R}^{3}$, there exists a $C^{\infty}$ isotopy $(\varphi^{t})_{t\in[0,1]}$ on $T^{*}\mathbb{R}^{3}$ with compact support such that $\varphi^{1}(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along $K_{1}$.
-See [17, Proposition 4.1].
+See [16, Proposition 4.1].
 As a special case, the answer to Question 1.1 when $K$ is the unknot is ‘no’.
-This can be deduced from the results by Ganatra-Pomerleano [10, Proposition 6.29], Smith-Wemyss [21, Proposition 4.10] or Asplund-Li [1, Theorem 1.3] about clean Lagrangian intersections in the plumbings of two copies of the cotangent bundle of $S^{3}$ along the unknots.
-These results are concerned with Smith’s question on the ‘persistence (or rigidity) of unknottedness of Lagrangian intersections’ [10, Question 1.5].
-For its relation to Question 1.1, see [17, Section 1].
-In a general case, under the assumption that $\varphi(L_{K})$ for $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ and $\mathbb{R}^{3}$ intersect cleanly along a knot in $\mathbb{R}^{3}$, the author proved [17, Theorem 1.1] by using the (framed) knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$, which is a differential graded algebra (DGA) defined by Ng [14].
-[17, Theorem 1.1] is stated by using the augmentation variety in $(\mathbb{C}^{*})^{2}$ derived from the set of augmentations of the knot DGA to $\mathbb{C}$.
-This theorem leads us to the same answer to Question 1.1 as above when $K$ is the unknot [17, Theorem 1.3] and we can also deduce some constraints on the knot type of $\varphi(L_{K})\cap\mathbb{R}^{3}$ when $K$ has certain non-trivial knot types [17, Theorem 1.4, 1.5].
-However, we cannot deduce any constraint from [17, Theorem 1.1] on the knot type of $K$ if we suppose that $\varphi(L_{K})\cap\mathbb{R}^{3}$ is the unknot in $\mathbb{R}^{3}$.
-For the reason, see Remark 1.5 below.
-Main results.
-Let $3_{1}$ denote the right-handed trefoil knot in $\mathbb{R}^{3}$ and $\overline{3_{1}}$ denote its mirror.
-The goal of this paper is to prove the next theorem which cannot be deduced from [17, Theorem 1.1].
-Theorem 1.1 (Theorem 3.10).
-Let $K$ be the connected sum of an arbitrary knot in $\mathbb{R}^{3}$ with $3_{1}$ or $\overline{3_{1}}$.
+This can be deduced from the result by Ganatra-Pomerleano [9, Proposition 6.29] about clean Lagrangian intersections in a plumbing of two copies of the cotangent bundle of $S^{3}$ along the unknots (see [16, Section 1]).
+Smith-Wemyss [20, Proposition 4.10] and Asplund-Li [1, Theorem 1.3] also studies clean Lagrangian intersections in plumbings of two copies of $T^{*}S^{3}$ along the unknots.
+These results are concerned with Smith’s question on the ‘persistence (or rigidity) of unknottedness of Lagrangian intersections’ [9, Question 1.5].
+In a general case, under the assumption that $\varphi(L_{K})$ for $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ intersects $\mathbb{R}^{3}$ cleanly along a knot $K^{\prime}$ in $\mathbb{R}^{3}$, the author proved [16, Theorem 1.1].
+It relates the augmentation variety in $(\mathbb{C}^{*})^{2}$ of $K$ with that of $K^{\prime}$.
+This result leads us to the same answer to Question 1.1 as above when $K$ is the unknot [16, Theorem 1.3].
+We can also deduce some constraints on the knot type of $K^{\prime}$ when $K$ has certain non-trivial knot types [16, Theorem 1.4, 1.5].
+Main results. In fact, the aforementioned result [16, Theorem 1.1] is useless to solve the following weaker version of Question 1.1. For the reason, see Remark 1.5 below.
+Question 1.2.
+If $K$ is not the unknot, is there $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{n})$ such that $\varphi(L_{K})$ intersects $\mathbb{R}^{3}$ cleanly along the unknot in $\mathbb{R}^{3}$?
+The goal of this paper is to give a negative answer to this question when $K$ has a certain knot type.
+As a notation, let $K_{1}\#K_{2}$ denote the connected sum of knots $K_{1}$ and $K_{2}$ in $\mathbb{R}^{3}$.
+Let $T_{(2,2m+1)}$ for $m\in\mathbb{Z}$ denote the $(2,2m+1)$-torus knot and let $4_{1}$ denote the figure-eight knot.
+Theorem 1.1 (Theorem 3.15 and Theorem 3.19).
+Let $K$ be $K^{\prime}\#T_{(2,2m+1)}$ or $K^{\prime}\#4_{1}$, where $K^{\prime}$ is an arbitrary knot in $\mathbb{R}^{3}$ and $m\in\mathbb{Z}\setminus\{0,-1\}$.
 Then, there is no $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K})$ intersects $\mathbb{R}^{3}$ cleanly along the unknot in $\mathbb{R}^{3}$.
-For the proof, the geometric arguments in the present paper are similar to those in [17], which are outlined as follows: Let $K_{0}$ and $K_{1}$ be knots in $\mathbb{R}^{3}$.
+For the proof, the geometric arguments in this paper are similar to those in [16], which are outlined as follows: Let $K_{0}$ and $K_{1}$ be knots in $\mathbb{R}^{3}$.
 For $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$, suppose that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ intersect cleanly along $K_{1}$.
 We use the results by Ekholm-Etnyre-Ng-Sullivan [6, 3] which show that the Chekanov-Eliashberg DGA of the unit conormal bundle of $K_{i}$ is stable tame isomorphic to the knot DGA of $K_{i}$ for $i=0,1$.
-The knot DGA, introduced by Ng [12, 13], is defined combinatorially from a braid representation of a knot.
+The knot DGA, introduced by Ng [12], is defined combinatorially from a braid representation of a knot.
+Here, we refer to the one with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$ defined in [14, 3, 15].
 We then apply a DGA map associated to an exact Lagrangian cobordism between the unit conormal bundles of $K_{0}$ and $K_{1}$.
@@
 Consequently, we obtain a DGA map from the knot DGA of $K_{0}$ to that of $K_{1}$.
-(Precisely, we need a base change on the coefficient ring of a DGA.)
-The feature of the present paper is that we utilize the knot DGA, which we denote here by $(\mathcal{A}^{\mathrm{knot}}_{*}(K),\partial_{K})$ for an oriented knot $K$, with coefficients in the ring
-$\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]\cong\mathbb{Z}[H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K})].$
-Here, $S^{*}\mathbb{R}^{3}$ denotes the unit cotangent bundle of $\mathbb{R}^{3}$ and $\Lambda_{K}$ denotes the unit conormal bundle of $K$.
-For details, we refer to [15, 4, 16].
-We note that in [17], referring to [14, 7], the knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]\cong\mathbb{Z}[H_{1}(\Lambda_{K})]$ was used.
-To prove Theorem 1.1, we show the following algebraic property of the knot DGA under an assumption on the isotopy class of $K$.
-Let $K^{\prime}\#3_{1}$ denote the connected sum of a knot $K^{\prime}$ in $\mathbb{R}^{3}$ and the trefoil knot $3_{1}$.
-Lemma 1.2 (Lemma 3.9).
-Let $\mathbb{k}$ be a field and $K^{\prime}$ be a knot.
-Then, for any augmentation $\varepsilon\colon\mathcal{A}_{*}^{\mathrm{knot}}(K^{\prime}\#3_{1})\to\mathbb{k}$ of the knot DGA of $K^{\prime}\#3_{1}$ to $\mathbb{k}$,
-$D_{\varepsilon}\coloneqq(\varepsilon(\mu)+1-\varepsilon(\mu)^{-1}\varepsilon(U))^{2}-4(\varepsilon(\mu)-1)$
-satisfies $D_{\varepsilon}\in\{k^{2}\mid k\in\mathbb{k}\}$.
-Here, $\varepsilon(\mu)$ and $\varepsilon(U)$ are abbreviations of $\varepsilon(\mu\cdot\mathbb{1})$ and $\varepsilon(U\cdot\mathbb{1})$ for the unit $\mathbb{1}$ of the knot DGA.
-In fact, $D_{\varepsilon}$ is the discriminant of a quadratic equation in $\mathbb{k}$ determined by $\varepsilon$ (see the proof of Lemma 3.9).
-Note that if $\varepsilon$ factors through the knot DGA over $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$ obtained by setting $U=1$, then $\varepsilon(U)=1$ and we have an identity
-$(\varepsilon(\mu)+1-\varepsilon(\mu)^{-1})^{2}-4(\varepsilon(\mu)-1)=(\varepsilon(\mu)-1+\varepsilon(\mu)^{-1})^{2},$
-so the constraint on $D_{\varepsilon}$ is trivial.
-For the proof of Theorem 1.1, we use the constraint from Lemma 1.2 in the case of $\mathbb{k}=\mathbb{Q}$, that is, $D_{\varepsilon}$ is the square of some rational number for any augmentation $\varepsilon$ of the knot DGA of $K^{\prime}\#3_{1}$ to $\mathbb{Q}$.
-(In fact, for this purpose, we may choose other fields such as $\mathbb{Q}[\sqrt{-1}]$ and $\mathbb{F}_{5}$.
-See Remark 3.11.) It is crucial that $\mathbb{Q}$ is not quadratically (and algebraically) closed, and we remark that the constraint on $D_{\varepsilon}$ is trivial for any $\varepsilon$ if $\mathbb{k}=\mathbb{C}$.
+From a simple description of the knot DGA of the unknot, we obtain the following result about augmentations of the knot DGA of $K_{0}$.
+Proposition 1.2 (Proposition 3.5).
+Let $\mathbb{k}$ be a field.
+Suppose that $K_{1}$ is the unknot.
+Then, for every $y,Z\in\mathbb{k}^{*}$ satisfying $y\notin\{Z^{n}\mid n\in\mathbb{Z}\}$, there exists an augmentation $\varepsilon$ of the knot DGA of $K_{0}$ to $\mathbb{k}$ such that $\varepsilon(\mu)=y$ and $\varepsilon(U)=Z$.
+There are two new points in algebraic arguments in this paper. Let us explain them by focusing on the knot $K^{\prime}\#T_{(2,2m+1)}$ for $m\in\mathbb{Z}_{\geq 1}$.
+The first point is that we find a constraint on augmentations of some knot DGAs in terms of roots of polynomials in $\mathbb{k}[T]$, where $\mathbb{k}$ is any field.
+Actually, we find a specific polynomial $P_{m}\in\mathbb{Z}[\mu,U][T]$ for every $m\in\mathbb{Z}_{\geq 1}$, whose degree in $T$ is $m+1$, with the following property (see Proposition 3.10): For any augmentation $\varepsilon$ of the knot DGA of $K^{\prime}\#T_{(2,2m+1)}$ to $\mathbb{k}$ such that $\varepsilon(\mu)=y$ and $\varepsilon(U)=Z$, the polynomial $\left.P_{m}\right|_{\mu=y,U=Z}\in\mathbb{k}[T]$ has a root in $\mathbb{k}$.
+About the derivation of $P_{m}$, see Remark 1.3 below.
+We will focus on the case $\mathbb{k}=\mathbb{Q}$.
+It is worth noting that this constraint on augmentations is trivial if $\mathbb{k}$ is algebraically closed.
+The second point is that we use a classical result in number theory.
+Indeed, Theorem 1.1 when $K=K^{\prime}\#T_{(2,2m+1)}$ for $m\in\mathbb{Z}_{\geq 1}$ is proved by a contradiction as follows: Suppose that $K_{0}=K^{\prime}\#T_{(2,2m+1)}$ and $K_{1}$ is the unknot.
+Then, Proposition 1.2 and the property of $P_{m}$ for $\mathbb{k}=\mathbb{Q}$ imply that $\left.P_{m}\right|_{\mu=2,U=Z}\in\mathbb{Q}[T]$ has a root in $\mathbb{Q}$ for every $Z\in\mathbb{Q}^{*}\setminus\{2^{\pm 1}\}$.
+On the other hand, by using Hilbert’s irreducibility theorem, we can show that there are infinitely many $Z\in\mathbb{Q}$ such that $\left.P_{m}\right|_{\mu=2,U=Z}\in\mathbb{Q}[T]$ does not have a root in $\mathbb{Q}$, which is a contradiction.
+(From the case of $m\geq 1$, we can deduce Theorem 1.1 when $K=K^{\prime}\#T_{(2,-2m-1)}$.)
+The proof of Theorem 1.1 when $K=K^{\prime}\#4_{1}$ goes in a similar way.
+We explicitly give a polynomial $P\in\mathbb{Z}[\mu^{\pm},U^{\pm}][T]$, whose degree in $T$ is $3$, with the following property (see Proposition 3.17): For any augmentation $\varepsilon$ of the knot DGA of $K^{\prime}\#4_{1}$ to $\mathbb{k}$ such that $\varepsilon(\mu)=y$ and $\varepsilon(U)=Z$, the polynomial $\left.P\right|_{\mu=y,U=Z}\in\mathbb{k}[T]$ has a root in $\mathbb{k}$.
+Using concrete calculations and the rational root theorem, we show that $\left.P\right|_{\mu=-1,U=-2}\in\mathbb{Q}[T]$ does not have a root in $\mathbb{Q}$.
+Combining with Proposition 1.2 for $\mathbb{k}=\mathbb{Q}$, Theorem 1.1 when $K=K^{\prime}\#4_{1}$ is deduced.
 Remark 1.3.
-In order to study the knot DGA of a knot $K$ (or the Chekanov-Eliashberg DGA of $\Lambda_{K}$), the set of augmentations to $\mathbb{k}$ has been used with $\mathbb{k}=\mathbb{C}$ or a finite field $\mathbb{F}_{p}$ for a prime number $p$.
-For examples of applications, see Ng [14, Section 5.2 and 5.3] and Cornwell [2] about the augmentation polynomial when $\mathbb{k}=\mathbb{C}$ and [14, Section 5.1] about the augmentation number when $\mathbb{k}=\mathbb{F}_{p}$.
-On the other hand, the rational numbers $D_{\varepsilon}$ from Lemma 1.2 when $\mathbb{k}=\mathbb{Q}$ are derived in a way different from those previous researches.
-Lastly, we provide a theorem which is a refinement of [17, Theorem 1.1].
-For an oriented knot $K$ in $\mathbb{R}^{3}$, let $V_{K}$ denote the augmentation variety defined by [16, Definition 5.1] as a subset of $(\mathbb{C}^{*})^{3}$.
+The polynomials $P_{m}$ for $K^{\prime}\#T_{(2,2m+1)}$ and $P$ for $K^{\prime}\#4_{1}$ are found from concrete computations on some part of (the homology of) the knot DGA which comes from the component $T_{(2,2m+1)}$ or $4_{1}$ of the connected sum.
+Our computation relies on braid representations of $T_{(2,2m+1)}$ and $4_{1}$, which are rather simple.
+It might be possible to find a polynomial with a similar property as $P_{m}$ for more complicated knots, such as the $(p,q)$-torus knot when $3\leq p<|q|$.
+We also provide a theorem which is a refinement of [16, Theorem 1.1].
+For an oriented knot $K$ in $\mathbb{R}^{3}$, let $V_{K}$ denote the augmentation variety defined by [15, Definition 5.1] as a subset of $(\mathbb{C}^{*})^{3}$.
 Let $(x,y,Z)$ be the coordinate of $\mathbb{C}^{3}$.
@@
 For oriented knots $K_{0}$ and $K_{1}$ in $\mathbb{R}^{3}$, suppose that there exists $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ intersects $\mathbb{R}^{3}$ cleanly along $K_{1}$.
-Then, there exist integers $b,c$ and $d$ for which we have an injective map
-$V_{K_{1}}\hookrightarrow V_{K_{0}}\colon(x,y,Z)\mapsto(xy^{b}Z^{c},yZ^{d},Z).$
+Then, there exist $a\in\{\pm 1\}$ and integers $b,c$ and $d$ for which we have an injective map
+$V_{K_{1}}\hookrightarrow V_{K_{0}}\colon(x,y,Z)\mapsto(x^{a}y^{b}Z^{c},y^{a}Z^{d},Z).$
 Remark 1.5.
-Identifying $\{(x,y,Z)\in(\mathbb{C}^{*})^{3}\mid Z=1\}$ with $(\mathbb{C}^{*})^{2}$, the augmentation variety considered in [17] coincides with $V_{K}\cap\{Z=1\}\subset(\mathbb{C}^{*})^{2}$.
+Identifying $\{(x,y,Z)\in(\mathbb{C}^{*})^{3}\mid Z=1\}$ with $(\mathbb{C}^{*})^{2}$, we obtain $V_{K}\cap\{Z=1\}\subset(\mathbb{C}^{*})^{2}$, which is a version of augmentation variety used in [16].
+This corresponds to the reduction of the coefficient ring of the knot DGAs from $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$ to $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$ by setting $U=1$.
 It follows from Theorem 1.4 that there exists an injective map
-$V_{K_{1}}\cap\{Z=1\}\hookrightarrow V_{K_{0}}\cap\{Z=1\}\colon(x,y)\mapsto(xy^{b},y)$
-for some $b\in\mathbb{Z}$.
-This can be seen as a slight refinement of [17, Theorem 1.1], but it gives no constraint on $K_{0}$ if we suppose that $K_{1}$ is the unknot.
-The reason is that $V_{K_{1}}\cap\{Z=1\}=\{x=1\}\cup\{y=1\}$ if $K_{1}$ is the unknot (see [16, Example 3.13]) and $\{x=1\}\cup\{y=1\}$ is a subset of $V_{K_{0}}\cap\{Z=1\}$ for any knot $K_{0}$ (see [14, Proposition 5.6] and [16, Exercise 3.17]).
+$V_{K_{1}}\cap\{Z=1\}\hookrightarrow V_{K_{0}}\cap\{Z=1\}\colon(x,y)\mapsto(x^{a}y^{b},y^{a})$
+for some $a\in\{\pm 1\}$ and $b\in\mathbb{Z}$.
+This is a slight refinement of [16, Theorem 1.1], but it gives no constraint on $K_{0}$ if we suppose that $K_{1}$ is the unknot.
+The reason is that $V_{K_{1}}\cap\{Z=1\}=\{x=1\}\cup\{y=1\}$ if $K_{1}$ is the unknot (see [15, Example 3.13]) and $\{x=1\}\cup\{y=1\}$ is a subset of $V_{K_{0}}\cap\{Z=1\}$ for every knot $K_{0}$ (see [13, Proposition 5.6] and [15, Exercise 3.17]).
 Notations.
@@
 In Section 2, for a contact manifold $Y$ and its Legendrian submanifold $\Lambda$ satisfying certain conditions, we introduce the Chekanov-Eliashberg DGA of $(Y,\Lambda)$ with coefficients in $\mathbb{Z}[H_{2}(Y,\Lambda)]$.
-We then define a DGA map associated to an exact Lagrangian cobordism in ×$\mathbb{R}\times Y$.
-In Section 2.5, we will see how to apply this DGA map to a setting with a clean Lagrangian intersection as in Question 1.1.
+We then define a DGA map associated to an exact Lagrangian cobordism in $\mathbb{R}\times Y$.
+In Section 2.5, we will see how to apply this DGA map to a setting with clean Lagrangian intersection as in Question 1.1.
 In Section 3, we combine the argument in Section 2.5 with the results of [3, 15] about the knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$.
-Theorem 1.4 is proved in Section 3.1.
-Theorem 1.1 is proved by using Lemma 1.2 (Lemma 3.9) in Section 3.2.
-In Appendix A, by using Floer cohomology group over $\mathbb{F}_{2}[\lambda^{\pm}]$, we prove Proposition 2.7 which gives a homological constraint on $1$-dimensional clean intersections.
-This enhances some results in [17].
-Nevertheless, the present proof is much simpler than that in [17].
+Proposition 1.2 and Theorem 1.4 are proved in Section 3.1.
+Theorem 1.1 for $K=K^{\prime}\#T_{(2,2m+1)}$ with $m\neq 0,-1$ (resp.
+$K=K^{\prime}\#4_{1}$) is proved by examining the knot DGA of $K^{\prime}\#T_{(2,2m+1)}$ in Section 3.2 (resp.
+$K^{\prime}\#4_{1}$ in Section 3.3).
+In Appendix A, by using Lagrangian Floer cohomology group over $\mathbb{F}_{2}[\lambda^{\pm}]$, we prove Proposition 2.7 which gives a homological constraint on $1$-dimensional clean intersections.
+This strengthen a result in [16].
+Nevertheless, the present proof is much simpler than that of [16].
 Acknowledgements.
 The author would like to thank Kei Irie for making valuable comments.
-He would also like to thank Fumikazu Nagasato and Kouki Sato for fruitful discussion before this work and for giving some insight from knot theory.
+He would also like to thank Fumikazu Nagasato and Kouki Sato for the discussion before this work and giving some insight from knot theory.
+He would also like to thank Lenhard Ng for the discussion during CBMS summer school 2025 and suggestions for improving the main result.
+The author appreciates the referee’s valuable comments.
 This work was supported by JSPS KAKENHI Grant Number JP23KJ1238.
+The author is currently supported by JSPS KAKENHI Grant Number JP25KJ0270.
 2 Symplectic field theory for exact Lagrangian cobordisms
-We fix a $2n$-dimensional simply connected manifold $P$ with a Liouville $1$-form $\lambda$ such that $P$ is the completion of a compact Liouville domain $(\bar{P},\left.\lambda\right|_{\bar{P}})$ satisfying $c_{1}(T\bar{P})=0$.
+We fix a $2n$-dimensional simply connected manifold $P$ with a Liouville $1$-form $\lambda$ such that $P$ is the completion of a compact Liouville domain $(\bar{P},\left.\lambda\right|_{\bar{P}})$ satisfying $2c_{1}(T\bar{P})=0$.
 Let us consider
@@
 Let $\pi_{P}\colon Y\to P$ denote the projection to $P$.
-The arguments from Section 2.1 to Section 2.4 are completely parallel to those in [17, Section 3], except the difference of the coefficient rings.
+The arguments from Section 2.1 to Section 2.4 are parallel to those in [16, Section 3], except the difference of the coefficient rings.
 In Section 2.5, we will discuss a particular case related to clean Lagrangian intersections.
 2.1 Preliminaries
-We fix a complex structure $j_{0}$ on ×$\mathbb{R}\times[0,1]$ such that $(j_{0})_{(s,t)}(\partial_{s})=\partial_{t}$, where $(s,t)$ is the coordinate of ×$\mathbb{R}\times[0,1]$.
+We fix a complex structure $j_{0}$ on $\mathbb{R}\times[0,1]$ such that $(j_{0})_{(s,t)}(\partial_{s})=\partial_{t}$, where $(s,t)$ is the coordinate of $\mathbb{R}\times[0,1]$.
 Let $D\coloneqq\{z\in\mathbb{C}\mid|z|\leq 1\}$.
@@
 with respect to $j_{0}$ and $j_{\kappa}$ such that $\lim_{s\to\infty}\psi_{0}(s,t)=p_{0}$ and $\lim_{s\to-\infty}\psi_{k}(s,t)=p_{k}$ uniformly on $t\in[0,1]$.
-Let $\overline{D_{m+1}}$ be a compactification of $D_{m+1}$ obtained by gluing $[0,\infty]\times[0,1]$ via $\psi_{0}$ and $[-\infty,0]\times[0,1]$ via $\psi_{k}$ for every $k=1,\dots,m$.
-In a natural way, $\overline{D_{m+1}}$ has a structure of a compact oriented surface with boundary and corner.
-Let $\Lambda$ be a compact Legendrian submanifold of $(Y,\alpha)$. Later, we will impose the following conditions on $\Lambda$:
-  1. 1.
-    $\Lambda$ is a connected and oriented manifold and admits a spin structure.
-  2. 2.
-    For the immersion $\left.\pi_{P}\right|_{\Lambda}\colon\Lambda\to P$, the self-intersection of $\pi_{P}(\Lambda)$ consists only of transversal double points.
-  3. 3.
-    For the immersed Lagrangian submanifold $\pi_{P}(\Lambda)$ of $P$, its Maslov class vanishes.
-The second condition is satisfied for generic $\Lambda$. We denote the set of Reeb chords of $\Lambda$ by
+Let $\overline{D_{m+1}}$ be a compactification of $D_{m+1}$ obtained by gluing $0,\infty]\times[0,1]$ via $\psi_{0}$ and $\infty,0]\times[0,1]$ via $\psi_{k}$ for every $k=1,\dots,m$.
+In a natural way, $\overline{D_{m+1}}$ has a structure of a compact oriented surface with boundary and corners.
+Let $\Lambda$ be a compact Legendrian submanifold of $(Y,\alpha)$. We denote the set of Reeb chords of $\Lambda$ by
 ${\mathcal{R}(\Lambda)\coloneqq\{a\colon[0,T]\to Y\mid T>0,\ \frac{da}{dt}(t)=(\partial_{z})_{a(t)}\text{ for every }t\in[0,T]\text{ and }a(0),a(T)\in\Lambda\}.}$
@@
 for every $t\in[0,1]$.
-When the second condition above is satisfied, the Conley-Zehnder index $\mu(a)\in\mathbb{Z}$ of $a$ is defined by using the capping path $\left.w_{a}\right|_{\partial D_{1}}$ for $a$ as described in [17, Section 3.1].
-We then define
-$|a|\coloneqq\mu(a)-1.$
 2.2 Moduli space of pseudo-holomorphic curves with boundary in an exact Lagrangian cobordism
 Let $\Lambda_{+}$ and $\Lambda_{-}$ be two compact Legendrian submanifolds of $(Y,\alpha)$.
-Recall that a submanifold $L$ of ×$\mathbb{R}\times Y$ is called an exact Lagrangian cobordism from $\Lambda_{-}$ to $\Lambda_{+}$ if the following are satisfied:
+Recall that a submanifold $L$ of $\mathbb{R}\times Y$ is called an exact Lagrangian cobordism from $\Lambda_{-}$ to $\Lambda_{+}$ if the following hold:
   • There exist $r_{+},r_{-}\in\mathbb{R}$ with $r_{-}<r_{+}$ such that
@@
     and $L\cap([r_{-},r_{+}]\times Y)$ is compact.
-  • There exists a $C^{\infty}$ function $f\colon L\to\mathbb{R}$ such that $df=\left.(e^{r}\alpha)\right|_{L}$ and
+  • There exists a $C^{\infty}$ function $f\colon L\to\mathbb{R}$ such that $df=\left.(e^{r}\alpha)\right|_{L}$ and its restriction
     $\left.f\right|_{(-\infty,r_{-}]\times\Lambda_{-}}\colon(-\infty,r_{-}]\times\Lambda_{-}\to\mathbb{R}$
-    is a constant function. Here, $r$ is the $\mathbb{R}$-coordinate of ×$\mathbb{R}\times Y$.
+    is a constant function. Here, $r$ is the $\mathbb{R}$-coordinate of $\mathbb{R}\times Y$.
 In the rest of this paper, we only consider the case where $\Lambda_{-}$ is connected.
@@
 $\tau_{s}\colon\mathbb{R}\times Y\to\mathbb{R}\times Y\colon(r,x)\mapsto(r+s,x).$
-Let $J$ be an almost complex structure on $P$ such that $d\lambda(\cdot,J\cdot)$ is a Riemannian metric on $P$.
-We associate to $J$ an almost complex structure $\widetilde{J}$ on ×$\mathbb{R}\times Y$ determined by $\tau_{s}^{*}\widetilde{J}=\widetilde{J}$ for every $s\in\mathbb{R}$, $\widetilde{J}(\operatorname{Ker}\alpha)=\operatorname{Ker}\alpha$$(\pi_{P})_{*}(\widetilde{J}(v))=J((\pi_{P})_{*}(v))$ for every $v\in\operatorname{Ker}\alpha$ and $\widetilde{J}(\partial_{z})=\partial_{r}$.
+Let $J$ be an almost complex structure on $P$ such that $d\lambda(\cdot,J\cdot)$ is a Riemannian metric on $P$ and adapted to $\Lambda$ in the sense of the definition in [5, Section 2.3].
+We associate to $J$ an almost complex structure $\widetilde{J}$ on $\mathbb{R}\times Y$ determined by $\tau_{s}^{*}\widetilde{J}=\widetilde{J}$ for every $s\in\mathbb{R}$, $\widetilde{J}(\operatorname{Ker}\alpha)=\operatorname{Ker}\alpha$, $(\pi_{P})_{*}(\widetilde{J}(v))=J((\pi_{P})_{*}(v))$ for every $v\in\operatorname{Ker}\alpha$ and $\widetilde{J}(\partial_{z})=\partial_{r}$.
 For any $a\in\mathcal{R}(\Lambda_{+})$, $b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda_{-})$ and $C\in H_{2}(\mathbb{R}\times Y,L)$, we define
-^$\hat{\mathcal{M}}_{L,C}(a;b_{1},\dots,b_{m})$
-to be the space of pairs $(u,\kappa)$ of $\kappa\in\mathcal{C}_{m+1}$ and a $C^{\infty}$ map $u\colon D_{m+1}\to P$ such that:
+$\hat{\mathcal{M}}_{L,C}(a;b_{1},\dots,b_{m})$
+to be the space of pairs $(u,\kappa)$ of $\kappa\in\mathcal{C}_{m+1}$ and a $C^{\infty}$ map $u\colon D_{m+1}\to\mathbb{R}\times Y$ such that:
   • $du+\widetilde{J}\circ du\circ j_{\kappa}=0$.
@@
     represents the homology class $C$. (More precisely, each term is regarded as a singular 2-chain after taking suitable triangulations of $\overline{D_{m+1}}$ and $\overline{D_{1}}$.)
-When $m\leq 1$, $\mathcal{C}_{m+1}$ consists of a single element and the group $\mathcal{A}_{m+1}$ of conformal automorphisms on $D_{m+1}$ acts on ^$\hat{\mathcal{M}}_{L,C}(a;b_{1},\dots,b_{m})$ by $(u,\kappa)\cdot\varphi\coloneqq(u\circ\varphi,\kappa)$ for every $\varphi\in\mathcal{A}_{m+1}$.
-As in [17, (10) in Section 3.2], let us define a moduli space to be
+When $m\leq 1$, $\mathcal{C}_{m+1}$ consists of a single element and the group $\mathcal{A}_{m+1}$ of conformal automorphisms on $D_{m+1}$ acts on $\hat{\mathcal{M}}_{L,C}(a;b_{1},\dots,b_{m})$ by $(u,\kappa)\cdot\varphi\coloneqq(u\circ\varphi,\kappa)$ for every $\varphi\in\mathcal{A}_{m+1}$.
+Let us define a moduli space to be
 $\mathcal{M}_{L,C}(a;b_{1},\dots,b_{m})\coloneqq\begin{cases}\hat{\mathcal{M}}_{L,C}(a;b_{1},\dots,b_{m})/\mathcal{A}_{m+1}&\text{ if }m=0,1,\\ \hat{\mathcal{M}}_{L,C}(a;b_{1},\dots,b_{m})&\text{ if }m\geq 2.\end{cases}$
-In the case where $\Lambda_{+}=\Lambda_{-}=\Lambda$ and $L=\mathbb{R}\times\Lambda$, for any $a,b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda)$ and $A\in H_{2}(Y,\Lambda)$, an $\mathbb{R}$-action on ×$\mathcal{M}_{\mathbb{R}\times\Lambda,(i_{+})_{*}(A)}(a;b_{1},\dots,b_{m})$ is defined by $s\cdot(u,\kappa)\coloneqq(\tau_{s}\circ u,\kappa)$ for every $s\in\mathbb{R}$.
+In the case where $\Lambda_{+}=\Lambda_{-}=\Lambda$ and $L=\mathbb{R}\times\Lambda$, for any $a,b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda)$ and $A\in H_{2}(Y,\Lambda)$, an $\mathbb{R}$-action on $\mathcal{M}_{\mathbb{R}\times\Lambda,(i_{+})_{*}(A)}(a;b_{1},\dots,b_{m})$ is defined by $s\cdot(u,\kappa)\coloneqq(\tau_{s}\circ u,\kappa)$ for every $s\in\mathbb{R}$.
 Let us denote the quotient space by
-¯$\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})\coloneqq\mathbb{R}\backslash\mathcal{M}_{\mathbb{R}\times\Lambda,(i_{+})_{*}(A)}(a;b_{1},\dots,b_{m}).$
+$\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})\coloneqq\mathbb{R}\backslash\mathcal{M}_{\mathbb{R}\times\Lambda,(i_{+})_{*}(A)}(a;b_{1},\dots,b_{m}).$
 2.3 Chekanov-Eliashberg DGA over $\mathbb{Z}[H_{2}(Y,\Lambda)]$
-We consider a compact Legendrian submanifold $\Lambda$ of $(Y,\alpha)$ satisfying the three conditions in Section 2.1.
-We also assume that $\Lambda$ is admissible in the sense of the definition in [6, Section 2.3].
-Let us take a spin structure $\mathfrak{s}$ on $\Lambda$.
-We take an almost complex structure $J$ on $P$ as in [6, Proposition 2.3] such that the following hold for every $a,b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda)$ and $A\in H_{2}(Y,\Lambda)$: The moduli space ¯$\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ is cut out transversely and it is a manifold of dimension $|a|-(|b_{1}|+\dots+|b_{m}|)-1$.
-Moreover, when $|b_{1}|+\dots+|b_{m}|=|a|-1$, $\coprod_{A\in H_{2}(Y,\Lambda)}\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ is a compact 00-dimensional manifold.
-(For the relation between $J$-holomorphic disks in $P$ defined as in [6, Section 2.3] and $\widetilde{J}$-holomorphic maps in ×$\mathbb{R}\times Y$ defined as in Section 2.2, see [3, Theorem 2.1] and [17, Appendix A.4].)
-In addition, ¯$\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ has an orientation depending on $\mathfrak{s}$ and a choice of an orientation of the determinant line of the capping operator for each Reeb chord of $\Lambda$.
-See [17, Definition A.4] for details which follow the conventions in [5] related to orientations.
-Let $\mathcal{A}_{*}(\Lambda)$ be the graded $\mathbb{Z}[H_{2}(Y,\Lambda)]$-algebra freely generated by $\mathcal{R}(\Lambda)$ whose graded is given by $|\cdot|\colon\mathcal{R}(\Lambda)\to\mathbb{Z}$.
-A derivative over $\mathbb{Z}[H_{2}(Y,\Lambda)]$ of degree $-1$
+We consider a compact Legendrian submanifold $\Lambda$ of $(Y,\alpha)$ satisfying the following conditions:
+  1. 1.
+    $\Lambda$ is a connected and oriented manifold and admits a spin structure.
+  2. 2.
+    For the immersion $\left.\pi_{P}\right|_{\Lambda}\colon\Lambda\to P$, the self-intersection of $\pi_{P}(\Lambda)$ consists only of transverse double points.
+  3. 3.
+    For the immersed Lagrangian submanifold $\pi_{P}(\Lambda)$ of $P$, its Maslov class vanishes.
+The second condition is satisfied for generic $\Lambda$.
+Under the second condition, the Conley-Zehnder index $\mu(a)\in\mathbb{Z}$ of $a$ is defined by using the capping path $\left.w_{a}\right|_{\partial D_{1}}$ for $a$ as in [16, Section 3.1].
+We then define
+$|a|\coloneqq\mu(a)-1.$
+We also assume that $\Lambda$ is admissible in the sense of the definition in [5, Section 2.3]. Let us fix a spin structure $\mathfrak{s}$ on $\Lambda$.
+We take an almost complex structure $J$ on $P$ as in [5, Proposition 2.3] such that the following hold for every $a,b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda)$ and $A\in H_{2}(Y,\Lambda)$ such that $(b_{1},\dots,b_{m})\neq(a)$: The moduli space $\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ is cut out transversely and it is a manifold of dimension $|a|-(|b_{1}|+\dots+|b_{m}|)-1$.
+Moreover, when $|b_{1}|+\dots+|b_{m}|=|a|-1$, $\coprod_{A\in H_{2}(Y,\Lambda)}\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ is a compact 0-dimensional manifold.
+(For the relation between $J$-holomorphic disks in $P$ defined as in [5, Section 2.3] and $\widetilde{J}-$holomorphic maps in $\mathbb{R}\times Y$ defined as in Section 2.2, see [2, Theorem 2.1] and [16, Appendix A.4].)
+In addition, $\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ has an orientation depending on $\mathfrak{s}$ and a choice of an orientation of the determinant line of the capping operator for each Reeb chord of $\Lambda$.
+See [16, Definition A.4] for details which follow the conventions in [4] related to orientations.
+Let $\mathcal{A}_{*}(\Lambda)$ be the graded $\mathbb{Z}[H_{2}(Y,\Lambda)]$-algebra freely generated by $\mathcal{R}(\Lambda)$ whose grading is given by $|\cdot|\colon\mathcal{R}(\Lambda)\to\mathbb{Z}$.
+A derivation over $\mathbb{Z}[H_{2}(Y,\Lambda)]$ of degree 1-1
 $\partial_{J,\mathfrak{s}}\colon\mathcal{A}_{*}(\Lambda)\to\mathcal{A}_{*-1}(\Lambda)$
@@
 $\partial_{J,\mathfrak{s}}(a)\coloneqq\sum_{|b_{1}|+\dots+|b_{m}|=|a|-1}\sum_{A\in H_{2}(Y,\Lambda)}(-1)^{(n-1)(|a|+1)}\left(\#_{\operatorname{sign}}\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})\right)e^{A}\cdot b_{1}\cdots b_{m}$
-and extended by the Leibniz rule. Here, #$\#_{\operatorname{sign}}$ is the number of the points counted with the signs induced by the orientation.
-We refer to [5, Theorem 4.1, Theorem 4.12] which takes $\mathbb{Z}[H_{1}(\Lambda)]$ as the coefficient ring of the Chekanov-Eliashberg DGA.
-We modify the arguments in their proofs to the present case where we take $\mathbb{Z}[H_{2}(Y,\Lambda)]$ as the coefficient ring.
-Then, we have the following theorem.
+and extended by the Leibniz rule. Here, $_{\operatorname{sign}}$ is the number of the points counted with the signs induced by the orientation.
+We refer to [4, Theorem 4.1, Theorem 4.12] which takes $\mathbb{Z}[H_{1}(\Lambda)]$ as the coefficient ring of the Chekanov-Eliashberg DGA.
+We modify the arguments in their proof to the present case where we take $\mathbb{Z}[H_{2}(Y,\Lambda)]$ as the coefficient ring.
+Then, we obtain the following theorem.
 Theorem 2.1.
@@
 Let us discuss the effect by changing the spin structure on $\Lambda$.
-For any two spin structures $\mathfrak{s}_{0}$ and $\mathfrak{s}_{1}$ on $\Lambda$, a cohomology class $d(\mathfrak{s}_{0},\mathfrak{s}_{1})\in H^{1}(\Lambda;\mathbb{Z}/2)$ is defined as in [5, Subsection 4.4.1].
+For any two spin structures $\mathfrak{s}_{0}$ and $\mathfrak{s}_{1}$ on $\Lambda$, a cohomology class $d(\mathfrak{s}_{0},\mathfrak{s}_{1})\in H^{1}(\Lambda;\mathbb{Z}/2)$ is defined as in [4, Subsection 4.4.1].
 Let $\partial\colon H_{2}(Y,\Lambda)\to H_{1}(\Lambda)$ denote the boundary map.
@@
 $\coprod_{|b_{1}|+\dots+|b_{m}|=|a|-1}\coprod_{A\in H_{2}(Y,\Lambda)\colon\partial A=\beta}\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$
-is changed by the sign $(-1)^{\langle d(\mathfrak{s},\mathfrak{s}^{\prime}),\beta\rangle}$. See the proof of [5, Theorem 4.29]. We define a ring isomorphism
+is changed by the sign $(-1)^{\langle d(\mathfrak{s},\mathfrak{s}^{\prime}),\beta\rangle}$. See the proof of [4, Theorem 4.29]. We define a ring isomorphism
 $\psi(\mathfrak{s},\mathfrak{s}^{\prime})\colon\mathbb{Z}[H_{2}(Y,\Lambda)]\to\mathbb{Z}[H_{2}(Y,\Lambda)]\colon e^{A}\mapsto(-1)^{\langle d(\mathfrak{s},\mathfrak{s}),\partial A\rangle}e^{A}.$ (2)
-Then, the next proposition easily follows as in the proof of [5, Theorem 4.30].
+Then, the next proposition holds as in the proof of [4, Theorem 4.30].
 Proposition 2.2.
@@
 We assume that $L$ is oriented and has a spin structure $\mathfrak{s}_{L}$.
-Then, on a boundary $\Lambda_{+}$ (resp.
-$\Lambda_{-}$) of $L$, an orientation and a spin structure $\mathfrak{s}_{+}$ (resp.
+Then, on positive end $\Lambda_{+}$ (resp.
+the negative end $\Lambda_{-}$) of $L$, an orientation and a spin structure $\mathfrak{s}_{+}$ (resp.
 $\mathfrak{s}_{-}$) are induced.
-See [17, Remark 3.6] for their precise descriptions.
+See [16, Remark 3.6] for their precise descriptions.
 Associated to $(\Lambda_{+},\mathfrak{s}_{+})$ and $(\Lambda_{-},\mathfrak{s}_{-})$, the Chekanov-Eliashberg DGAs
@@
 and we regard it as a DGA over $\mathbb{Z}[H_{2}(Y,\Lambda_{+})]$ via $(i_{+})_{*}$.
-By [8, Lemma 3.8], the moduli space $\mathcal{M}_{L,C}(a;b_{1},\dots,b_{m})$ is cut out transversely for generic $J$ and it is a manifold of dimension $|a|-(|b_{1}|+\dots+|b_{m}|)$ for any $a\in\mathcal{R}(\Lambda_{+})$, $b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda_{-})$ and $C\in H_{2}(\mathbb{R}\times Y,L)$.
-By the Gromov compactness, if $|b_{1}|+\dots+|b_{m}|=|a|$, $\coprod_{C\in H_{2}(\mathbb{R}\times Y,L)}\mathcal{M}_{L,C}(a;b_{1},\dots,b_{m})$ is a compact 00-dimensional manifold.
+By [7, Lemma 3.8], the moduli space $\mathcal{M}_{L,C}(a;b_{1},\dots,b_{m})$ is cut out transversely for generic $J$ and it is a manifold of dimension $|a|-(|b_{1}|+\dots+|b_{m}|)$ for any $a\in\mathcal{R}(\Lambda_{+})$, $b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda_{-})$ and $C\in H_{2}(\mathbb{R}\times Y,L)$.
+By the Gromov compactness, if $|b_{1}|+\dots+|b_{m}|=|a|$, $\coprod_{C\in H_{2}(\mathbb{R}\times Y,L)}\mathcal{M}_{L,C}(a;b_{1},\dots,b_{m})$ is a compact 0-dimensional manifold.
 In addition, it has an orientation depending on $\mathfrak{s}_{L}$ and a choice of an orientation of the determinant line of the capping operator for each Reeb chord of $\Lambda_{+}$ and $\Lambda_{-}$.
-See [17, Section A.3].
-We refer to the proof of [17, Theorem 3.7] which takes $\mathbb{Z}[H_{1}(\Lambda_{\pm})]$ as the coefficient rings.
-We modify the arguments in [17, Appendix A.6] to the present case where we take $\mathbb{Z}[H_{2}(Y,\Lambda_{\pm})]$ as the coefficient rings.
-(As referred in [17, Appendix A], see also [11, Theorem 2.5] about the DGA map over $\mathbb{Z}$, but note that some conventions in [11] related to orientations are different from those in [5, 17].) Then, we obtain the following theorem.
+See [16, Section A.3].
+We refer to the proof of [16, Theorem 3.7] which takes $\mathbb{Z}[H_{1}(\Lambda_{\pm})]$ as the coefficient rings.
+We modify the arguments in [16, Appendix A.6] to the present case where we take $\mathbb{Z}[H_{2}(Y,\Lambda_{\pm})]$ as the coefficient rings.
+(As referred in [16, Appendix A], see also [10, Theorem 2.5] about the DGA map over $\mathbb{Z}$, but note that some conventions in [10] related to orientations are different from those in [4, 16].) Then, we obtain the following theorem.
 Theorem 2.3.
@@
 2.5.1 Chekanov-Eliashberg DGA of the unit conormal bundle of a knot in $\mathbb{R}^{3}$
-Let $(q_{1},q_{2},q_{3})$ be the coordinate on $\mathbb{R}^{3}$.
+Let $(q_{1},q_{2},q_{3})$ be the coordinate of $\mathbb{R}^{3}$.
 The manifold $\mathbb{R}^{3}$ is equipped with a standard Riemannian metric $\langle\cdot,\cdot\rangle\coloneqq\sum_{i=1}^{3}dq_{i}\otimes dq_{i}$ and an orientation such that $dq_{1}\wedge dq_{2}\wedge dq_{3}$ is a positive volume form.
-A coordinate $((q_{1},q_{2},q_{3}),(p_{1},p_{2},p_{3}))$ on $T^{*}\mathbb{R}^{3}$ is defined via the diffeomorphism
+A coordinate $((q_{1},q_{2},q_{3}),(p_{1},p_{2},p_{3}))$ of $T^{*}\mathbb{R}^{3}$ is defined via the diffeomorphism
 $\mathbb{R}^{3}\times\mathbb{R}^{3}\to T^{*}\mathbb{R}^{3}\colon((q_{1},q_{2},q_{3}),(p_{1},p_{2},p_{3}))\mapsto\sum_{i=1}^{3}p_{i}(dq_{i})_{(q_{1},q_{2},q_{3})}.$
@@
 We focus on a special kind of Legendrian submanifolds.
-For any knot $K$ in $\mathbb{R}^{3}$, recall the conormal bundle $L_{K}$ defined in Section 1.
+For any knot $K$ in $\mathbb{R}^{3}$, recall its conormal bundle $L_{K}$ defined in Section 1.
 Then, the unit conormal bundle of $K$ is defined by
@@
 We identify a disk bundle of $L_{K}$ with a tubular neighborhood of $K$ in $\mathbb{R}^{3}$ via the map $L_{K}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to\mathbb{R}^{3}\colon(q,p)\mapsto q+p$ for sufficiently small $\varepsilon>0$.
-Then, an orientation on $L_{K}$ is determined from the orientation on $\mathbb{R}^{3}$.
+Then, an orientation of $L_{K}$ is determined from the orientation of $\mathbb{R}^{3}$.
 We also fix a spin structure $\mathfrak{s}_{L_{K}}$ given by a trivial $\mathrm{Spin}(3)$-bundle on $L_{K}$.
-As the positive end of $L_{K}$, an orientation and a spin structure $\mathfrak{s}_{K}$ on $\Lambda_{K}$ are determined as in [17, Remark 3.6].
-We fix an orientation on $K$.
+As the positive end of $L_{K}$, an orientation and a spin structure $\mathfrak{s}_{K}$ on $\Lambda_{K}$ are determined as in [16, Remark 3.6].
+We fix an orientation of $K$.
 Let $l\colon K\to\Lambda_{K}$ be a section which gives the zero framing of $K$ and $m\colon S^{1}\to\Lambda_{K}$ be the meridian of $K$.
@@
 $\{(q,p)\in S^{*}\mathbb{R}^{3}\mid q\in K,\ p(V_{q})\geq 0\}.$
-Therefore, the homology class $[\Lambda_{K}]\in H_{2}(S^{*}\mathbb{R}^{3})$ vanishes. For the pair $(S^{*}\mathbb{R}^{3},\Lambda_{K})$, we have a short exact sequence
-[…]
+Therefore, the homology class $\Lambda_{K}]\in H_{2}(S^{*}\mathbb{R}^{3})$ vanishes. For the pair $(S^{*}\mathbb{R}^{3},\Lambda_{K})$, we have a short exact sequence
+${H_{2}(\Lambda_{K})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{0}$${H_{2}(S^{*}\mathbb{R}^{3})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\partial}$${H_{1}(\Lambda_{K})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{1}(S^{*}\mathbb{R}^{3})=0.}$
 We choose a basis $\{\widehat{l},\widehat{m},\widehat{s}\}$ of $H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K})$ such that $\partial(\widehat{l})=[l]$, $\partial(\widehat{m})=[m]$ and $\widehat{s}$ is represented by the fiber $S^{*}_{q}\mathbb{R}^{3}$ of any $q\in\mathbb{R}^{3}$.
-For this choice, we refer to the remark after the proof of Theorem 1.2 in [4, Section 3.7], which means that we use a knot isotopic to $K$ and transversal to the contact structure $\operatorname{Ker}(dq_{3}-q_{1}dq_{2}+q_{2}dq_{1})$ on $\mathbb{R}^{3}$.
+For this choice, we refer to the remark after the proof of Theorem 1.2 in [3, Section 3.7], using a knot which is isotopic to $K$ and transverse to the contact structure $\operatorname{Ker}(dq_{3}-q_{1}dq_{2}+q_{2}dq_{1})$ on $\mathbb{R}^{3}$.
 Remark 2.4.
@@
 Let us make a setup to deal with Question 1.1.
-Let $K_{0}$ and $K_{1}$ be oriented knots in $\mathbb{R}^{3}$ and tak $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$.
+Let $K_{0}$ and $K_{1}$ be oriented knots in $\mathbb{R}^{3}$ and let $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$.
 Suppose that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along $K_{1}$.
-From [17, Lemma 2.3], we may assume that there exists $\varepsilon>0$ such that
+By moving $\varphi(L_{K_{0}})$ by a Hamiltonian isotopy supported in a neighborhood of $\varphi(L_{K_{0}})\cap\mathbb{R}^{3}=K_{1}$, we may assume that there exists $\varepsilon>0$ such that
 $\varphi(L_{K_{0}})\cap D_{\varepsilon}^{*}\mathbb{R}^{3}=L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}$
-by moving $\varphi(L_{K_{0}})$ by a Hamiltonian isotopy supported in a neighborhood of $\varphi(L_{K_{0}})\cap\mathbb{R}^{3}=K_{1}$. We consider a diffeomorphism
+For the proof, see [16, Lemma 2.3]. We consider a diffeomorphism
 $F\colon\mathbb{R}\times S^{*}\mathbb{R}^{3}\to T^{*}\mathbb{R}^{3}\setminus\mathbb{R}^{3}\colon(r,(q,p))\mapsto(q,e^{r}p),$
@@
 $L_{\varphi}\coloneqq F^{-1}(\varphi(L_{K_{0}})\setminus K_{1})$
-is an exact Lagrangian cobordism from $\Lambda_{K_{1}}$ to $\Lambda_{K_{0}}$. The inclusion maps (1) are written as
-$i_{+}\colon(S^{*}\mathbb{R}^{3},\Lambda_{K_{0}})\to(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi})\colon(q,p)\mapsto(r_{+},(q,p)),$
-$i_{-}\colon(S^{*}\mathbb{R}^{3},\Lambda_{K_{1}})\to(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi})\colon(q,p)\mapsto(r_{-},(q,p)),$
-for some $r_{+},r_{-}\in\mathbb{R}$ with $r_{+}>r_{-}$.
-The Lagrangian filling $\varphi(L_{K_{0}})$ of $\Lambda_{K_{0}}$ in $T^{*}\mathbb{R}^{3}$ is obtained by gluing the Lagrangian cobordism $L_{\varphi}$ in ×$\mathbb{R}\times S^{*}\mathbb{R}^{3}$ and the Lagrangian filling $L_{K_{1}}$ in $T^{*}\mathbb{R}^{3}$ along $\Lambda_{K_{1}}$.
+is an exact Lagrangian cobordism from $\Lambda_{K_{1}}$ to $\Lambda_{K_{0}}$. We have the inclusion maps (1)
+$\begin{array}[]{cc}i_{+}\colon(S^{*}\mathbb{R}^{3},\Lambda_{K_{0}})\to(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi}),&i_{-}\colon(S^{*}\mathbb{R}^{3},\Lambda_{K_{1}})\to(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi}).\end{array}$
+The Lagrangian filling $\varphi(L_{K_{0}})$ of $\Lambda_{K_{0}}$ in $T^{*}\mathbb{R}^{3}$ is obtained by gluing the Lagrangian cobordism $L_{\varphi}$ in $\mathbb{R}\times S^{*}\mathbb{R}^{3}$ and the Lagrangian filling $L_{K_{1}}$ in $T^{*}\mathbb{R}^{3}$ along $\Lambda_{K_{1}}$.
 Therefore, as a part of the Mayer-Vietoris exact sequence, we have a short exact sequence
-[…]
+${0=H_{3}(T^{*}\mathbb{R}^{3},\varphi(L_{K_{0}}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K_{1}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\begin{matrix}H_{2}(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi})\\ \oplus H_{2}(T^{*}\mathbb{R}^{3},L_{K_{1}})\end{matrix}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{u}$${H_{2}(T^{*}\mathbb{R}^{3},\varphi(L_{K_{0}})).}$
 Note that $H_{2}(T^{*}\mathbb{R}^{3},L_{K_{1}})\cong H_{1}(L_{K_{1}})=\mathbb{Z}\cdot[K_{1}]$ and $H_{2}(T^{*}\mathbb{R}^{3},\varphi(L_{K_{0}}))\cong H_{1}(L_{K_{0}})=\mathbb{Z}\cdot[K_{0}]$.
-The left map is a surjection since $(i_{+})_{*}(\widehat{l})\in H_{2}(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi})$ is mapped to a generator of $H_{2}(T^{*}\mathbb{R}^{3},\varphi(L_{K_{0}}))$.
+The map $u$ is a surjection since $(i_{+})_{*}(\widehat{l})\in H_{2}(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi})$ is mapped to a generator of $H_{2}(T^{*}\mathbb{R}^{3},\varphi(L_{K_{0}}))$.
 It follows that $H_{2}(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi})$ is a free abelian group of rank $3$ with a basis $\{\widehat{l}_{0},\widehat{m}_{0},\widehat{s}_{0}\}$ defined by
-$\widehat{l}_{0}$ $\coloneqq(i_{+})_{*}(\widehat{l})\text{ for }\widehat{l}\in H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K_{0}}),$
-$\widehat{m}_{0}$ $\coloneqq(i_{-})_{*}(\widehat{m})\text{ for }\widehat{m}\in H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K_{1}}),$
-$\widehat{s}_{0}$ [$\coloneqq[\{0\}\times S^{*}_{q}\mathbb{R}^{3}]\text{ for any }q\in\mathbb{R}^{3}.$
+$\begin{cases}\widehat{l}_{0}\coloneqq(i_{+})_{*}(\widehat{l})&\text{for }\widehat{l}\in H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K_{0}}),\\ \widehat{m}_{0}\coloneqq(i_{-})_{*}(\widehat{m})&\text{for }\widehat{m}\in H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K_{1}}),\\ \widehat{s}_{0}\coloneqq[\{0\}\times S^{*}_{q}\mathbb{R}^{3}]&\text{for any }q\in\mathbb{R}^{3}.\end{cases}$
 We abbreviate
@@
 The open embedding $\varphi^{-1}\colon L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to L_{K_{0}}$ preserves the orientations of $L_{K_{1}}$ and $L_{K_{0}}$.
-Moreover, $\varphi_{*}([K_{0}])=a\cdot[K_{0}]$ in $H_{1}(\varphi(L_{K_{0}}))$ for $a\in\{1,-1\}$.
+Moreover, $(\left.\varphi\right|_{L_{K_{0}}})_{*}([K_{0}])=a\cdot[K_{0}]$ in $H_{1}(\varphi(L_{K_{0}}))$ for $a\in\{1,-1\}$.
 The proof is left to Appendix A.
-Note that the first assertion about orientations was shown in [17, Lemma 4.4] and a version of the second assertion with coefficients in $\mathbb{F}_{2}$ was shown by [17, Proposition 2.8, Lemma 4.5] using Floer cohomology group over $\mathbb{F}_{2}$ with the triangle product.
-It should be possible to prove Proposition 2.7 in a way parallel to [17, Section 2.3] by lifting the coefficient ring from $\mathbb{F}_{2}$ to $\mathbb{Z}$.
+Note that the first assertion about orientations was shown in [16, Lemma 4.4] and a version of the second assertion with coefficients in $\mathbb{F}_{2}$ was shown by [16, Proposition 2.8, Lemma 4.5] using Floer cohomology group over $\mathbb{F}_{2}$ together with the triangle product.
+It should be possible to prove Proposition 2.7 in a way parallel to [16, Section 2.3] by lifting the coefficient ring from $\mathbb{F}_{2}$ to $\mathbb{Z}$.
 However, we will give a simpler proof in Appendix A by using Floer cohomology with coefficients in $\mathbb{F}_{2}[H_{1}(L_{K_{0}})]\cong\mathbb{F}_{2}[\lambda^{\pm}]$.
@@
 It is clear that $(i_{\pm})_{*}\colon R_{\pm}\to R_{0}$ maps $U$ to $U_{0}$.
-The circle $\varphi(K_{0})$ is oriented via $\varphi\colon K_{0}\to\varphi(K_{0})$.
-From Proposition 2.7, we take $a\in\{\pm 1\}$ such that $[K_{1}]=a[\varphi(K_{0})]$.
+The circle $\varphi(K_{0})$ in $\varphi(L_{K_{0}})$ is oriented via the diffeomorphism $\left.\varphi\right|_{K_{0}}\colon K_{0}\to\varphi(K_{0})$.
+From Proposition 2.7, we take $a\in\{\pm 1\}$ such that $K_{1}]=a\cdot[\varphi(K_{0})]$ in $H_{1}(\varphi(L_{K_{0}}))$.
 Consider the map induced by $i_{\pm}$ on the first homology groups
 $\begin{array}[]{ll}(i_{+})_{*}\colon H_{1}(\Lambda_{K_{0}})\to H_{1}(L_{\varphi}),&(i_{-})_{*}\colon H_{1}(\Lambda_{K_{1}})\to H_{1}(L_{\varphi}).\end{array}$
-A basis of $H_{1}(L_{\varphi})$ is given by $\{l_{0},m_{0}\}$ for $l_{0}\coloneqq(i_{+})_{*}([l])$ and $m_{0}\coloneqq(i_{-})_{*}([m])$. It suffices to show that for some $b\in\mathbb{Z}$,
-$\begin{array}[]{cc}(i_{-})_{*}([l])=a\cdot l_{0}+b\cdot m_{0},&(i_{+})_{*}([m])=a\cdot m_{0}.\end{array}$ (5)
-By the map $H_{1}(L_{\varphi})\to H_{1}(\varphi(L_{K_{0}}))$ induced by $\left.F\right|_{L_{\varphi}}\colon L_{\varphi}\to\varphi(L_{K_{0}})$, $l_{0}$ is mapped to $[\varphi(K_{0})]$ and $m_{0}$ is mapped to 00.
-Since $(\left.F\right|_{L_{\varphi}}\circ i_{-})_{*}([l])=[K_{1}]=a[\varphi(K_{0})]$ in $H_{1}(\varphi(L_{K_{0}}))$, the coefficient of $l_{0}$ in $(i_{-})_{*}([l])$ is $a$.
-This shows the first equation of (5) for some $b\in\mathbb{Z}$.
+A basis of $H_{1}(L_{\varphi})$ is given by $\{l_{0},m_{0}\}$ for $l_{0}\coloneqq(i_{+})_{*}([l])$ and $m_{0}\coloneqq(i_{-})_{*}([m])$. It suffices to show that the equations
+$\begin{array}[]{cc}(i_{-})_{*}([l])=a\cdot l_{0}+b\cdot m_{0},&(i_{+})_{*}([m])=a\cdot m_{0},\end{array}$ (5)
+hold for some $b\in\mathbb{Z}$.
+By the map $H_{1}(L_{\varphi})\to H_{1}(\varphi(L_{K_{0}}))$ induced by $\left.F\right|_{L_{\varphi}}\colon L_{\varphi}\to\varphi(L_{K_{0}})$, $l_{0}$ is mapped to $\varphi(K_{0})]$ and $m_{0}$ is mapped to 0.
+Since $(\left.F\right|_{L_{\varphi}}\circ i_{-})_{*}([l])=[K_{1}]=a\cdot[\varphi(K_{0})]$ in $H_{1}(\varphi(L_{K_{0}}))$, the coefficient of $l_{0}$ in $(i_{-})_{*}([l])$ is $a$.
+This shows that the first equation of (5) holds for some $b\in\mathbb{Z}$.
 Next, take an embedded disk $D\subset\varphi(L_{K_{0}})$ which bounds the circle $(F\circ i_{+}\circ m)(S^{1})$. $D$ is oriented so that
-$[\partial D]=(\left.F\right|_{L_{\varphi}})_{*}((i_{+})_{*}([m]))\text{ in }H_{1}(\varphi(L_{K_{0}})\setminus K_{1})$
-for $\left.F\right|_{L_{\varphi}}\colon L_{\varphi}\to\varphi(L_{K_{0}})\setminus K_{1}$.
+[…]
+with respect to the diffeomorphism $\left.F\right|_{L_{\varphi}}\colon L_{\varphi}\to\varphi(L_{K_{0}})\setminus K_{1}$.
 We may assume that $D$ intersects $K_{1}$ transversely.
-Then, for sufficiently small $\varepsilon>0$, $D\cap(L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3})=\coprod_{i\in I}D^{\prime}_{i}$, where $D^{\prime}_{i}\subset D$ is a small disk and $I$ is a finite set.
-Then, using $a\in\{\pm 1\}$ determined by $[K_{1}]=a[\varphi(K_{0})]$, we have
+Then, their algebraic intersection number is equal to $a\in\{\pm 1\}$.
+For sufficiently small $\varepsilon>0$, $D\cap(L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3})=\coprod_{i\in I}D^{\prime}_{i}$, where $D^{\prime}_{i}\subset D$ is a small disk and $I$ is a finite set.
+The summand of $\partial D_{i}]\in H_{1}(\varphi(L_{K_{0}})\setminus K_{1})$ for all $i\in I$ satisfy
 $\sum_{i\in I}[\partial D_{i}]=a\cdot(\left.F\right|_{L_{\varphi}}\circ i_{-})_{*}([m])=a\cdot(\left.F\right|_{L_{\varphi}})_{*}(m_{0})\text{ in }H_{1}(\varphi(L_{K_{0}})\setminus K_{1}).$
-For the first equality, note that the inclusion map $L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to\varphi(L_{K_{0}})$ preserves orientations by Proposition 2.7 when $\varphi(L_{K_{0}})$ is oriented via the map $\varphi\colon L_{K_{0}}\to\varphi(L_{K_{0}})$.
-Let $\Sigma$ be the closure of $F^{-1}(D\setminus\coprod_{i\in I}D^{\prime}_{i})$ in $L_{\varphi}$.
+For the first equality, note that the inclusion map $L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to\varphi(L_{K_{0}})$ preserves orientations by Proposition 2.7 when $\varphi(L_{K_{0}})$ is oriented via the diffeomorphism $\left.\varphi\right|_{L_{K_{0}}}\colon L_{K_{0}}\to\varphi(L_{K_{0}})$.
+Let $\Sigma$ be the closure of $(\left.F\right|_{L_{\varphi}})^{-1}(D\setminus\coprod_{i\in I}D^{\prime}_{i})$ in $L_{\varphi}$.
 Then,
@@
 $L_{K_{0}}$ is equipped with the spin structure $\mathfrak{s}_{L_{K_{0}}}$ given by a trivial $\mathrm{Spin}(3)$-bundle.
-It is restricted on $L_{K_{0}}\setminus\varphi^{-1}(K_{1})$, then via the diffeomorphism $F^{-1}\circ\varphi\colon L_{K_{0}}\setminus\varphi^{-1}(K_{1})\to L_{\varphi}$, we obtain a spin structure $s_{L_{\varphi}}$ on the exact Lagrangian cobordism $L_{\varphi}$.
+We restrict $\mathfrak{s}_{L_{K_{0}}}$ on $L_{K_{0}}\setminus\varphi^{-1}(K_{1})$.
+Then via the diffeomorphism $F^{-1}\circ\varphi\colon L_{K_{0}}\setminus\varphi^{-1}(K_{1})\to L_{\varphi}$, we obtain a spin structure $s_{L_{\varphi}}$ on the exact Lagrangian cobordism $L_{\varphi}$.
 This induces $\mathfrak{s}_{K_{0}}$ on the positive end $\Lambda_{K_{0}}$ and $\mathfrak{s}_{K_{1}}$ on the negative end $\Lambda_{K_{0}}$.
 (Here, we need to care about the underlying orientations of spin manifolds.
-See the arguments in [17, Subsection 4.2.3].) Then, the existence of a DGA map $\Phi_{L_{\varphi}}$ follows from Theorem 2.3 and the description of the base change from $R_{-}$ to $R_{+}$ follows from Lemma 2.6.
+See the arguments in [16, Subsection 4.2.3].) Then, the existence of a DGA map $\Phi_{L_{\varphi}}$ follows from Theorem 2.3 and the description of the base change from $R_{-}$ to $R_{+}$ follows from Lemma 2.6.
 ∎
-Lastly, let us introduce a spin structure $\mathfrak{s}^{0}_{K}$ on $\Lambda_{K}$ as in [7, Section 6.2].
+Lastly, let us introduce a spin structure $\mathfrak{s}^{0}_{K}$ on $\Lambda_{K}$ as in [6, Section 6.2].
 When $T\Lambda_{K}$ is trivialized as an oriented vector bundle, then $\mathfrak{s}^{0}_{K}$ is given by the trivial $\mathrm{Spin}(2)$-bundle on $\Lambda_{K}$.
-It determines the orientations of moduli spaces which are used to define the differential of the DGAs in [4, Section 2.3.4] and [7, Section 3.6].
+It determines the orientations of moduli spaces which are used to define the differential of the DGAs in [3, Section 2.3.4] and [6, Section 3.6].
 The ring isomorphism $\psi(\mathfrak{s}^{0}_{K},\mathfrak{s}_{K})$ from (2) is given by
@@
 We will combine Proposition 2.8 with the results in [3, 14, 15] about the knot DGA.
-Some conventions about the knot DGA in these papers are slightly different from each other, as summarized in [16, Appendix A].
+Some conventions about the knot DGA in these papers are slightly different from each other, as summarized in [15, Appendix A].
 In this section, we mainly refer to [15] for the definition of the knot DGA.
 Let $K$ be an oriented knot in $\mathbb{R}^{3}$ given by the closure of a braid $B$.
-Then, a DGA over $R$, called the knot DGA of $K$, is constructed from $B$ as in [16, Definition 3.11].
+Then, a DGA over $R$, called the knot DGA of $K$, is constructed from $B$ as in [15, Definition 3.11].
 Its stable tame isomorphism class as a DGA over $R$ depends only on the isotopy class of $K$ as an oriented knot.
-See [15, Theorem 1.4] and [16, Theorem 3.1].
+See [14, Theorem 1.4] and [15, Theorem 3.1].
 Let us denote this DGA by $(\mathcal{A}^{\mathrm{knot}}_{*}(K),\partial_{K})$ and its homology $\operatorname{Ker}\partial_{K}/\operatorname{Im}\partial_{K}$ by $HC_{*}(K)$.
-We will review a description of the 00-th degree part $HC_{0}(K)$ in Section 3.2.
+We will review a description of the 0-th degree part $HC_{0}(K)$ in Section 3.2.
 Theorem 3.1 (Theorem 1.2 of [3], Theorem 3.14 and Appendix A of [15] ).
@@
 $\psi^{\prime}\colon R\to R\colon\lambda\mapsto-\lambda,\ \mu\mapsto-\mu,\ U\mapsto U.$
-The base change by $\psi^{\prime}$ is due to the difference of conventions between [4] and [16]. See [16, Appendix A]. It follows from this theorem that there are unital graded ring homomorphisms
+The base change by $\psi^{\prime}$ is due to the difference of conventions between [3] and [15]. See [15, Appendix A]. It follows from this theorem that there are unital graded ring homomorphisms
 $\begin{array}[]{cc}I_{K}\colon\mathcal{A}_{*}(\Lambda_{K})\to\mathcal{A}^{\mathrm{knot}}_{*}(K),&\bar{I}_{K}\colon\mathcal{A}^{\mathrm{knot}}_{*}(K)\to\mathcal{A}_{*}(\Lambda_{K}),\end{array}$ (8)
@@
 For $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$, suppose that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along $K_{1}$.
-Then we have the following theorem analogous to [17, Proposition 4.6].
+Then we have the following theorem analogous to [16, Proposition 4.6].
 Theorem 3.2.
@@
 $\Psi_{\varphi}(r\cdot x)=\psi(r)\cdot\Psi_{\varphi}(x)$ (9)
-for every $x\in\mathcal{A}^{\mathrm{knot}}_{*}(K_{0})$ and $r\in R$, where $\psi$ a ring isomorphism defined by
+for every $x\in\mathcal{A}^{\mathrm{knot}}_{*}(K_{0})$ and $r\in R$, where $\psi$ is a ring isomorphism defined by
 $\psi\colon R\to R\colon\lambda\mapsto\lambda^{a}\mu^{b}U^{c},\ \mu\mapsto\mu^{a}U^{d},\ U\mapsto U.$
 Proof.
-A unital graded ring homomorphism $\Psi_{\varphi}$ satisfying $\Psi_{\varphi}\circ\partial_{K_{0}}=\partial_{K_{1}}\circ\Psi_{\varphi}$ is defined as the composition of the following morphisms of differential graded rings:
-[…]
+A unital graded ring homomorphism $\Psi_{\varphi}$ satisfying $\Psi_{\varphi}\circ\partial_{K_{0}}=\partial_{K_{1}}\circ\Psi_{\varphi}$ is defined by the composition of the following morphisms of differential graded rings:
+${(\mathcal{A}^{\mathrm{knot}}_{*}(K_{0}),\partial_{K_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\bar{I}_{K_{0}}}$${(\mathcal{A}_{*}(\Lambda_{K_{0}}),\partial_{J,\mathfrak{s}^{0}_{K_{0}}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$$\scriptstyle{I(\mathfrak{s}^{0}_{K_{0}},\mathfrak{s}_{K_{0}})}$${(\mathcal{A}_{*}(\Lambda_{K_{0}}),\partial_{J,\mathfrak{s}_{K_{0}}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{L_{\varphi}}}$${(\mathcal{A}^{\mathrm{knot}}_{*}(K_{1}),\partial_{K_{1}})}$${(\mathcal{A}_{*}(\Lambda_{K_{1}}),\partial_{J,\mathfrak{s}^{0}_{K_{1}}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{I_{K_{1}}}$${(\mathcal{A}_{*}(\Lambda_{K_{1}}),\partial_{J,\mathfrak{s}_{K_{1}}}).\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$$\scriptstyle{I(\mathfrak{s}^{0}_{K_{1}},\mathfrak{s}_{K_{1}})^{-1}}$
 Here, $\bar{I}_{K_{0}}$ and $I_{K_{1}}$ are the morphisms (8) from Theorem 3.1.
@@
 $\psi(\mathfrak{s}^{0}_{K_{i}},\mathfrak{s}_{K_{i}})\circ(\psi^{\prime})^{-1}\colon\lambda\mapsto-\lambda,\ \mu\mapsto\mu,\ U\mapsto U,$
-and $\psi$ is described as in the theorem for the integers $a,b,c$ and $d$ which determine (3). ∎
-We recall the notion of an augmentation of a DGA. See, for instance, [16, Definition 2.6].
+then $\psi$ is described as in the theorem by using the integers $a,b,c$ and $d$ which determine (3). ∎
+We recall the definition of augmentations of a DGA. See, for instance, [15, Definition 2.6].
 Definition 3.3.
-Let $\mathbb{k}$ be a field and $(\mathcal{A}_{*},\partial)$ be a DGA over $R$.
+Let $\mathbb{k}$ be a field and let $(\mathcal{A}_{*},\partial)$ be a unital non-commutative DGA over $R$.
 A ring homomorphism $\varepsilon\colon\mathcal{A}_{*}\to\mathbb{k}$ is called an augmentation of $(\mathcal{A}_{*},\partial)$ to $\mathbb{k}$ if $\varepsilon(\mathbb{1})=1$, $\varepsilon(x)=0$ for any $x\in\mathcal{A}_{p}$ with $p\neq 0$, and $\varepsilon\circ\partial\colon\mathcal{A}_{1}\to\mathbb{k}$ is the zero map.
 Here, $\mathbb{1}$ is the unit of $\mathcal{A}_{*}$.
-A knot invariant is defined from the set of augmentations of the knot DGA to $\mathbb{C}$.
+In the following, given an augmentation $\varepsilon$ as in the definition, we abbreviate $\varepsilon(r\cdot\mathbb{1})$ by $\varepsilon(r)$ for any $r\in R$.
+Let us prepare a lemma which is necessary to deduce Proposition 3.10 and Proposition 3.17.
+Lemma 3.4.
+Let $(\mathcal{A}_{*},\partial)$ be a unital non-commutative DGA over $R$ and let $\mathbb{k}$ be a field.
+Let $H^{\mathrm{ab}}_{0}(\mathcal{A}_{*},\partial)$ denote the commutative $R$-algebra obtained by the abelianization of the 0-th homology of $(\mathcal{A}_{*},\partial)$.
+Suppose that there exist a polynomial $P\in R[T]$ and a unital $R$-algebra map
+$J\colon R[T]/(P)\to H^{\mathrm{ab}}_{0}(\mathcal{A}_{*},\partial).$
+Then, for any augmentation $\varepsilon$ of $(\mathcal{A}_{*},\partial)$ to $\mathbb{k}$ such that $(\varepsilon(\lambda),\varepsilon(\mu),\varepsilon(U))=(x,y,Z)\in(\mathbb{k}^{*})^{3}$, the polynomial $\left.P\right|_{\lambda=x,\mu=y,U=Z}\in\mathbb{k}[T]$ has a root in $\mathbb{k}$.
+Proof.
+Any augmentation $\varepsilon$ of $(\mathcal{A}_{*},\partial)$ to $\mathbb{k}$ such that $(\varepsilon(\lambda),\varepsilon(\mu),\varepsilon(U))=(x,y,Z)$ induces a unital ring homomorphism $\bar{\varepsilon}\colon HC^{\mathrm{ab}}_{0}(\mathcal{A}_{*},\partial)\to\mathbb{k}$ such that $(\bar{\varepsilon}(\lambda),\bar{\varepsilon}(\mu),\bar{\varepsilon}(U))=(x,y,Z)$.
+We obtain a unital ring homomorphism $\bar{\varepsilon}\circ J\colon R[T]/(P)\to\mathbb{k}$, and $(\bar{\varepsilon}\circ J)(T)\in\mathbb{k}$ is a root of $\left.P\right|_{\lambda=x,\mu=y,U=Z}\in\mathbb{k}[T]$.
+∎
+The next proposition is used to prove Theorem 3.15 and Theorem 3.19 by taking $\mathbb{k}=\mathbb{Q}$.
+Proposition 3.5.
+Suppose that $K_{1}$ is the unknot.
+Then, for any $y,Z\in\mathbb{k}^{*}$ satisfying $y\neq Z^{n}$ for every $n\in\mathbb{Z}$, there exists an augmentation $\varepsilon$ of $(\mathcal{A}^{\mathrm{knot}}_{*}(K_{0}),\partial_{K_{0}})$ to $\mathbb{k}$ such that $(\varepsilon(\mu),\varepsilon(U))=(y,Z)$.
+Proof.
+If $K_{1}$ is the unknot, from the trivial braid on a single strand, we can compute as in [15, Example 3.13] that $\mathcal{A}^{\mathrm{knot}}_{0}(K_{1})=R=\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$ and $\partial_{K_{1}}(\mathcal{A}_{1}^{\mathrm{knot}}(K_{1}))$ is the ideal generated by
+$U-\lambda-\mu+\lambda\mu.$
+Let us take $a\in\{\pm 1\}$ and $b,c,d\in\mathbb{Z}$ to be the integers in Theorem 3.2.
+For any $y,Z\in\mathbb{k}^{*}$ satisfying $y\notin\{Z^{n}\mid n\in\mathbb{Z}\}$, we take an augmentation $\varepsilon_{1}\colon\mathcal{A}^{\mathrm{knot}}_{*}(K_{1})\to\mathbb{k}$ determined by
+$\begin{array}[]{ccc}\varepsilon_{1}(\lambda)=\displaystyle{\frac{Z-(y\cdot Z^{-d})^{a}}{1-(y\cdot Z^{-d})^{a}}},&\varepsilon_{1}(\mu)=(y\cdot Z^{-d})^{a},&\varepsilon_{1}(U)=Z.\end{array}$
+Note that $y\cdot Z^{-d}\neq 1$ and $y\cdot Z^{-d}\neq Z^{a}$ from the condition on $y$ and $Z$.
+Composing with the map $\Psi_{\varphi}$ of Theorem 3.2, we obtain $\varepsilon\coloneqq\varepsilon_{1}\circ\Psi_{\varphi}\colon\mathcal{A}^{\mathrm{knot}}_{*}(K_{0})\to\mathbb{k}$.
+It is an augmentation of the knot DGA of $K_{0}$ to $\mathbb{k}$ such that
+$\begin{array}[]{cc}\varepsilon(\mu)=\varepsilon_{1}(\mu^{a}U^{d})=y,&\varepsilon(U)=\varepsilon_{1}(U)=Z.\end{array}$
+Therefore, $\varepsilon$ satisfies the required condition. ∎
+Let us also consider a knot invariant defined from $\mathbb{C}$-valued augmentations of the knot DGA.
 Definition 3.6 (Definition 5.1 of [15]).
 For an oriented knot $K$ in $\mathbb{R}^{3}$, its augmentation variety is a subset of $(\mathbb{C}^{*})^{3}$ defined by
-$V_{K}\coloneqq\left\{(x,y,Z)\in(\mathbb{C}^{*})^{3}\ \middle|\begin{array}[]{l}\text{There exists an augmentation }\varepsilon\text{ of }(\mathcal{A}^{\mathrm{knot}}_{*}(K),\partial_{K})\text{ to }\mathbb{C}\\ \text{such that }\varepsilon(\lambda)=x,\ \varepsilon(\mu)=y\text{ and }\varepsilon(U)=Z\end{array}\right\}.$
-Here, $\varepsilon(r\cdot\mathbb{1})$ for any $r\in R$ is abbreviated by $\varepsilon(r)$.
-From Theorem 3.2, we obtain a refinement of [17, Theorem 1.1].
+[…]
+From Theorem 3.2, we obtain a refinement of [16, Theorem 1.1].
 Corollary 3.7.
-There exist integers $b^{\prime},c^{\prime}$ and $d^{\prime}$ for which we have an injective map
-$V_{K_{1}}\hookrightarrow V_{K_{0}}\colon(x,y,Z)\mapsto(xy^{b^{\prime}}Z^{c^{\prime}},yZ^{d^{\prime}},Z).$
+There exist $a\in\{\pm 1\}$ and integers $b,c$ and $d^{\prime}$ for which we have an injective map
+$V_{K_{1}}\hookrightarrow V_{K_{0}}\colon(x,y,Z)\mapsto(x^{a}y^{b}Z^{c},yZ^{d},Z).$
 Proof.
-By Theorem 3.2, for any augmentation $\varepsilon$ of $(\mathcal{A}_{*}^{\mathrm{knot}}(K_{1}),\partial_{K_{1}})$ to $\mathbb{C}$ with $(\varepsilon(\lambda),\varepsilon(\mu),\varepsilon(U))=(x,y,Z)$, the composite map $\varepsilon\circ\Psi_{\varphi}\colon\mathcal{A}_{*}^{\mathrm{knot}}(K_{0})\to\mathbb{C}$ is an augmentation of $(\mathcal{A}_{*}^{\mathrm{knot}}(K_{0}),\partial_{K_{0}})$ to $\mathbb{C}$ such that
+Using $\Psi_{\varphi}$ of Theorem 3.2, it follows that for any augmentation $\varepsilon$ of $(\mathcal{A}_{*}^{\mathrm{knot}}(K_{1}),\partial_{K_{1}})$ to $\mathbb{C}$ such that $(\varepsilon(\lambda),\varepsilon(\mu),\varepsilon(U))=(x,y,Z)$, the composite map $\varepsilon\circ\Psi_{\varphi}\colon\mathcal{A}_{*}^{\mathrm{knot}}(K_{0})\to\mathbb{C}$ is an augmentation of $(\mathcal{A}_{*}^{\mathrm{knot}}(K_{0}),\partial_{K_{0}})$ to $\mathbb{C}$ such that
 $(\varepsilon\circ\Psi_{\varphi}(\lambda),\varepsilon\circ\Psi_{\varphi}(\mu),\varepsilon\circ\Psi_{\varphi}(U))=(x^{a}y^{b}Z^{c},y^{a}Z^{d},Z).$
-From the definition of $V_{K_{i}}$ for $i\in\{0,1\}$, this implies that we have a map
-$V_{K_{1}}\to V_{K_{0}}\colon(x,y,Z)\mapsto(x^{a}y^{b}Z^{c},y^{a}Z^{d},Z).$
-We use the following fact by [15, Proposition 4.3]: Up to chain isomorphisms of DGA over $R$, $(\mathcal{A}_{*}^{\mathrm{knot}}(K_{0}),\partial_{K_{0}})$ is unchanged by the substitution $\lambda\mapsto\lambda^{-1}U$, $\mu\mapsto\mu^{-1}U$ and $U\mapsto U$.
-Hence, we have a bijection
-$V_{K_{0}}\to V_{K_{0}}\colon(x,y,Z)\mapsto(x^{-1}Z,y^{-1}Z,Z).$
-If $a=1$, we take $b^{\prime}\coloneqq b$, $c^{\prime}\coloneqq c$ and $d^{\prime}\coloneqq d$.
-If $a=-1$, we take $b^{\prime}\coloneqq-b$, $c^{\prime}\coloneqq-c+1$ and $d^{\prime}\coloneqq-d+1$.
-Then, the assertion of the corollary holds.
-∎
-3.2 Proof of Theorem 1.1
+From the definition of $V_{K_{i}}$ for $i\in\{0,1\}$, we obtain a map from $V_{K_{1}}$ to $V_{K_{0}}$ as in the assertion. ∎
+3.2 Proof of Theorem 1.1 for $K^{\prime}\#T_{(2,2m+1)}$ ($m\neq 0,-1$)
+3.2.1 Review of the definition of $HC_{0}(K)$
 First, let us review how to compute $HC_{0}(K)$ for an oriented knot $K$.
-We refer to [16, Section 3].
+We refer to [15, Section 3].
 Note that since $\mathcal{A}^{\mathrm{knot}}_{p}(K)=0$ for $p<0$, $HC_{0}(K)=\mathcal{A}^{\mathrm{knot}}_{0}(K)/\partial_{K}(\mathcal{A}^{\mathrm{knot}}_{1}(K))$.
-For any $n\in\mathbb{Z}_{\geq 1}$, let $B_{n}$ denote the group of braids on $n$ strands, which is generated by $\{\sigma_{1},\dots,\sigma_{n-1}\}$ with relations $\sigma_{i}\sigma_{i+1}\sigma_{i}=\sigma_{i+1}\sigma_{i}\sigma_{i+1}$ for $21\leq i\leq n-$ 2 and $\sigma_{i}\sigma_{j}=\sigma_{j}\sigma_{i}$ for $|i-j|\geq 2$.
+For any $n\in\mathbb{Z}_{\geq 1}$, let $B_{n}$ denote the group of braids on $n$ strands, which is generated by $\{\sigma_{1},\dots,\sigma_{n-1}\}$ with relations $\sigma_{i}\sigma_{i+1}\sigma_{i}=\sigma_{i+1}\sigma_{i}\sigma_{i+1}$ for $1\leq i\leq n-2$ and $\sigma_{i}\sigma_{j}=\sigma_{j}\sigma_{i}$ for $|i-j|\geq 2$.
 Let $\mathcal{A}_{n}$ be the $R$-algebra freely generated by the set $\{a_{i,j}\mid i,j\in\{1,\dots,n\},\ i\neq j\}$.
-Then, a group homomorphism $B_{n}\to\mathrm{Aut}(\mathcal{A}_{n})\colon B\mapsto\phi_{B}$ is defined as in [16, Definition 3.3].
+Then, a group homomorphism $B_{n}\to\mathrm{Aut}(\mathcal{A}_{n})\colon B\mapsto\phi_{B}$ is defined as in [15, Definition 3.3].
 It is determined by $\phi_{\sigma_{k}}\colon\mathcal{A}_{n}\to\mathcal{A}_{n}$ for each generator $\sigma_{k}$, and $\phi_{\sigma_{k}}$ is given by
-$\phi_{\sigma_{k}}\colon\begin{cases}a_{i,j}\mapsto a_{i,j}&\text{if }i,j\neq k,k+1,\\ a_{k+1,i}\mapsto a_{k,i}&\text{ if }i\neq k,k+1,\\ a_{i,k+1}\mapsto a_{i,k}&\text{ if }i\neq k,k+1,\\ a_{k,k+1}\mapsto-a_{k+1,k},&\\ a_{k+1,k}\mapsto-a_{k,k+1},&\\ a_{k,i}\mapsto a_{k+1,i}-a_{k+1,k}a_{k,i}&\text{ if }i\neq k,k+1,\\ a_{i,k}\mapsto a_{i,k+1}-a_{i,k}a_{k,k+1}&\text{ if }i\neq k,k+1.\end{cases}$ (10)
-See also [7, Section 2.1]. We remark that $\phi_{B\cdot B^{\prime}}=\phi_{B}\circ\phi_{B^{\prime}}$ for any $B,B^{\prime}\in B_{n}$.
-The automorphism $\phi_{B}\in\mathrm{Aut}(\mathcal{A}_{n})$ for $B\in B_{n}$ is extended to an automorphism on $\mathcal{A}_{n+1}$ as follows: We consider the group homomorphism $B_{n+1}\to\mathrm{Aut}(\mathcal{A}_{n+1})$ defined as above and an injective map $B_{n}\to B_{n+1}$ defined by adding the extra $(n+1)$-th strand to any braid on $n$ strands (see [16, Remark 3.5]).
+[…]
+See also [6, Section 2.1]. We remark that $\phi_{B\cdot B^{\prime}}=\phi_{B}\circ\phi_{B^{\prime}}$ for any $B,B^{\prime}\in B_{n}$.
+The automorphism $\phi_{B}\in\mathrm{Aut}(\mathcal{A}_{n})$ for $B\in B_{n}$ is extended to an automorphism on $\mathcal{A}_{n+1}$ as follows: We consider the group homomorphism $B_{n+1}\to\mathrm{Aut}(\mathcal{A}_{n+1})$ defined as above and an injective map $B_{n}\to B_{n+1}$ defined by adding the extra $(n+1)$-th strand to any braid on $n$ strands (see [15, Remark 3.5]).
 Then, for any $B\in B_{n}$, $\phi_{B}\in\mathrm{Aut}(\mathcal{A}_{n+1})$ is defined via the composite map $B_{n}\hookrightarrow B_{n+1}\to\mathrm{Aut}(\mathcal{A}_{n+1})$.
-In addition, let us label the $(n+1)$-st strand by * and denote $a_{i,n+1}$ by $a_{i,*}$ and $a_{n+1,i}$ by $a_{*,i}$ for every $i\in\{1,\dots,n\}$.
+In addition, let us label the $(n+1)$-th strand by * and denote $a_{i,n+1}$ by $a_{i,*}$ and $a_{n+1,i}$ by $a_{*,i}$ for every $i\in\{1,\dots,n\}$.
 We define $n\times n$ matrices $\mathbb{A}$ and $\widehat{\mathbb{A}}$ by
-$\begin{array}[]{ll}\mathbb{A}_{i,j}\coloneqq\begin{cases}a_{i,j}&\text{ if }i<j,\\ 1-\mu&\text{ if }i=j,\\ -\mu a_{i,j}&\text{ if }i>j,\end{cases}&\widehat{\mathbb{A}}_{i,j}\coloneqq\begin{cases}Ua_{i,j}&\text{ if }i<j,\\ U-\mu&\text{ if }i=j,\\ -\mu a_{i,j}&\text{ if }i>j.\end{cases}\end{array}$
+[…]
 Fix $B\in B_{n}$ which has the writhe $w$ (the sum of the exponents in the braid word).
-Let $\bm{\Lambda}$ be an $n\times n$ diagonal matrix with $\bm{\Lambda}_{1,1}=\lambda\mu^{w}U^{-(w-n+1)/2}$ and $\bm{\Lambda}_{i,i}=1$ for every $i\in\{2,\dots,n\}$.
+Let $\bm{\Lambda}$ be an $n\times n$ diagonal matrix such that $\bm{\Lambda}_{1,1}=\lambda\mu^{w}U^{-(w-n+1)/2}$ and $\bm{\Lambda}_{i,i}=1$ for every $i\in\{2,\dots,n\}$.
 Let $\bm{\Phi}^{L}_{B}$ be an $n\times n$ matrix whose entries $(\bm{\Phi}^{L}_{B})_{i,j}\in\mathcal{A}_{n}$ for $i,j\in\{1,\dots,n\}$ are determined by
@@
 Let $K$ be the oriented knot in $\mathbb{R}^{3}$ given by the closure of the braid $B$.
-By [16, Theorem 4.1, Remark 4.2], $HC_{0}(K)$ is the quotient of $\mathcal{A}_{n}$ by the two-sided ideal generated by all entries of the two matrices
+By [15, Theorem 4.1, Remark 4.2], $HC_{0}(K)$ is the quotient of $\mathcal{A}_{n}$ by the two-sided ideal generated by all entries of the two matrices
 $\begin{array}[]{cc}\mathbb{A}-\bm{\Lambda}\cdot\phi_{B}(\mathbb{A})\cdot\bm{\Lambda}^{-1},&\widehat{\mathbb{A}}-\bm{\Lambda}\cdot\bm{\Phi}^{L}_{B}\cdot\mathbb{A},\end{array}$ (12)
 where $(\phi_{B}(\mathbb{A}))_{i,j}\coloneqq\phi_{B}(\mathbb{A}_{i,j})$ for $i,j\in\{1,\dots,n\}$.
-The isomorphism class of $HC_{0}(K)$ as a unital $R$-algebra depends only on the isotopy class of $K$.
-See [16, Theorem 3.15] and [15, Theorem 1.4 (2)].
+The isomorphism class of $HC_{0}(K)$ as a unital $R$-algebra depends only on the isotopy class of $K$ as an oriented knot.
+See [15, Theorem 3.15] and [14, Theorem 1.4 (2)].
 We consider a commutative $R$-algebra $HC_{0}^{\mathrm{ab}}(K)$ defined as the quotient of $HC_{0}(K)$ by the relations $\xi\eta=\eta\xi$ for all $\xi,\eta\in HC_{0}(K)$.
-Using the $R$-algebra $R[a_{i,j}\mid 1\leq i,j\leq n,\ i\neq j]$ of polynomials and a natural map $\pi_{n}\colon\mathcal{A}_{n}\to R[a_{i,j}\mid 1\leq i,j\leq n,\ i\neq j]\colon a_{i,j}\mapsto a_{i,j}$, we have
+Using the $R$-algebra $R[a_{i,j}\mid 1\leq i,j\leq n,\ i\neq j]$ of polynomials and a natural map
+$\pi_{n}\colon\mathcal{A}_{n}\to R[a_{i,j}\mid 1\leq i,j\leq n,\ i\neq j]\colon a_{i,j}\mapsto a_{i,j},$ (13)
+we have
 $HC_{0}^{\mathrm{ab}}(K)=R[a_{i,j}\mid 1\leq i,j\leq n,\ i\neq j]/\mathcal{I}_{B},$
 where $\mathcal{I}_{B}$ is the ideal generated by the images under $\pi_{n}$ of all entries of the two matrices (12).
+3.2.2 On augmentations of $\mathcal{A}^{\mathrm{knot}}_{*}(K\#T_{(2,2m+1)})$ for $m\geq 0$
+Fix any $B\in B_{n}$ whose closure gives a knot $K$ in $\mathbb{R}^{3}$. We consider a braid $B^{\#}_{m}$ on $(n+2)$ strands defined by
+$B^{\#}_{m}\coloneqq\sigma_{n}\cdot(\sigma_{n+1})^{2m+1}\cdot B\in B_{n+2}$
+for every $m\in\mathbb{Z}_{\geq 0}$.
+Then, the closure of $B^{\#}_{m}$ gives the connected sum $K\#T_{(2,2m+1)}$ of $K$ and the $(2,2m+1)$-torus knot.
+See Figure 1 in the case of $m=1$.
+(When $m=0$, $T_{(2,1)}$ is the unknot.)
 […]
-Figure. 1: The braid $B^{\#}\in B_{n+2}$. The box with ‘$B$’ in the center represents the braid $B$ on $n$ strands.
-Associated to $B$, consider a braid $B^{\#}$ on $(n+2)$ strands defined by
-$B^{\#}\coloneqq\sigma_{n}\cdot(\sigma_{n+1})^{3}\cdot B\in B_{n+2}.$
-See Figure 1. The closure of $B^{\#}$ gives a knot which is the connected sum $K\#3_{1}$ of $K$ and the right-handed trefoil knot $3_{1}$.
-In general, $HC^{\mathrm{ab}}_{0}(K\#3_{1})$ is a complicated $R$-algebra. However, it has the following property.
+Figure. 1: The braid $B^{\#}_{m}\in B_{n+2}$ for $m=1$. The box with ‘$B$’ in the center represents the braid $B$ on $n$ strands.
+By using (10), let us compute $\phi_{B^{\#}_{m}}(a_{n+2,*})=\left(\phi_{\sigma_{n}}\circ\phi_{(\sigma_{n+1})^{2m+1}}\circ\phi_{B}\right)(a_{n+2,*})$.
+First, since $\phi_{\sigma_{k}}(a_{n+2,*})=a_{n+2,*}$ for every $k\in\{1,\dots,n-1\}$, $\phi_{B}(a_{n+2,*})=a_{n+2,*}$.
+Next, we compute $\phi_{(\sigma_{n+1})^{2m+1}}(a_{n+2,*})$ by the induction on $m=0,1,\dots$.
 Lemma 3.8.
-We define an $R$-coefficient polynomial $F\in R[X]$ by
-$F(X)\coloneqq X^{2}+(\mu+1-\mu^{-1}U)X+(\mu-1).$
-Then, there exists a unital $R$-algebra map
-$J\colon R[X]/(F)\to HC^{\mathrm{ab}}_{0}(K\#3_{1}).$
+Let $\hat{f}_{m}$ and $\hat{g}_{m}$ be elements of the free ring $\mathbb{Z}\langle a_{n+1,n+2},a_{n+2,n+1}\rangle$ for every $m\in\mathbb{Z}_{\geq 0}$ determined inductively on $m=0,1,\dots$ by $\hat{f}_{0}=1$, $\hat{g}_{0}=0$, and
+$\begin{cases}\hat{f}_{m+1}=\hat{f}_{m}\cdot(1-a_{n+1,n+2}a_{n+2,n+1})-\hat{g}_{m}\cdot a_{n+2,n+1},\\ \hat{g}_{m+1}=\hat{f}_{m}\cdot a_{n+1,n+2}+\hat{g}_{m}.\end{cases}$ (14)
+Then, for every $m\in\mathbb{Z}_{\geq 0}$,
+$\phi_{(\sigma_{n+1})^{2m+1}}(a_{n+2,*})=\hat{f}_{m}\cdot a_{n+1,*}+\hat{g}_{m}\cdot a_{n+2,*}.$ (15)
 Proof.
-We compute $\phi_{B^{\#}}(a_{n+2,*})$ by (10). Since $\phi_{\sigma_{k}}(a_{n+2,*})=a_{n+2,*}$ for every $k\in\{1,\dots,n-1\}$, $\phi_{B}(a_{n+2,*})=a_{n+2,*}$. Furthermore,
-${a_{n+2,*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{B}}$${a_{n+2,*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{\sigma_{n+1}}}$${a_{n+1,*}}$$\scriptstyle{\phi_{\sigma_{n+1}}}$${a_{n+2,*}-a_{n+2,n+1}a_{n+1,*}}$$\scriptstyle{\phi_{\sigma_{n+1}}}$${a_{n+1,*}+a_{n+1,n+2}(a_{n+2,*}-a_{n+2,n+1}a_{n+1,*})}$$\scriptstyle{\phi_{\sigma_{n}}}$${a_{n,*}+a_{n,n+2}(a_{n+2,*}-a_{n+2,n}a_{n,*}).}$
-Let us write $P=a_{n,n+2}$ and $Q=a_{n+2,n}$. Then, the bottom $(n+2)$-th row of $\bm{\Phi}^{L}_{B^{\#}}$ is
-$\begin{pmatrix}0&\cdots&0&1-PQ&0&P\end{pmatrix}.$
-It follows that the $(n+2,n)$-th entry of $\widehat{\mathbb{A}}-\bm{\Lambda}\cdot\bm{\Phi}^{L}_{B^{\#}}\cdot\mathbb{A}$ is
-$\widehat{\mathbb{A}}_{n+2,n}-((1-PQ)\mathbb{A}_{n,n}+P\mathbb{A}_{n+2,n})$ $=-\mu Q-((1-PQ)(1-\mu)+P(-\mu Q))$
-$=PQ-\mu Q-(1-\mu),$
-and the $(n+2,n+2)$-th entry of $\widehat{\mathbb{A}}-\bm{\Lambda}\cdot\bm{\Phi}^{L}_{B^{\#}}\cdot\mathbb{A}$ is
-$\widehat{\mathbb{A}}_{n+2,n+2}-((1-PQ)\mathbb{A}_{n,n+2}+P\mathbb{A}_{n+2,n+2})$ $=(U-\mu)-((1-PQ)P+P(1-\mu))$
-$=PQP-(2-\mu)P+(U-\mu).$
-Hence $P^{2}Q^{2}-(2-\mu)PQ+(U-\mu)Q\in\mathcal{I}_{B^{\#}}$ and by substituting $PQ$ with $\mu Q+(1-\mu)$, we can find an element of $\mathcal{I}_{B^{\#}}$
-$(\mu Q+(1-\mu))^{2}-(2-\mu)(\mu Q+(1-\mu))+(U-\mu)Q,$
-which is equal to
-$\mu^{2}Q^{2}+(2\mu(1-\mu)-(2-\mu)\mu+(U-\mu))Q+(1-\mu)^{2}-(2-\mu)(1-\mu)$
-=\ = $\mu^{2}Q^{2}+(-\mu^{2}-\mu+U)Q+(\mu-1).$
-Therefore, if we define $F(X)=X^{2}+(\mu+1-\mu^{-1}U)X+(\mu-1)$, we have a well-defined unital $R$-algebra map
-$J\colon R[X]/(F)\to HC^{\mathrm{ab}}_{0}(K\#3_{1})$
-which sends $X$ to $-\mu Q=-\mu a_{n+2,n}$. ∎
-Let us check this property in the simplest case.
+We prove by induction on $m$.
+Since $\phi_{\sigma_{n+1}}(a_{n+2,*})=a_{n+1,*}$, (15) for $m=0$ is satisfied by setting $\hat{f}_{0}=1$ and $\hat{g}_{0}=0$.
+Suppose that (15) holds for some $m\in\mathbb{Z}_{\geq 0}$.
+For readability, let us rewrite
+$\begin{array}[]{cccc}a_{n+1,n+2}=p,&a_{n+2,n+1}=q,&\hat{f}_{m}=\hat{f}_{m}(p,q),&\hat{g}_{m}=\hat{g}_{m}(p,q).\end{array}$
+Note that $\phi_{\sigma_{n+1}}$ maps $p$ to q-q and $q$ to p-p.
+Let $\hat{f}_{m}(-q,-p),\hat{g}_{m}(-q,-p)\in\mathbb{Z}\langle p,q\rangle$ be the elements obtained from $\hat{f}_{m}(p,q)$ and $\hat{g}_{m}(p,q)$ by replacing $p$ with q-q and $q$ with p-p.
+Then, we can compute that
+$\phi_{(\sigma_{n+1})^{2m+2}}(a_{n+2},*)$ $=\phi_{\sigma_{n+1}}\left(\hat{f}_{m}(p,q)\cdot a_{n+1,*}+\hat{g}_{m}(p,q)\cdot a_{n+2,*}\right)$
+$=\hat{f}_{m}(-q,-p)\cdot(a_{n+2,*}-q\cdot a_{n+1,*})+\hat{g}_{m}(-q,-p)\cdot a_{n+1,*}$
+$=\left(-\hat{f}_{m}(-q,-p)q+\hat{g}_{m}(-q,-p)\right)\cdot a_{n+1,*}+\hat{f}_{m}(-q,-p)\cdot a_{n+2,*},$
+$\phi_{(\sigma_{n+1})^{2m+3}}(a_{n+2},*)$ $=\left(-\hat{f}_{m}(p,q)(-p)+\hat{g}_{m}(p,q)\right)\cdot(a_{n+2,*}-q\cdot a_{n+1,*})+\hat{f}_{m}(p,q)\cdot a_{n+1,*}$
+$=\left(\hat{f}_{m}(p,q)(1-pq)-\hat{g}_{m}(p,q)q\right)\cdot a_{n+1,*}+\left(\hat{f}_{m}(p,q)p+\hat{g}_{m}(p,q)\right)\cdot a_{n+2,*}.$
+Therefore, if we set $\hat{f}_{m+1}$ and $\hat{g}_{m+1}$ by (14), then (15) with $m$ replaced by $m+1$ holds. ∎
+We apply $\phi_{\sigma_{n}}$ to (15), then
+$\phi_{B^{\#}_{m}}(a_{n+2,*})$ $=\left(\phi_{\sigma_{n}}\circ\phi_{(\sigma_{n+1})^{2m+1}}\circ\phi_{B}\right)(a_{n+2,*})$
+$=\left(\phi_{\sigma_{n}}\circ\phi_{(\sigma_{n+1})^{2m+1}}\right)(a_{n+2,*})$
+$=\phi_{\sigma_{n}}(\hat{f}_{m})\cdot a_{n,*}+\phi_{\sigma_{n}}(\hat{g}_{m})\cdot a_{n+2,*}.$
+Since $\phi_{\sigma_{n}}$ maps $a_{n+1,n+2}$ to $a_{n,n+2}$ and $a_{n+2,n+1}$ to $a_{n+2,n}$, the elements $\phi_{\sigma_{n}}(\hat{f}_{m}),\phi_{\sigma_{n}}(\hat{g}_{m})\in\mathbb{Z}\langle a_{n,n+2},a_{n+2,n}\rangle$ are obtained from $\hat{f}_{m}$ and $\hat{g}_{m}$ by replacing $a_{n+1,n+2}$ with $a_{n,n+2}$ and $a_{n+2,n+1}$ with $a_{n+2,n}$.
+It follows that the bottom ($n+2$)-th row of $\bm{\Phi}^{L}_{B^{\#}_{m}}$ is
+$\begin{pmatrix}0&\cdots&0&\phi_{\sigma_{n}}(\hat{f}_{m})&0&\phi_{\sigma_{n}}(\hat{g}_{m})\end{pmatrix}$.
+Let $c^{m}_{i,j}\in\mathcal{A}_{n+2}$ denote the $(i,j)$-th entry of $\widehat{\mathbb{A}}-\bm{\Lambda}\cdot\bm{\Phi}^{L}_{B^{\#}_{m}}\cdot\mathbb{A}$.
+We focus on the $(n+2,n)$-th and $(n+2,n+2)$-th entries:
+$\begin{split}c^{m}_{n+2,n}&=\widehat{\mathbb{A}}_{n+2,n}-\left(\phi_{\sigma_{n}}(\hat{f}_{m})\cdot\mathbb{A}_{n,n}+\phi_{\sigma_{n}}(\hat{g}_{m})\cdot\mathbb{A}_{n+2,n}\right)\\ &=-\mu a_{n+2,n}-\left(\phi_{\sigma_{n}}(\hat{f}_{m})\cdot(1-\mu)+\phi_{\sigma_{n}}(\hat{g}_{m})\cdot(-\mu a_{n+2,n})\right),\\ c^{m}_{n+2,n+2}&=\widehat{\mathbb{A}}_{n+2,n+2}-\left(\phi_{\sigma_{n}}(\hat{f}_{m})\cdot\mathbb{A}_{n,n+2}+\phi_{\sigma_{n}}(\hat{g}_{m})\cdot\mathbb{A}_{n+2,n+2}\right)\\ &=(U-\mu)-\left(\phi_{\sigma_{n}}(\hat{f}_{m})\cdot a_{n,n+2}+\phi_{\sigma_{n}}(\hat{g}_{m})\cdot(1-\mu)\right).\end{split}$ (16)
+Using the map $\pi_{n+2}$ of (13), let us define elements of $R[a_{i,j}\mid 1\leq i,j\leq n+2,\ i\neq j]$ by
+$\begin{array}[]{cc}f_{m}\coloneqq\pi_{n+2}(\phi_{\sigma_{n}}(\hat{f}_{m})),&g_{m}\coloneqq\pi_{n+2}(\phi_{\sigma_{n}}(\hat{f}_{m})),\end{array}$
+and denote $X=a_{n,n+2}$ and $Y=a_{n+2,n}$. By (14), $f_{m}$ and $g_{m}$ are polynomials in $\mathbb{Z}[X,Y]$ determined inductively by $f_{0}=1$, $g_{0}=0$ and
+$\begin{array}[]{cc}f_{m+1}=(1-XY)f_{m}-Yg_{m},&g_{m+1}=Xf_{m}+g_{m},\end{array}$ (18)
+for every $m\in\mathbb{Z}_{\geq 0}$. Let us also define $F_{m},G_{m}\in\mathbb{Z}[\mu][X,Y]$ by
+$\begin{array}[]{cc}F_{m}\coloneqq(1-\mu)f_{m}-\mu Yg_{m},&G_{m}\coloneqq Xf_{m}+(1-\mu)g_{m}.\end{array}$
+Then, $\pi_{n+2}(c^{m}_{n+2,n})=-\mu Y-F_{m}$ and $\pi_{n+2}(c^{m}_{n+2,n+2})=U-\mu-G_{m}$ by (16).
+Lemma 3.9.
+Let us define $h_{m},k_{m}\in\mathbb{Z}[\mu][T]$ inductively on $m=0,1,\dots$ by $h_{0}=1-\mu$, $k_{0}=1$ and
+$\begin{array}[]{cc}h_{m+1}=h_{m}-T\cdot k_{m},&k_{m+1}=h_{m}+(1-T)\cdot k_{m},\end{array}$ (20)
+Then, for every $m\in\mathbb{Z}_{\geq 0}$,
+$\begin{array}[]{cc}F_{m}(X,Y)=h_{m}(XY),&G_{m}(X,Y)=X\cdot k_{m}(XY).\end{array}$ (22)
+Proof.
+We prove by induction on $m$.
+Since $F_{0}=1-\mu$ and $G_{0}=X$, (22) for $m=0$ is satisfied by setting $h_{0}\coloneqq 1-\mu$ and $k_{0}\coloneqq 1$.
+Suppose that (22) holds for some $m\in\mathbb{Z}_{\geq 0}$.
+Using (18), we compute that
+$F_{m+1}$ $=(1-\mu)f_{m+1}-\mu Yg_{m+1}$
+$=(1-\mu)((1-XY)f_{m}-Yg_{m})-\mu Y(Xf_{m}+g_{m})$
+$=(1-\mu-XY)f_{m}-Yg_{m}$
+$=((1-\mu)f_{m}-\mu Yg_{m})-Y(Xf_{m}+(1-\mu)g_{m})$
+$=F_{m}-YG_{m}$
+$=h_{m}(XY)-XY\cdot k_{m}(XY),$
+$G_{m+1}$ $=Xf_{m+1}+(1-\mu)g_{m+1}$
+$=X((1-XY)f_{m}-Yg_{m})+(1-\mu)(Xf_{m}+g_{m})$
+$=X(2-\mu-XY)f_{m}+(1-\mu-XY)g_{m}$
+$=X((1-\mu)f_{m}-\mu Yg_{m})+(1-XY)(Xf_{m}+(1-\mu)g_{m})$
+$=XF_{m}+(1-XY)G_{m}$
+$=X\left(h_{m}(XY)+(1-XY)\cdot k_{m}(XY)\right).$
+Therefore, if we set $h_{m+1},k_{m+1}\in\mathbb{Z}[\mu][T]$ by (20), then (22) with $m$ replaced by $m+1$ holds. ∎
+For every $m\in\mathbb{Z}_{\geq 0}$, let us define a polynomial $P_{m}\in\mathbb{Z}[\mu,U][T]\subset R[T]$ by
+$P_{m}\coloneqq(U-\mu)\cdot h_{m}+\mu T\cdot k_{m}.$ (23)
+Proposition 3.10.
+Let $\mathbb{k}$ be a field.
+For any augmentation $\varepsilon$ of $\mathcal{A}^{\mathrm{knot}}_{*}(K\#T_{(2,2m+1)})$ to $\mathbb{k}$ such that $(\varepsilon(\mu),\varepsilon(U))=(y,Z)\in(\mathbb{k}^{*})^{2}$, the polynomial $\left.P_{m}\right|_{\mu=y,U=Z}\in\mathbb{k}[T]$ has a root in $\mathbb{k}$.
+Proof.
+By definition, the ideal $\mathcal{I}_{B^{\#}_{m}}$ contains $\pi_{n+2}(c^{m}_{n+2,n})=-\mu Y-F_{m}$ and $\pi_{n+2}(c^{m}_{n+2,n+2})=U-\mu-G_{m}$. Moreover, we have an identity
+$P_{m}(XY)$ $=(U-\mu)\cdot h_{m}(XY)+\mu XY\cdot k_{m}(XY)$
+$=(U-\mu)\left(h_{m}(XY)+\mu Y\right)-\mu Y\left(U-\mu-X\cdot k_{m}(XY)\right)$
+$=(U-\mu)(\mu Y+F_{m}(X,Y))-\mu Y(U-\mu-G_{m}(X,Y)),$
+which implies that $P_{m}(XY)\in\mathcal{I}_{B^{\#}_{m}}$. Therefore, there exists a well-defined unital $R$-algebra map
+$J_{m}\colon R[T]/(P_{m})\to R[a_{i,j}\mid 1\leq i,j\leq n+2,\ i\neq j]/\mathcal{I}_{B^{\#}_{m}}=HC^{\mathrm{ab}}_{0}(K\#T_{(2,2m+1)})$
+which maps $T$ to $XY=a_{n,n+2}a_{n+2,n}$. Now the proposition follows from Lemma 3.4. ∎
 Example 3.11.
-From a braid representation $\sigma_{1}^{3}\in B_{2}$ of $3_{1}$, it is not difficult to compute $HC_{0}(3_{1})$ as outlined in [16, Exercise 3.10].
-$HC_{0}(3_{1})$ is isomorphic to a commutative $R$-algebra $R[Y]/(G_{1},G_{2})$, where
-$\begin{array}[]{cc}G_{1}(Y)\coloneqq UY^{2}-\mu UY+\lambda\mu^{3}(1-\mu),&G_{2}(Y)\coloneqq UY^{2}+\lambda\mu^{2}Y+\lambda\mu^{2}(\mu-U).\end{array}$
-See [16, Exercise 5.5]. We have identities
-$YG_{1}+\mu G_{2}$ $=UY^{3}+\lambda\mu^{3}(2-\mu)Y+\lambda\mu^{3}(\mu-U),$
-$UY(YG_{1}+\mu G_{2})$ $=(UY^{2})^{2}+\lambda\mu^{3}(2-\mu)(UY^{2})+\lambda\mu^{3}(\mu-U)UY,$
-$UY^{2}$ $=G_{1}+(\mu UY-\lambda\mu^{3}(1-\mu)),$
-so the ideal $(G_{1},G_{2})$ has an element
-$(\mu UY-\lambda\mu^{3}(1-\mu))^{2}+\lambda\mu^{3}(2-\mu)(\mu UY-\lambda\mu^{3}(1-\mu))+\lambda\mu^{3}(\mu-U)UY$
-=\ = $(\mu UY)^{2}+(\lambda\mu^{4}+\lambda\mu^{3}-\lambda\mu^{2}U)(\mu UY)+(\lambda\mu^{3})^{2}(\mu-1).$
-This means that a unital $R$-algebra map
-$R[X]/(F)\to HC^{\mathrm{ab}}_{0}(3_{1})\cong R[Y]/(G_{1},G_{2})\colon X\mapsto\lambda^{-1}\mu^{-2}UY=(\lambda\mu^{3})^{-1}(\mu UY)$
-is well-defined for the polynomial $F$ in Lemma 3.6, since
-$(\lambda\mu^{3})^{2}F(\lambda^{-1}\mu^{-2}UY)=(\mu UY)^{2}+(\lambda\mu^{3})(\mu+1-\mu^{-1}U)(\mu UY)+(\lambda\mu^{3})^{2}(\mu-1)\in(G_{1},G_{2}).$
+Let us check the existence of a map as $J_{m}$ in a simple case. We have $h_{1}=1-\mu-T$, $k_{1}=(1-\mu)+(1-T)=2-\mu-T$ and
+$P_{1}(T)$ $=(U-\mu)(1-\mu-T)+\mu T(2-\mu-T)$
+$=-\mu T^{2}+(3\mu-\mu^{2}-U)T+(U-\mu)(1-\mu).$
+From a braid representation $(\sigma_{1})^{3}\in B_{2}$ of $T_{(2,3)}$, it is not difficult to compute $HC_{0}(T_{(2,3)})$ as outlined in [15, Exercise 3.10].
+$HC_{0}(T_{(2,3)})$ is isomorphic to a commutative $R$-algebra $R[X]/(G_{1},G_{2})$, where
+$\begin{array}[]{cc}G_{1}(X)\coloneqq UX^{2}-\mu UX+\lambda\mu^{3}(1-\mu),&G_{2}(X)\coloneqq UX^{2}+\lambda\mu^{2}X+\lambda\mu^{2}(\mu-U).\end{array}$
+See [15, Exercise 5.5]. We have identities
+$XG_{1}+\mu G_{2}$ $=UX^{3}+\lambda\mu^{3}(2-\mu)X+\lambda\mu^{3}(\mu-U),$
+$UX(XG_{1}+\mu G_{2})$ $=(UX^{2})^{2}+\lambda\mu^{3}(2-\mu)(UX^{2})+\lambda\mu^{3}(\mu-U)UX,$
+$\lambda\mu^{2}(\mu-U)G_{1}$ $=\lambda\mu^{2}(\mu-U)(UX^{2})-\lambda\mu^{3}(\mu-U)UX+\lambda^{2}\mu^{5}(1-\mu)(\mu-U),$
+and
+$UX(XG_{1}+\mu G_{2})+\lambda\mu^{2}(\mu-U)G_{1}$
+$=\$$(UX^{2})^{2}+\lambda\mu^{2}(3\mu-\mu^{2}-U)(UX^{2})+\lambda^{2}\mu^{5}(1-\mu)(\mu-U)$
+[…]
+Therefore, we have a well-defined unital $R$-algebra map
+$R[T]/(P_{1})\to HC^{\mathrm{ab}}_{0}(T_{(2,3)})\cong R[X]/(G_{1},G_{2})\colon T\mapsto-\lambda^{-1}\mu^{-3}UX^{2}.$
 Remark 3.12.
-If we reduce the coefficient ring $R$ to $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$ by setting $U=1$, Lemma 3.6 implies the existence of a unital $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra map
-$\mathbb{Z}[\lambda^{\pm},\mu^{\pm}][X]/(\left.F\right|_{U=1})\to\left.HC^{\mathrm{ab}}_{0}(K\#3_{1})\right|_{U=1}$
-induced by $J$.
-(Here, the notation $\left.HC_{0}(K\#3_{1})\right|_{U=1}$ for an $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra is the same one as in [16, Section 4].) However, this assertion is trivial.
-Indeed,
-$\left.F\right|_{U=1}(X)=X^{2}+(\mu+1-\mu^{-1})X+(\mu-1)=(X+\mu)(X+1-\mu^{-1})$
-and thus, for any unital $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra $A$, there exists a unital $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra map $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}][X]/(\left.F\right|_{U=1})\to A\colon X\mapsto-\mu$.
-Lemma 3.6 implies the following constraint on augmentations.
+Consider the reduction of the coefficient ring from $R$ to $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$ by setting $U=1$.
+By the substitution $U=1$, we have $\left.P_{0}\right|_{U=1}=(1-\mu)^{2}+\mu T$, and if we define $h^{\prime}_{m},k^{\prime}_{m}\in\mathbb{Z}[\mu^{\pm}]$ by $h^{\prime}_{m}\coloneqq\left.h_{m}\right|_{T=-\mu^{-1}(1-\mu)^{2}}$ and $k^{\prime}_{m}\coloneqq\left.k_{m}\right|_{T=-\mu^{-1}(1-\mu)^{2}}$, then
+$\left.P_{m+1}\right|_{U=1,T=-\mu^{-1}(1-\mu)^{2}}$ $=(1-\mu)\cdot h^{\prime}_{m+1}+\mu(-\mu^{-1}(1-\mu)^{2})\cdot k^{\prime}_{m+1}$
+$=(1-\mu)(h^{\prime}_{m}+\mu^{-1}(1-\mu)^{2}k^{\prime}_{m})-(1-\mu)^{2}(h^{\prime}_{m}+(1+\mu^{-1}(1-\mu)^{2})k^{\prime}_{m})$
+$=\mu\left((1-\mu)h^{\prime}_{m})-(1-\mu)^{2}k^{\prime}_{m}\right)$
+$=\mu\left.P_{m}\right|_{U=1,T=-\mu^{-1}(1-\mu)^{2}}.$
+Here, the second equality follows from (20).
+By induction on $m=0,1,\dots$, we can see that the polynomial $\left.P_{m}\right|_{U=1}\in\mathbb{Z}[\mu^{\pm}][T]$ has a root $T=-\mu^{-1}(1-\mu)^{2}$.
+Hence, Proposition 3.10 is trivial for if an augmentation $\varepsilon$ satisfies $\varepsilon(U)=1$.
+3.2.3 Application of $\mathbb{Q}$-valued augmentations
+Notation. For every $q\in\mathbb{C}(T)$, we define its degree $\deg q\coloneqq\max\{\deg f,\deg g\}$, where $f,g\in\mathbb{C}[T]$ are coprime polynomials such that $g\neq 0$ and $q=\frac{f}{g}$.
+Remark 3.13.
+Let us take $q=\frac{f}{g}\in\mathbb{C}(T)$ as above.
+Then, $\deg q\in\mathbb{Z}_{\geq 0}$ coincides with the degree of the continuous map $\hat{q}\colon\mathbb{C}P^{1}\to\mathbb{C}P^{1}$ such that $\hat{q}([z:1])=[f(z):g(z)]$ for every $z\in\mathbb{C}$ and
+$\hat{q}([1:0])=\begin{cases}[\lim_{z\to\infty}\frac{f(z)}{g(z)}:1]&\text{ if }\deg f\leq\deg g,\\ [1:0]&\text{ if }\deg f>\deg g.\end{cases}$
+Consider $r=\frac{f^{\prime}}{g^{\prime}}\in\mathbb{C}(T^{\prime})$ for $f^{\prime},g^{\prime}\in\mathbb{C}(T^{\prime})$ such that $\left.g^{\prime}\right|_{T^{\prime}=q}\neq 0$ in $\mathbb{C}(T)$.
+Then, the degree of $\left.r\right|_{T^{\prime}=q}\in\mathbb{C}(T)$ is computed by $\deg r\cdot\deg q$, which is the degree of the composite map $\hat{r}\circ\hat{q}\colon\mathbb{C}P^{1}\to\mathbb{C}P^{1}$.
+In the following, we fix $y_{0}\in\mathbb{Q}^{*}\setminus\{1\}$ (one may choose, for instance, $y_{0}=2$). By the substitution $\mu=y_{0}$, let us define $\bar{h}_{m},\bar{k}_{m}\in\mathbb{Q}[T]$ as
+$\begin{array}[]{cc}\bar{h}_{m}\coloneqq\left.h_{m}\right|_{\mu=y_{0}},&\bar{k}_{m}\coloneqq\left.k_{m}\right|_{\mu=y_{0}}.\end{array}$
 Lemma 3.14.
-Let $\mathbb{k}$ be a field. For any augmentation $\varepsilon$ of $(\mathcal{A}^{\mathrm{knot}}_{*}(K\#3_{1}),\partial_{K\#3_{1}})$ to $\mathbb{k}$,
-$D_{\varepsilon}\coloneqq(\varepsilon(\mu)+1-\varepsilon(\mu)^{-1}\varepsilon(U))^{2}-4(\varepsilon(\mu)-1)$
-satisfies $D_{\varepsilon}\in\{k^{2}\mid k\in\mathbb{k}\}$.
+For every $m\in\mathbb{Z}_{\geq 1}$, $\deg\bar{h}_{m}=\deg\bar{k}_{m}=m$, and $\bar{h}_{m}$ and $T\cdot\bar{k}_{m}$ are coprime polynomials in $\mathbb{C}[T]$. In particular, if we define
+$r_{m}\coloneqq\frac{T\cdot\bar{k}_{m}}{\bar{h}_{m}}\in\mathbb{C}(T),$
+then $\deg r_{m}=\deg(T\cdot\bar{k}_{m})=m+1$ for every $m\in\mathbb{Z}_{\geq 1}$.
 Proof.
-Since $HC_{0}(K\#3_{1})=\mathcal{A}^{\mathrm{knot}}_{0}(K\#3_{1})/\partial_{K\#3_{1}}(\mathcal{A}^{\mathrm{knot}}_{1}(K\#3_{1}))$, any augmentation $\varepsilon$ to $\mathbb{k}$ of $(\mathcal{A}^{\mathrm{knot}}_{*}(K\#3_{1}),\partial_{K\#3_{1}})$ induces a ring homomorphism
-$\bar{\varepsilon}\colon HC^{\mathrm{ab}}_{0}(K\#3_{1})\to\mathbb{k}$
-such that $\bar{\varepsilon}(r)=\varepsilon(r)$ for every $r\in R$. Composing with $J$ in Lemma 3.6, we obtain a map
-$\bar{\varepsilon}\circ J\colon R[X]/(F)\to\mathbb{k}.$
-Let us denote $\zeta=\bar{\varepsilon}\circ J(X)$. Since $\bar{\varepsilon}\circ J(F)=0$ and $J$ is a unital $R$-algebra map, we obtain an equation in $\mathbb{k}$
-$\zeta^{2}+(\varepsilon(\mu)+1-\varepsilon(\mu)^{-1}\varepsilon(U))\zeta+(\varepsilon(\mu)-1)=0.$
-Then, $D_{\varepsilon}$ in the lemma is the discriminant of this quadratic equation for $\zeta$.
-In order that this equation has a solution in $\mathbb{k}$, it is necessary that $D_{\varepsilon}$ is the square of some element of $\mathbb{k}$.
+By Lemma 3.9, we have $\bar{h}_{0}=1-y_{0}$, $\bar{k}_{0}=1$ and
+$\begin{array}[]{cc}\bar{h}_{m+1}=\bar{h}_{m}-T\cdot\bar{k}_{m},&\bar{k}_{m+1}=\bar{h}_{m}+(1-T)\cdot\bar{k}_{m},\end{array}$
+for every $m\in\mathbb{Z}_{\geq 0}$. Then, it is easy to see that:
+  • For every $m\in\mathbb{Z}_{\geq 1}$, both $\bar{h}_{m}$ and $\bar{k}_{m}$ have $(-1)^{m}\cdot T^{m}$ as the term with the highest degree.
+  • For every $m\in\mathbb{Z}_{\geq 0}$, $\bar{h}_{m}=(1-T)\cdot\bar{h}_{m+1}+T\cdot\bar{k}_{m+1}$ and $T\cdot\bar{k}_{m}=T\cdot(\bar{k}_{m+1}-\bar{h}_{m+1})$.
+From the first assertion, $\deg\bar{k}_{m}=\deg\bar{h}_{m}=m$ for every $m\in\mathbb{Z}_{\geq 1}$ From the second assertion, if we suppose that $\bar{h}_{m}$ and $T\cdot\bar{k}_{m}$ have a common root $T=c$ in $\mathbb{C}$ for some $m\in\mathbb{Z}_{\geq 1}$, then $\bar{h}_{m^{\prime}}(c)=c\cdot\bar{k}_{m^{\prime}}(c)=0$ for every $m^{\prime}\in\{0,\dots,m\}$, which is a contradiction when $m^{\prime}=0$ since $1-y_{0}\neq 0$.
+Thus, $\bar{h}_{m}$ and $T\cdot\bar{k}_{m}$ are coprime in $\mathbb{C}[T]$ for every $m\in\mathbb{Z}_{\geq 1}$.
 ∎
-Combining Theorem 3.2, Lemma 3.9 and a computation of the knot DGA of the unknot, we obtain a constraint on the isotopy class of clean Lagrangian intersections.
+Combining Proposition 3.5 and Proposition 3.10 with $\mathbb{k}=\mathbb{Q}$, we obtain the following theorem.
 Theorem 3.15.
-Let $K_{0}$ be the connected sum of an arbitrary knot $K$ in $\mathbb{R}^{3}$ with $3_{1}$ or its mirror $\overline{3_{1}}$.
+Let $K_{0}=K\#T_{(2,2m+1)}$, where $K$ is an arbitrary knot in $\mathbb{R}^{3}$ and $m\in\mathbb{Z}\setminus\{0,-1\}$.
 Then, there is no $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along the unknot in $\mathbb{R}^{3}$.
 Proof.
-Assume that there exists $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection $K_{1}=\varphi(L_{K_{0}})\cap\mathbb{R}^{3}$ and $K_{1}$ is the unknot in $\mathbb{R}^{3}$.
-In the case of $K_{0}=K\#3_{1}$, we can deduce a contradiction as follows: From the trivial braid on a single strand, we can compute as in [16, Example 3.13] that $\mathcal{A}^{\mathrm{knot}}_{0}(K_{1})=R$ and $\partial_{K_{1}}(\mathcal{A}_{1}^{\mathrm{knot}}(K_{1}))$ is the ideal generated by
-$U-\lambda-\mu+\lambda\mu.$
-Let us choose an augmentation $\varepsilon_{1}\colon\mathcal{A}^{\mathrm{knot}}_{*}(K_{1})\to\mathbb{Q}$ determined by
-$\begin{array}[]{ccc}\varepsilon_{1}(\lambda)=3,&\varepsilon_{1}(\mu)=2,&\varepsilon_{1}(U)=-1.\end{array}$
-Composing with the map $\Psi_{\varphi}$ in Theorem 3.2, we obtain $\varepsilon\coloneqq\varepsilon_{1}\circ\Psi_{\varphi}\colon\mathcal{A}^{\mathrm{knot}}_{*}(K_{0})\to\mathbb{Q}$.
-It is an augmentation of the knot DGA of $K_{0}=K\#3_{1}$ to $\mathbb{Q}$ such that
-$\begin{array}[]{ccc}\varepsilon(\lambda)=3^{a}2^{b}(-1)^{c},&\varepsilon(\mu)=2^{a}(-1)^{d},&\varepsilon(U)=-1.\end{array}$
-We apply Lemma 3.9 for $\mathbb{k}=\mathbb{Q}$. Then,
-$D_{\varepsilon}=(\varepsilon(\mu)+1+\varepsilon(\mu)^{-1})^{2}-4(\varepsilon(\mu)-1)$
-must be the square of some rational number. However, we can calculate as follows:
-$D_{\varepsilon}=\begin{cases}(\frac{7}{2})^{2}-4\cdot 1=\frac{33}{4}&\text{ if}\varepsilon(\mu)=2,\\ (-\frac{3}{2})^{2}-4\cdot(-3)=\frac{57}{4}&\text{ if }\varepsilon(\mu)=-2,\\ (\frac{7}{2})^{2}-4\cdot(-\frac{1}{2})=\frac{57}{4}&\text{ if }\varepsilon(\mu)=\frac{1}{2},\\ (-\frac{3}{2})^{2}-4\cdot(-\frac{3}{2})=\frac{33}{4}&\text{ if }\varepsilon(\mu)=-\frac{1}{2},\end{cases}$
-and this is a contradiction.
-In the case of $K_{0}=K\#\overline{3_{1}}$, we can deduce a contradiction as follows: Take a Hamiltonian $H\colon T^{*}\mathbb{R}^{3}\times[0,1]\to\mathbb{R}$ such that $\varphi=\varphi^{1}_{H}$.
-Consider an involution
+For a knot $K_{0}$, suppose that there exists $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection $K_{1}=\varphi(L_{K_{0}})\cap\mathbb{R}^{3}$ and $K_{1}$ is the unknot in $\mathbb{R}^{3}$.
+Assuming that $K_{0}=K\#T_{(2,2m+1)}$, we will deduce contradictions in two cases: the case $m\geq 1$ and the case $m\leq-2$.
+(The case $m\geq 1$) By Proposition 3.5, for every $Z_{0}\in\{r\in\mathbb{Q}^{*}\mid r^{n}\neq y_{0}\text{ for every }n\in\mathbb{Z}\}$, there exists an augmentation $\varepsilon$ of $(\mathcal{A}^{\mathrm{knot}}_{*}(K_{0}),\partial_{K_{0}})$ to $\mathbb{Q}$ such that $\varepsilon(\mu)=y_{0}$ and $\varepsilon(U)=Z_{0}$.
+Then, it follows from Proposition 3.10 that $\left.P_{m}\right|_{\mu=y_{0},U=Z_{0}}\in\mathbb{Q}[T]$ has a root in $\mathbb{Q}$.
+From the definition (23) of $P_{m}$, it is equivalent to that the equation for $T$
+$(Z_{0}-y_{0})\cdot\bar{h}_{m}(T)+y_{0}T\cdot\bar{k}_{m}(T)=0$ (24)
+has a solution in $\mathbb{Q}$.
+As a polynomial in $\mathbb{Q}(Z)[T]$, let us consider $(Z-y_{0})\cdot\bar{h}_{m}+y_{0}T\cdot\bar{k}_{m}$. Let $p_{1},\dots,p_{e}\in\mathbb{Q}(Z)[T]$ be irreducible elements such that
+$(Z-y_{0})\cdot\bar{h}_{m}+y_{0}T\cdot\bar{k}_{m}=p_{1}\cdots p_{e}.$
+We claim that the degree of $p_{i}$ as a polynomial in $T$ is greater than or equal to $2$ for every $i\in\{1,\dots,e\}$.
+It is proved by a contradiction as follows: Assume that there exists $i\in\{1,\dots,e\}$ such that $p_{i}=q^{\prime}(T-q)$ for some $q,q^{\prime}\in\mathbb{Q}(Z)$ with $q^{\prime}\neq 0$.
+Then, we have an equation
+$(y_{0}^{-1}Z-1)\cdot\left.\bar{h}_{m}\right|_{T=q}+q\cdot\left.\bar{k}_{m}\right|_{T=q}=0$
+in $\mathbb{C}(Z)$.
+Since $\bar{h}_{m}$ and $T\cdot\bar{k}_{m}$ does not have a common root in $\mathbb{C}$ by Lemma 3.14, $q$ cannot be a constant and thus, $\left.\bar{h}_{m}\right|_{T=q}\neq 0$ in $\mathbb{C}(Z)$.
+Therefore, we obtain an equation in $\mathbb{C}(Z)$
+$\left.r_{m}\right|_{T=q}=1-y_{0}^{-1}Z$
+for $r_{m}=(T\cdot\bar{k}_{m})/\bar{h}_{m}\in\mathbb{C}(T)$. We can compute that
+${\deg r_{m}\cdot\deg q=\deg\left(\left.r_{m}\right|_{T=q}\right)=\deg(1-y_{0}^{-1}Z)=1.}$
+For the first equality, see Remark 3.13. However, $\deg r_{m}\geq 2$ by Lemma 3.14, and we get a contradiction.
+By Hilbert’s irreduciblity theorem (see e.g.
+[11, Chapter 9, Corollary 2.4], which shows that $\mathbb{Q}$ is a Hilbertian field), there are infinitely many $Z_{0}\in\mathbb{Q}$ such that all of the specializations
+$\left.p_{1}\right|_{Z=Z_{0}},\dots,\left.p_{e}\right|_{Z=Z_{0}}\in\mathbb{Q}[T]$
+are irreducible in $\mathbb{Q}[T]$.
+We note that the set $\bigcup_{i=1}^{e}\{Z_{0}\in\mathbb{Q}\mid\deg(\left.p_{i}\right|_{Z=Z_{0}})\leq 1\}$ is finite from the above observation of the degree of $p_{1},\dots,p_{m}$ in $T$.
+In addition, for the fixed rational number $y_{0}\neq 1$, the set $\{r\in\mathbb{Q}^{*}\mid r^{n}=y_{0}\text{ for some }n\in\mathbb{Z}\}$ is also finite.
+As a consequence, there are infinitely many $Z_{0}\in\mathbb{Q}^{*}\setminus\{r\in\mathbb{Q}^{*}\mid r^{n}=y_{0}\text{ for some }n\in\mathbb{Z}\}$ such that $\left.p_{i}\right|_{Z=Z_{0}}\in\mathbb{Q}[T]$ is irreducible and $\deg(\left.p_{i}\right|_{Z=Z_{0}})\geq 2$ for every $i\in\{1,\dots,e\}$, which means that the equation (24) has no solution in $\mathbb{Q}$.
+This contradicts Proposition 3.5 and Proposition 3.10.
+(The case $m\leq-2$) Take a Hamiltonian $H\colon T^{*}\mathbb{R}^{3}\times[0,1]\to\mathbb{R}$ such that $\varphi=\varphi^{1}_{H}$. Consider an involution
 $\iota\colon T^{*}\mathbb{R}^{3}\to T^{*}\mathbb{R}^{3}\colon(q,p)=((q_{1},q_{2},q_{3}),(p_{1},p_{2},p_{3}))\mapsto((q_{1},q_{2},-q_{3}),(p_{1},p_{2},-p_{3}))$
@@
 $\overline{K_{i}}\coloneqq\{(q_{1},q_{2},-q_{3})\in\mathbb{R}^{3}\mid(q_{1},q_{2},q_{3})\in K_{i}\}\text{ for }i=0,1,$
-$\varphi^{1}_{\overline{H}}(L_{\overline{K_{0}}})$ and $\mathbb{R}^{3}=\iota(\mathbb{R}^{3})$ intersect cleanly along $\overline{K_{1}}$.
-Since $\overline{K_{1}}$ is the unknot and $\overline{K_{0}}$ is the connected sum of the mirror of $K$ and $3_{1}$, this contradicts to the result in the previous case.
+$\varphi^{1}_{\overline{H}}(L_{\overline{K_{0}}})$ intersects $\mathbb{R}^{3}=\iota(\mathbb{R}^{3})$ cleanly along $\overline{K_{1}}$.
+Since $\overline{K_{1}}$ is the unknot and $\overline{K_{0}}$ is the connected sum of the mirror of $K$ and $T_{(2,-2m-1)}$, this contradicts to the result from the above case.
 ∎
 Remark 3.16.
-For the proof of the theorem, we may choose other field $\mathbb{k}$ and other augmentation $\varepsilon_{1}$ of the knot DGA of $K\#3_{1}$ to $\mathbb{k}$ such that $\varepsilon(U)\neq 1$.
-Let us see two examples.
-  1. 1.
-    Choose $\mathbb{k}=\mathbb{Q}[\sqrt{-1}]$ and $\varepsilon_{1}$ determined by
-    $\begin{array}[]{ccc}\varepsilon_{1}(\lambda)=2-\sqrt{-1},&\varepsilon_{1}(\mu)=2,&\varepsilon_{1}(U)=\sqrt{-1},\end{array}$
-    in $\mathbb{Q}[\sqrt{-1}]$. For $\varepsilon\coloneqq\varepsilon_{1}\circ\Psi_{\varphi}$, let us write $D_{\varepsilon}=D_{y}$ if $\varepsilon(\mu)=y$. Then,
-    $D_{y}=(y+1-\sqrt{-1}y^{-1})^{2}-4(y-1)=y^{2}-2y-2\sqrt{-1}y^{-1}-y^{-2}+(5-2\sqrt{-1}),$
-    and $y$ varies in the set {±$\{\pm 2,\pm 2\sqrt{-1},\pm\frac{1}{2},\pm\frac{\sqrt{-1}}{2}\}$. Noting that $D_{\sqrt{-1}y^{-1}}=D_{y}$, we can compute that $D_{y}$ takes its value in the set
-    ${S\coloneqq\left\{\frac{19}{4}-3\sqrt{-1},\frac{51}{4}-\sqrt{-1},\frac{1}{4}-6\sqrt{-1},\frac{9}{4}+2\sqrt{-1}\right\}.}$
-    In general, suppose that $D=a+\sqrt{-1}b$ for $a,b\in\mathbb{Q}$ satisfies $D\in\{k^{2}\mid k\in\mathbb{Q}[\sqrt{-1}]\}$.
-Then, $a^{2}+b^{2}=D\cdot\bar{D}\in\{r^{2}\mid r\in\mathbb{Q}\}$, where $\bar{D}$ is the complex conjugate of $D$.
-However, for any $D=a+\sqrt{-1}b\in S$, we can check that $a^{2}+b^{2}\notin\{r^{2}\mid r\in\mathbb{Q}\}$.
-Therefore, $D_{y}\notin\{k^{2}\mid k\in\mathbb{Q}[\sqrt{-1}]\}$ for any $y\in\{\pm 2,\pm 2\sqrt{-1},\pm\frac{1}{2},\pm\frac{\sqrt{-1}}{2}\}$.
-  2. 2.
-    Choose $\mathbb{k}=\mathbb{F}_{5}\cong\mathbb{Z}/5$ and $\varepsilon_{1}$ determined by
-    $\begin{array}[]{ccc}\varepsilon_{1}(\lambda)=3,&\varepsilon_{1}(\mu)=2,&\varepsilon_{1}(U)=-1,\end{array}$
-    in $\mathbb{F}_{5}$, similar to the one in the proof of Theorem 3.10.
-In general, if $D\in\mathbb{Z}/5$ satisfies $D\equiv k^{2}\mod 5$ for some $k\in\mathbb{Z}$, then $D\equiv 0,1$ or $4\mod 5$.
-However, for $\varepsilon\coloneqq\varepsilon_{1}\circ\Psi_{\varphi}$,
-    $D_{\varepsilon}\equiv 33\cdot 4\equiv 2\text{ or }D_{\varepsilon}\equiv 57\cdot 4\equiv 3\mod 5.$
-    Therefore, $D_{\varepsilon}\notin\{k^{2}\mid k\in\mathbb{F}_{5}\}$.
-Remark 3.12.
-To the author’s knowledge, it seems difficult to show Theorem 3.10, even when $K_{0}=3_{1}$, by only using the knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$.
-Indeed, for the trefoil knot $3_{1}$ and the unknot $O$,
-$\left.HC_{0}(3_{1})\right|_{U=1}$ $\cong\mathbb{Z}[\lambda^{\pm},\mu^{\pm}][Y]/(Y^{2}-\mu Y+\lambda\mu^{3}(1-\mu),\ Y^{2}+\lambda\mu^{2}Y+\lambda\mu^{2}(\mu-1)),$
+To the author’s knowledge, it seems difficult to show Theorem 3.15, even when $K_{0}=T_{(2,3)}$, by only using the knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$.
+Indeed, for the trefoil knot $T_{(2,3)}$ and the unknot $O$,
+$\left.HC_{0}(T_{(2,3)})\right|_{U=1}$ $\cong\mathbb{Z}[\lambda^{\pm},\mu^{\pm}][X]/(X^{2}-\mu X+\lambda\mu^{3}(1-\mu),\ X^{2}+\lambda\mu^{2}X+\lambda\mu^{2}(\mu-1)),$
 $\left.HC_{0}(O)\right|_{U=1}$ $\cong\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]/((1-\lambda)(1-\mu)),$
 and we can find a unital $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra map
-$\left.HC_{0}(3_{1})\right|_{U=1}\to\left.HC_{0}(O)\right|_{U=1}\colon Y\mapsto\mu-\mu^{2}.$
-Therefore, we cannot deduce a contradiction from the existence of a unital $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra map which is induced on the 00-the degree part of the homology from the DGA map as in Theorem 3.2.
+$\left.HC_{0}(T_{(2,3)})\right|_{U=1}\to\left.HC_{0}(O)\right|_{U=1}\colon X\mapsto\mu-\mu^{2}.$
+Therefore, we cannot deduce a contradiction from the existence of a unital $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra map between the 0-th homology which is induced from the DGA map as in Theorem 3.2.
+3.3 Proof of Theorem 1.1 for $K^{\prime}\#4_{1}$
+Fix any $B\in B_{n}$ whose closure gives a knot $K$. We consider a braid on $(n+3)$ strands
+$B^{\#}\coloneqq\sigma_{n}\left(\sigma_{n+1}(\sigma_{n+2})^{-1}\sigma_{n+1}(\sigma_{n+2})^{-1}\right)B\in B_{n+3}.$
+Then, the closure of $B^{\#}$ is the connected sum of $K$ and the figure-eight knot $4_{1}$.
+Proposition 3.17.
+We define $P\in\mathbb{Z}[\mu^{\pm},U^{\pm}][T]$ by
+$\begin{split}P(T)\coloneqq&-(1-\mu)(U-\mu)+\left(-2+2\mu+\mu^{2}+(\mu^{-1}-1-\mu)U-\mu^{3}U^{-1}\right)T\\ &+(-2\mu^{2}+\mu U)T^{2}-\mu UT^{3}.\end{split}$ (25)
+Let $\mathbb{k}$ be a field.
+Then, for any augmentation $\varepsilon$ of $\mathcal{A}^{\mathrm{knot}}_{*}(K\#4_{1})$ to $\mathbb{k}$ such that $(\varepsilon(\mu),\varepsilon(U))=(y,Z)$, the polynomial $\left.P\right|_{\mu=y,U=Z}\in\mathbb{k}[T]$ has a root in $\mathbb{k}$.
+Proof.
+First, let us compute $\phi_{B^{\#}}(a_{n+2,*})$ and $\phi_{B^{\#}}(a_{n+3,*})$.
+Since $\phi_{\sigma_{k}}(a_{n+2,*})=a_{n+2,*}$ and $\phi_{\sigma_{k}}(a_{n+3,*})=a_{n+3,*}$ for every $k\in\{1,\dots,n-1\}$, we have $\phi_{B}(a_{n+2,*})=a_{n+2,*}$ and $\phi_{B}(a_{n+3,*})=a_{n+3,*}$.
+Referring to (10), we compute that
+$(\phi_{\sigma_{n+2}})^{-1}\colon\begin{cases}a_{n+1,*}\mapsto a_{n+1,*},\\ a_{n+2,*}\mapsto a_{n+3,*},\\ a_{n+3,*}\mapsto a_{n+2,*}-a_{n+2,n+3}a_{n+3,*},\\ a_{n+1,n+3}\mapsto a_{n+1,n+2}-a_{n+1,n+3}a_{n+3,n+2},\end{cases}$
+and
+${a_{n+2,*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\phi_{\sigma_{n+2}})^{-1}}$${a_{n+3,*}}$${a_{n+3,*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\phi_{\sigma_{n+2}})^{-1}}$${a_{n+2,*}-a_{n+2,n+3}a_{n+3,*}}$$\scriptstyle{\phi_{\sigma_{n+1}}}$${a_{n+3,*}}$$\scriptstyle{\phi_{\sigma_{n+1}}}$${a_{n+1,*}-a_{n+1,n+3}a_{n+3,*}}$$\scriptstyle{(\phi_{\sigma_{n+2}})^{-1}}$${a_{n+2,*}-a_{n+2,n+3}a_{n+3,*}}$$\scriptstyle{(\phi_{\sigma_{n+2}})^{-1}}$${a_{n+1,*}-M^{\prime\prime}(a_{n+2,*}-a_{n+2,n+3}a_{n+3,*})}$$\scriptstyle{\phi_{\sigma_{n+1}}}$${a_{n+1,*}-a_{n+1,n+3}a_{n+3,*}}$$\scriptstyle{\phi_{\sigma_{n+1}}}$${\begin{array}[]{l}a_{n+2,*}-a_{n+2,n+1}a_{n+1,*}\\ -M^{\prime}(a_{n+1,*}-a_{n+1,n+3}a_{n+3,*})\end{array}}$$\scriptstyle{\phi_{\sigma_{n}}}$${a_{n,*}-a_{n,n+3}a_{n+3,*},}$$\scriptstyle{\phi_{\sigma_{n}}}$${a_{n+2,*}-a_{n+2,n}a_{n,*}-M(a_{n,*}-a_{n,n+3}a_{n+3,*}),}$
+where
+$M^{\prime\prime}$ $\coloneqq(\phi_{\sigma_{n+2}})^{-1}(a_{n+1,n+3})=a_{n+1,n+2}-a_{n+1,n+3}a_{n+3,n+2},$
+$M^{\prime}$ $\coloneqq\phi_{\sigma_{n+1}}(M^{\prime\prime})=-a_{n+2,n+1}-(a_{n+2,n+3}-a_{n+2,n+1}a_{n+1,n+3})a_{n+3,n+1},$
+$M$ $\coloneqq\phi_{\sigma_{n}}(M^{\prime})=-a_{n+2,n}-(a_{n+2,n+3}-a_{n+2,n}a_{n,n+3})a_{n+3,n}.$
+As a consequence, we obtain
+$\phi_{B^{\#}}(a_{n+2,*})$ $=a_{n,*}-a_{n,n+3}a_{n+3,*},$
+$\phi_{B^{\#}}(a_{n+3,*})$ $=(-a_{n+2,n}-M)a_{n,*}+a_{n+2,*}+(Ma_{n,n+3})a_{n+3,*},$
+for
+$M=-a_{n+2,n}(1-a_{n,n+3}a_{n+3,n})-a_{n+2,n+3}a_{n+3,n}.$ (26)
+The $(n+2)$-th and $(n+3)$-th row of $\bm{\Phi}^{L}_{B^{\#}}$ are given by
+$\begin{pmatrix}0&\cdots&0&1&0&0&-a_{n,n+3}\\ 0&\cdots&0&-a_{n+2,n}-M&0&1&Ma_{n,n+3}\end{pmatrix}$.
+Let $c_{i,j}$ denote the $(i,j)$-th entry of $\widehat{\mathbb{A}}-\bm{\Lambda}\cdot\bm{\Phi}^{L}_{B^{\#}}\cdot\mathbb{A}$. Then,
+$c_{n+2,n}$ $=\widehat{\mathbb{A}}_{n+2,n}-(1\cdot\mathbb{A}_{n,n}-a_{n,n+3}\mathbb{A}_{n+3,n})$
+$=-\mu a_{n+2,n}-(1-\mu)-\mu a_{n,n+3}a_{n+3,n},$
+$c_{n+2,n+3}$ $=\widehat{\mathbb{A}}_{n+2,n+3}-(1\cdot\mathbb{A}_{n,n+3}-a_{n,n+3}\mathbb{A}_{n+3,n+3})$
+$=Ua_{n+2,n+3}-a_{n,n+3}+(1-\mu)a_{n,n+3}$
+$=Ua_{n+2,n+3}-\mu a_{n,n+3},$
+$c_{n+3,n}$ $=\widehat{\mathbb{A}}_{n+3,n}-((-a_{n+2,n}-M)\mathbb{A}_{n,n}+1\cdot\mathbb{A}_{n+2,n}+Ma_{n,n+3}\mathbb{A}_{n+3,n})$
+$=-\mu a_{n+3,n}+(1-\mu)(a_{n+2,n}+M)+\mu a_{n+2,n}+\mu Ma_{n,n+3}a_{n+3,n}$
+$=-\mu a_{n+3,n}+a_{n+2,n}+M(1-\mu+\mu a_{n,n+3}a_{n+3,n}),$
+$c_{n+3,n+3}$ $=\widehat{\mathbb{A}}_{n+3,n+3}-((-a_{n+2,n}-M)\mathbb{A}_{n,n+3}+1\cdot\mathbb{A}_{n+2,n+3}+Ma_{n,n+3}\mathbb{A}_{n+3,n+3})$
+$=(U-\mu)+(a_{n+2,n}+M)a_{n,n+3}-a_{n+2,n+3}-(1-\mu)Ma_{n,n+3}$
+$=(U-\mu)+a_{n+2,n}a_{n,n+3}-a_{n+2,n+3}+\mu Ma_{n,n+3}.$
+Let us denote $X=a_{n,n+3}$ and $Y=a_{n+3,n}$.
+The ideal $\mathcal{I}_{B^{\#}}$ contains $\pi_{n+3}(c_{i,j})$ for every $i,j\in\{1,\dots,n+3\}$, where $\pi_{n+3}$ is defined as (13).
+In particular, $\mathcal{I}_{B^{\#}}$ contains
+$\mu^{-1}\cdot\pi_{n+3}(c_{n+2,n})$ $=-a_{n+2,n}-\mu^{-1}+1-XY,$
+$U^{-1}\cdot\pi_{n+3}(c_{n+2,n+3})$ $=a_{n+2,n+3}-\mu U^{-1}X.$
+Replacing $a_{n+2,n}$ with $\mu^{-1}+1-XY$ and $a_{n+2,n+3}$ with $\mu U^{-1}X$, we obtain from $\pi_{n+3}(M)$
+$\bar{M}$ $\coloneqq-(-\mu^{-1}+1-XY)(1-XY)-\mu U^{-1}XY$
+$=\mu^{-1}(1-\mu)+(\mu^{-1}-\mu U^{-1})XY-(XY)^{2},$
+and by the same replacement, we obtain from $\pi_{n+3}(c_{n+3,n})$ and $\pi_{n+3}(c_{n+3,n+3})$
+$F$ $\coloneqq-\mu Y+(-\mu^{-1}+1-XY)+\bar{M}(1-\mu+\mu XY),$
+$G$ $\coloneqq(U-\mu)+(-\mu^{-1}+1-XY)X-\mu U^{-1}X+\mu\bar{M}X$
+$=(U-\mu)+(-\mu^{-1}+1-\mu U^{-1}-XY)X+\mu\bar{M}X.$
+Then, both $F$ and $G$ are contained in $\mathcal{I}_{B^{\#}}$.
+We focus on the first terms $\mu Y$ and $(U-\mu)$ of $F$ and $G$. Then, we compute
+$(U-\mu)\cdot F+\mu Y\cdot G=$ $(U-\mu)(-\mu^{-1}+1-XY)+(U-\mu)\bar{M}(1-\mu+\mu XY)$
+$+\mu(-\mu^{-1}+1-\mu U^{-1}-XY)XY+\mu^{2}\bar{M}XY$
+$=$ $-\mu^{-1}(1-\mu)(U-\mu)+(-1+2\mu-U-\mu^{2}U^{-1})XY-\mu(XY)^{2}$
+$+\bar{M}\left((1-\mu)(U-\mu)+\mu UXY\right)$
+and
+$\bar{M}\left((1-\mu)(U-\mu)+\mu UXY\right)$
+$=$ $\left(\mu^{-1}(1-\mu)+(\mu^{-1}-\mu U^{-1})XY-(XY)^{2}\right)\left((1-\mu)(U-\mu)+\mu UXY\right)$
+$=$ $\mu^{-1}(1-\mu)^{2}(U-\mu)+(1-\mu)(U\mu^{-1}-1-\mu+U+\mu^{2}U^{-1})XY$
+$+(\mu-2\mu^{2}+\mu U)(XY)^{2}-\mu U(XY)^{3}.$
+Therefore,
+$(U-\mu)\cdot F+\mu Y\cdot G$
+$=$ $-(1-\mu)(U-\mu)+\left(-2+2\mu+\mu^{2}+(\mu^{-1}-1-\mu)U-\mu^{3}U^{-1}\right)XY$
+$+(-2\mu^{2}+\mu U)(XY)^{2}-\mu U(XY)^{3}.$
+If we define $P\in\mathbb{Z}[\mu^{\pm},U][T]$ by (25), then
+$(U-\mu)\cdot F+\mu Y\cdot G=P(XY),$
+which implies that $P(XY)$ is contained in $\mathcal{I}_{B^{\#}}$. Therefore, we have a well-defined unital $R$-algebra map
+$J\colon R[T]/(P)\to R[a_{i,j}\mid 1\leq i,j\leq n+3,\ i\neq j]/\mathcal{I}_{B^{\#}}=HC^{\mathrm{ab}}_{0}(K\#4_{1})$
+which maps $T$ to $XY=a_{n,n+3}a_{n+3,n}$. Now, the proposition follows from Lemma 3.4. ∎
+Remark 3.18.
+By the substitution $U=1$,
+$\left.P\right|_{U=1}(T)=$ $-(1-\mu)^{2}+(\mu^{-1}-3+\mu+\mu^{2}-\mu^{3})T+(-2\mu^{2}+\mu)T^{2}-\mu T^{3}$
+$=$ $-(T+\mu^{-1}(1-\mu)^{2})\left(\mu T^{2}+(-1+\mu+\mu^{2})T+\mu\right).$
+Therefore, $\left.P\right|_{U=1}\in\mathbb{Z}[\mu^{\pm}][T]$ has a root $T=-\mu^{-1}(1-\mu)^{2}$. Hence, Proposition 3.17 is trivial if an augmentation $\varepsilon$ satisfies $\varepsilon(U)=1$.
+Theorem 3.19.
+Let $K_{0}=K\#4_{1}$, where $K$ is an arbitrary knot in $\mathbb{R}^{3}$.
+Then, there is no $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along the unknot in $\mathbb{R}^{3}$.
+Proof.
+For a knot $K_{0}$, suppose that there exists $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection $K_{1}=\varphi(L_{K_{0}})\cap\mathbb{R}^{3}$ and $K_{1}$ is the unknot in $\mathbb{R}^{3}$.
+Assuming that $K_{0}=K\#4_{1}$, we deduce a contradiction as follows: By Proposition 3.5, there exists an augmentation $\varepsilon$ of $(\mathcal{A}^{\mathrm{knot}}_{*}(K_{0}),\partial_{K_{0}})$ to $\mathbb{Q}$ such that $\varepsilon(\mu)=-1$ and $\varepsilon(U)=-2$.
+Then, it follows from Proposition 3.17 that the polynomial $\bar{P}\coloneqq\left.P\right|_{\mu=-1,U=-2}\in\mathbb{Q}[T]$ has a root in $\mathbb{Q}$.
+On the other hand, from the definition of $P$,
+$\bar{P}(T)=2-\frac{3}{2}T-2T^{3}$
+Multiplying by $2$, we obtain a polynomial with integer coefficients
+$4-3T-4T^{3}.$
+By the rational root theorem, the only possible rational roots of this polynomial are of the form $\frac{p}{q}$, where $p\in\mathbb{Z}$ is a divisor of $4$ and $q\in\mathbb{Z}$ is a divisor of 4-4.
+The set of such rational numbers consists of the integers
+$\pm 1,\pm 2,\pm 4,$
+and those which are not integers
+$\pm\frac{1}{2},\pm\frac{1}{4}.$
+We note that $2\cdot\bar{P}(T)=4-T(4T^{2}+3)$.
+For every $n\in\{\pm 1,\pm 2,\pm 4\}$, the integer $4n^{2}+3$ is not a divisor of $4$, which implies that $2\cdot\bar{P}(n)\neq 0$.
+One can also check that
+$\begin{array}[]{llll}\bar{P}(\frac{1}{2})=1,&\bar{P}(-\frac{1}{2})=3,&\bar{P}(\frac{1}{4})=\frac{51}{32},&\bar{P}(-\frac{1}{4})=\frac{77}{32}.\end{array}$
+Therefore, we can conclude that $\bar{P}$ does not have a root in $\mathbb{Q}$, which contradicts Proposition 3.5 and Proposition 3.17. ∎
 A Proof of Proposition 2.7
@@
 $(CF^{*}(\psi(\mathbb{R}^{3}),L_{K_{0}};H),d_{J})$
-with coefficients in the Laurent polynomial ring $\mathbb{F}_{2}[\lambda^{\pm}]$. It is a slight enhancement of the Floer cochain complex over $\mathbb{F}_{2}$ constructed as in [20, Section 8].
+with coefficients in the Laurent polynomial ring $\mathbb{F}_{2}[\lambda^{\pm}]$. It is a slight enhancement of the Floer cochain complex over $\mathbb{F}_{2}$ constructed as in [19, Section 8].
 We choose a Hamiltonian $H\colon T^{*}\mathbb{R}^{3}\times[0,1]\to\mathbb{R}$ with compact support such that the Lagrangian submanifold $P_{H}\coloneqq(\varphi^{1}_{H}\circ\psi)(\mathbb{R}^{3})$ intersects $L_{K_{0}}$ transversely.
@@
 From the orientations of $P_{H}$ and $L_{K_{0}}$, a $\mathbb{Z}/2$-grading $|x|\in\mathbb{Z}/2$ is defined for every $x\in P_{H}\cap L_{K_{0}}$.
-See [19, Example 2.7, Section 2.d].
+See [18, Example 2.7, Section 2.d].
 Note that the grading is shifted by $1$ if we use the opposite orientation of $P_{H}$.
@@
 For any $x\in P_{H}\cap L_{K_{0}}$, we choose a path $c_{x}\colon[0,1]\to L_{K_{0}}$ such that $c_{x}(0)=z_{0}$ and $c_{x}(1)=x$.
-Let $c_{x}^{-1}$ denote the path $[0,1]\to L_{K_{0}}\colon t\mapsto c_{x}(1-t)$.
+Let $c_{x}^{-1}$ denote the path $0,1]\to L_{K_{0}}\colon t\mapsto c_{x}(1-t)$.
 For every $x,y\in P_{H}\cap L_{K_{0}}$ and a homology class $A\in H_{1}(L_{K_{0}})$, we define $\mathcal{M}_{(H,J),A}(x,y)$ to be the space of maps $u\colon\mathbb{R}\times[0,1]\to T^{*}\mathbb{R}^{3}$ satisfying:
-  • su($\partial_{s}u(s,t)+J_{u(s,t)}(\partial_{t}u(s,t)-(X^{t}_{H})_{u(s,t)})=0$ for every $(s,t)\in\mathbb{R}\times[0,1]$.
+  • $\partial_{s}u(s,t)+J_{u(s,t)}(\partial_{t}u(s,t)-(X^{t}_{H})_{u(s,t)})=0$ for every $(s,t)\in\mathbb{R}\times[0,1]$.
   • $u(\mathbb{R}\times\{0\})\subset\psi(\mathbb{R}^{3})$ and $u(\mathbb{R}\times\{1\})\subset L_{K_{0}}$.
   • $\lim_{s\to-\infty}u(s,t)=\varphi^{t}_{H}((\varphi^{1}_{H})^{-1}(x))$ and $\lim_{s\to\infty}u(s,t)=\varphi_{H}^{t}((\varphi^{1}_{H})^{-1}(y))$ $C^{\infty}$-uniformly on $t\in[0,1]$.
-  • $A\in H_{1}(L_{K_{0}})$ is represented by a loop based at $z_{0}$ obtained by concatenating the three paths in $L_{K_{0}}$: $c_{x}$, the path $[-\infty,\infty]\to L_{K_{0}}\colon s\mapsto u(s,1)$ from $x$ to $y$, and $c_{y}^{-1}$.
+  • $A\in H_{1}(L_{K_{0}})$ is represented by a loop based at $z_{0}$ obtained by concatenating the three paths in $L_{K_{0}}$: $c_{x}$, the path $\infty,\infty]\to L_{K_{0}}\colon s\mapsto u(s,1)$ from $x$ to $y$, and $c_{y}^{-1}$.
 Figure 2 depicts $u(\mathbb{R}\times[0,1])\subset T^{*}\mathbb{R}^{3}$ and the three paths in $L_{K_{0}}$.
 The group $\mathbb{R}$ acts on $\mathcal{M}_{(H,J),A}(x,y)$ by $(u\cdot\sigma)(s,t)=u(s+\sigma,t)$ for every $\sigma\in\mathbb{R}$.
-To define the $\mathbb{Z}/2$-graded Floer cochain complex, we refer to [19, Section 2.f].
-We take the pair $(H,J)$ to be regular so that the moduli space $\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ is cut out transversely.
-The dimension of $\coprod_{A\in H_{1}(L_{K_{0}})}\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ is equal to $|x|-|y|-1$ modulo $2$.
-Moreover, by the Gromov compactness, the union of 00-dimensional components of $\coprod_{A\in H_{1}(L_{K_{0}})}\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ is a finite set.
-We define $n_{(H,J),A}(x,y)\in\mathbb{F}_{2}$ to be the modulo $2$ count of the number of the 00-dimensional components of $\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$.
+To define the $\mathbb{Z}/2$-graded Floer cochain complex, we refer to [18, Section 2.f].
+We take the pair $(H,J)$ to be regular so that the moduli space $\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ is cut out transversely for every $x,y\in P_{H}\cap L_{K_{0}}$ such that $x\neq y$.
+The dimension of $\coprod_{A\in H_{1}(L_{K_{0}})}\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ is equal to |x||y|1|x|-|y|-1 modulo $2$.
+Moreover, by the Gromov compactness, the union of 0-dimensional components of $\coprod_{A\in H_{1}(L_{K_{0}})}\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ is a finite set.
+We define $n_{(H,J),A}(x,y)\in\mathbb{F}_{2}$ to be the modulo $2$ count of the number of the 0-dimensional components of $\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$.
 We define $CF^{*}(\psi(\mathbb{R}^{3}),L_{K_{0}};H)$ to be the $\mathbb{Z}/2$-graded $\mathbb{F}_{2}[\lambda^{\pm}]$-module freely generated by the set $P_{H}\cap L_{K_{0}}$ whose grading is given by $|\cdot|\colon P_{H}\cap L_{K_{0}}\to\mathbb{Z}/2$.
@@
 for every $y\in P_{H}\cap L_{K_{0}}$.
-By a standard argument using the Gromov compactification of $\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ for each $A\in H_{1}(L_{K_{0}})$ when its dimension is $1$, we can prove that $d_{J}\circ d_{J}=0$.
+By a standard argument using the Gromov compactification of the $1$-dimensional components of $\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ for each $A\in H_{1}(L_{K_{0}})$, we can prove that $d_{J}\circ d_{J}=0$.
 From the cochain complex, we obtain a cohomology group
 $HF^{*}(\psi(\mathbb{R}^{3}),L_{K_{0}};(H,J))\coloneqq\operatorname{Ker}d_{J}/\operatorname{Im}d_{J}.$
-Its isomorphism class as a $\mathbb{Z}/2$-graded $\mathbb{F}_{2}[\lambda^{\pm}]$-module is independent of the choice of the paths $\{c_{x}\mid x\in P_{H}\cap L_{K_{0}}\}$.
-To see this claim, we choose another path $c^{\prime}_{x}$ instead of $c_{x}$ for every $x\in P_{H}\cap L_{K_{0}}$ and define $d^{\prime}_{J}$ by using $\{c^{\prime}_{x}\mid x\in P_{H}\cap L_{K_{0}}\}$ in the same way as $d_{J}$.
+We claim that its isomorphism class as a $\mathbb{Z}/2$-graded $\mathbb{F}_{2}[\lambda^{\pm}]$-module is independent of the choice of the paths $\{c_{x}\mid x\in P_{H}\cap L_{K_{0}}\}$.
+To see this, we choose another path $c^{\prime}_{x}$ instead of $c_{x}$ for every $x\in P_{H}\cap L_{K_{0}}$ and define $d^{\prime}_{J}$ by using $\{c^{\prime}_{x}\mid x\in P_{H}\cap L_{K_{0}}\}$ in the same way as $d_{J}$.
 Take $m_{x}\in\mathbb{Z}$ such that $m_{x}[K_{0}]\in H_{1}(L_{K_{0}})$ is represented by a loop obtained by concatenating $c^{\prime}_{x}$ and $c_{x}^{-1}$.
@@
 We may assume that for sufficiently small $\varepsilon>0$, $\psi(L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3})\subset L_{K_{0}}$.
-We fix an orientation of $\psi(K_{1})$ via $\psi\colon K_{1}\to\psi(K_{1})$ and define $a\in\mathbb{Z}$ by $[\psi(K_{1})]=a[K_{0}]\in H_{1}(L_{K_{0}})$.
-Let $(-1)^{\sigma}$ denote the sign of orientation of the open embedding $\psi\colon L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to L_{K_{0}}$.
+We fix an orientation of $\psi(K_{1})$ via the diffeomorphism $\left.\psi\right|_{K_{1}}\colon K_{1}\to\psi(K_{1})$ and define $a\in\mathbb{Z}$ by $\psi(K_{1})]=a[K_{0}]\in H_{1}(L_{K_{0}})$.
+Let $(-1)^{\sigma}$ denote the sign of orientation of the embedding $\left.\psi\right|_{L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}}\colon L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to L_{K_{0}}$.
 Then, the claim of the proposition is equivalent to that $a\in\{1,-1\}$ and $\sigma\equiv 0\mod 2$.
+Since $\sigma\equiv 0\mod 2$ is already shown in [16, Lemma 4.4], we only give a proof for $a\in\{1,-1\}$ (see also Remark A.2 below).
 We choose a Mose function $f\colon\psi(K_{1})\to\mathbb{R}$ which has only two critical points $x_{0},x_{1}\in\psi(K_{1})$ such that $f$ takes its minimum at $x_{0}$ and its maximum at $x_{1}$.
@@
   • $\lim_{s\to-\infty}\gamma(s)=x_{0}$ and $\lim_{s\to\infty}\gamma(s)=x_{1}$.
-Since $\psi(\mathbb{R}^{3})$ and $L_{K_{0}}$ intersects cleanly along $\psi(K_{1})$ and $\psi(K_{1})$ is connected, we can apply [18, Theorem 3.4.11, Lemma 3.4.12] to get the following results: From the pair $(f,g)$ with $||f||_{C^{1}}$ sufficiently small, we can construct a Hamiltonian $H_{f}\colon T^{*}\mathbb{R}^{3}\to\mathbb{R}$ supported in a neighborhood of $\psi(K_{1})$ and an almost complex structure $J_{g}$ on $T^{*}\mathbb{R}^{3}$ such that the pair $(H_{f},J_{g})$ is regular and the following hold:
+Since $\psi(\mathbb{R}^{3})$ and $L_{K_{0}}$ intersects cleanly along $\psi(K_{1})$ and $\psi(K_{1})$ is connected, we can apply [17, Theorem 3.4.11, Lemma 3.4.12] to get the following results: From the pair $(f,g)$ with $f||_{C^{1}}$ sufficiently small, we can construct a Hamiltonian $H_{f}\colon T^{*}\mathbb{R}^{3}\to\mathbb{R}$ supported in a neighborhood of $\psi(K_{1})$ and an almost complex structure $J_{g}$ on $T^{*}\mathbb{R}^{3}$ such that the pair $(H_{f},J_{g})$ is regular and the following hold:
   • $P_{H_{f}}\cap L_{K_{0}}=\{x_{0},x_{1}\}$.
@@
 $A=k[K_{0}]+[\psi(K_{1})]=(k+a)[K_{0}]$
-since $[\psi(K_{1})]\in H_{1}(L_{K_{0}})$ is represented by a loop obtained by concatenating the two paths: $[-\infty,\infty]\to L_{K_{0}}\colon s\mapsto\gamma_{1}(-s)$ and $[-\infty,\infty]\to L_{K_{0}}\colon s\mapsto\gamma_{2}(s)$.
+since $\psi(K_{1})]\in H_{1}(L_{K_{0}})$ is represented by a loop obtained by concatenating the two paths: $\infty,\infty]\to L_{K_{0}}\colon s\mapsto\gamma_{1}(-s)$ and $\infty,\infty]\to L_{K_{0}}\colon s\mapsto\gamma_{2}(s)$.
 From this observation, the differential $d_{J_{g}}\colon\mathbb{F}_{2}[\lambda^{\pm}]\cdot x_{1}\to\mathbb{F}_{2}[\lambda^{\pm}]\cdot x_{0}$ is given by
@@
 $\begin{array}[]{lll}HF^{0+\sigma_{\psi}}(\psi(\mathbb{R}^{3}),L_{K_{0}})=0,&HF^{1+\sigma_{\psi}}(\psi(\mathbb{R}^{3}),L_{K_{0}})\cong\mathbb{F}_{2}[\lambda^{\pm}]/(1+\lambda^{a})&\text{ if }a\neq 0,\\ HF^{0+\sigma_{\psi}}(\psi(\mathbb{R}^{3}),L_{K_{0}})\cong\mathbb{F}_{2}[\lambda^{\pm}],&HF^{1+\sigma_{\psi}}(\psi(\mathbb{R}^{3}),L_{K_{0}})\cong\mathbb{F}_{2}[\lambda^{\pm}]&\text{ if }a=0.\end{array}$
-Here, $\sigma_{\psi}\in\mathbb{Z}/2$ is the shift of degrees.
-Note that $\sigma_{\psi}$ depends only on the orientation of $L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}$ induced by the open embedding $\psi\colon L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to L_{K_{0}}$ and the fixed orientation of $L_{K_{0}}$.
-As $\mathbb{F}_{2}$-vector spaces,
+Here, $\sigma_{\psi}\in\mathbb{Z}/2$ is the shift of degrees. As $\mathbb{F}_{2}$-vector spaces,
 $\begin{array}[]{lll}\dim_{\mathbb{F}_{2}}HF^{0+\sigma_{\psi}}(\psi(\mathbb{R}^{3}),L_{K_{0}})=0,&\dim_{\mathbb{F}_{2}}HF^{1+\sigma_{\psi}}(\psi(\mathbb{R}^{3}),L_{K_{0}})=|a|&\text{ if }a\neq 0,\\ \dim_{\mathbb{F}_{2}}HF^{0+\sigma_{\psi}}(\psi(\mathbb{R}^{3}),L_{K_{0}})=\infty,&\dim_{\mathbb{F}_{2}}HF^{1+\sigma_{\psi}}(\psi(\mathbb{R}^{3}),L_{K_{0}})=\infty&\text{ if }a=0.\end{array}$ (29)
@@
 $\begin{array}[]{ll}\dim_{\mathbb{F}_{2}}HF^{0+\sigma_{\operatorname{id}}}(\mathbb{R}^{3},L_{K_{0}})=0,&\dim_{\mathbb{F}_{2}}HF^{1+\sigma_{\operatorname{id}}}(\mathbb{R}^{3},L_{K_{0}})=1.\end{array}$ (31)
-From the definition of the orientation sign $(-1)^{\sigma}$, the shift of degrees $\sigma_{\psi}$ is equal to $\sigma_{\operatorname{id}}+\sigma$ modulo $2$.
-From (15), (17) and Theorem A.1, it follows that $a\in\{1,-1\}$ and $\sigma\equiv 0\mod 2$.
-∎
+From (29), (31) and Theorem A.1, it follows that $\sigma_{\psi}\equiv\sigma_{\operatorname{id}}\mod 2$ and $a\in\{1,-1\}$. ∎
+Remark A.2.
+The fact $\sigma\equiv 0\mod 2$ was proved in [16, Lemma 4.4] by introducing $\mathbb{Z}$-gradings of Lagrangian submanifolds which are lifts of orientations of Lagrangian submanifolds.
+Analogous to [16, Remark 2.9, Lemma 4.4], by an argument on $\mathbb{Z}/2$-grading in the present case, we can deduce $\sigma\equiv 0\mod 2$ from $\sigma_{\psi}\equiv\sigma_{\operatorname{id}}\mod 2$.

### v3

@@
 Abstract
-Let $K$ be the connected sum in $\mathbb{R}^{3}$ of an arbitrary knot and the $(2,q)$-torus knot for $q\neq\pm 1$, and consider its conormal bundle $L_{K}$ in $T^{*}\mathbb{R}^{3}$.
-We prove that there is no compactly supported Hamiltonian diffeomorphism $\varphi$ on $T^{*}\mathbb{R}^{3}$ such that $\varphi(L_{K})$ and the zero section $\mathbb{R}^{3}$ intersect cleanly along the unknot in $\mathbb{R}^{3}$.
-For the proof, we examine the knot DGA, whose homology was proved to be isomorphic to the knot contact homology by Ekholm, Etnyre, Ng and Sullivan, with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$.
-We find an algebraic constraint imposed on every $\mathbb{Q}$-valued augmentation of the knot DGA of $K$ when $q\geq 3$.
-This constraint involves rational roots of polynomials in $\mathbb{Q}[T]$, and we apply Hilbert’s irreducibility theorem to complete the proof.
-Using a similar idea, we also prove an analogous result replacing the $(2,q)$-torus knot with the figure-eight knot.
+Let $K$ be a knot in $\mathbb{R}^{3}$ which has the $(2,q)$-torus knot for $q\neq\pm 1$ or the figure-eight knot as a component of connected sum.
+For its conormal bundle $L_{K}$ in $T^{*}\mathbb{R}^{3}$, we show that there is no compactly supported Hamiltonian diffeomorphism $\varphi$ on $T^{*}\mathbb{R}^{3}$ such that $\varphi(L_{K})$ intersects the zero section $\mathbb{R}^{3}$ cleanly along the unknot in $\mathbb{R}^{3}$.
+Using symplectic field theory, the proof is reduced to studying the augmentation variety $V_{\mathbb{k}}(K)$ of $K$ over a filed $\mathbb{k}$.
+The key point of this paper is finding an algebraic constraint on $V_{\mathbb{k}}(K)$ which is valid only when $\mathbb{k}$ is not algebraically closed, and the proof is completed by some arithmetic argument with $\mathbb{k}=\mathbb{Q}$.
 1 Introduction
-Background.
+Backgraound.
 Let us consider the cotangent bundle $T^{*}\mathbb{R}^{3}$ of $\mathbb{R}^{3}$ with a canonical symplectic structure.
-We identify the image of the zero section with $\mathbb{R}^{3}$.
-Given any knot $K$ in $\mathbb{R}^{3}$, its conormal bundle is defined by
-$L_{K}\coloneqq\{(q,p)\in T^{*}\mathbb{R}^{3}\mid q\in K,\ \langle p,v\rangle=0\text{ for every }v\in T_{q}K\}.$
+We identify the zero section with $\mathbb{R}^{3}$.
+Given any knot $K$ in $\mathbb{R}^{3}$, let $L_{K}$ denote its conormal bundle.
 It is a Lagrangian submanifold of $T^{*}\mathbb{R}^{3}$ and intersects $\mathbb{R}^{3}$ cleanly along the knot $K$.
 Let $\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$ denote the set of Hamiltonian diffeomorphisms on $T^{*}\mathbb{R}^{3}$ associated to compactly supported Hamiltonians on $T^{*}\mathbb{R}^{3}$.
-Using these notations, we ask the following question, which was observed by the author in [16].
-Question 1.1 (Question 1.1 of [16]).
-Is there $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K})$ and $\mathbb{R}^{3}$ have a clean intersection along a knot in $\mathbb{R}^{3}$ whose isotopy class is different from $K$?
-We remark that it is essential that $\varphi$ is a Hamiltonian diffeomorphism.
-Indeed, for any pair of knots $(K_{0},K_{1})$ in $\mathbb{R}^{3}$, there exists a $C^{\infty}$ isotopy $(\varphi^{t})_{t\in[0,1]}$ on $T^{*}\mathbb{R}^{3}$ with compact support such that $\varphi^{1}(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along $K_{1}$.
-See [16, Proposition 4.1].
-As a special case, the answer to Question 1.1 when $K$ is the unknot is ‘no’.
-This can be deduced from the result by Ganatra-Pomerleano [9, Proposition 6.29] about clean Lagrangian intersections in a plumbing of two copies of the cotangent bundle of $S^{3}$ along the unknots (see [16, Section 1]).
-Smith-Wemyss [20, Proposition 4.10] and Asplund-Li [1, Theorem 1.3] also studies clean Lagrangian intersections in plumbings of two copies of $T^{*}S^{3}$ along the unknots.
-These results are concerned with Smith’s question on the ‘persistence (or rigidity) of unknottedness of Lagrangian intersections’ [9, Question 1.5].
-In a general case, under the assumption that $\varphi(L_{K})$ for $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ intersects $\mathbb{R}^{3}$ cleanly along a knot $K^{\prime}$ in $\mathbb{R}^{3}$, the author proved [16, Theorem 1.1].
-It relates the augmentation variety in $(\mathbb{C}^{*})^{2}$ of $K$ with that of $K^{\prime}$.
-This result leads us to the same answer to Question 1.1 as above when $K$ is the unknot [16, Theorem 1.3].
-We can also deduce some constraints on the knot type of $K^{\prime}$ when $K$ has certain non-trivial knot types [16, Theorem 1.4, 1.5].
-Main results. In fact, the aforementioned result [16, Theorem 1.1] is useless to solve the following weaker version of Question 1.1. For the reason, see Remark 1.5 below.
-Question 1.2.
+Using these notations, we ask the following question, which was observed by the author in [15]: Is there $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K})$ intersects $\mathbb{R}^{3}$ cleanly along a knot whose isotopy class in $\mathbb{R}^{3}$ is different from $K$?
+As a special case, the answer to this question when $K$ is the unknot is ‘no’.
+This can be deduced from the result by Ganatra-Pomerleano [8, Proposition 6.29] about clean Lagrangian intersections in a plumbing of two copies of the cotangent bundle of $S^{3}$ along the unknots (see [15, Section 1]).
+Smith-Wemyss [19, Proposition 4.10] and Asplund-Li [1, Theorem 1.3] also studied clean Lagrangian intersections in plumbings of two copies of $T^{*}S^{3}$ along the unknots.
+These results are concerned with Smith’s question on the ‘persistence (or rigidity) of unknottedness of Lagrangian intersections’ [8, Question 1.5].
+In a general case, under the assumption that $\varphi(L_{K})$ for $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ intersects $\mathbb{R}^{3}$ cleanly along a knot $K^{\prime}$ in $\mathbb{R}^{3}$, the author proved a theorem [15, Theorem 1.1], which relates the augmentation variety in $(\mathbb{C}^{*})^{2}$ of $K$ with that of $K^{\prime}$.
+This result leads us to the same answer to the question as above when $K$ is the unknot [15, Theorem 1.3].
+We can also deduce some constraints on $K^{\prime}$ when $K$ has certain non-trivial knot types [15, Theorem 1.4, 1.5].
+In fact, the aforementioned result [15, Theorem 1.1] is useless to solve the following question which asks the persistence of knottedness of clean intersections. For the reason, see Remark 1.3.
+Question 1.1.
 If $K$ is not the unknot, is there $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{n})$ such that $\varphi(L_{K})$ intersects $\mathbb{R}^{3}$ cleanly along the unknot in $\mathbb{R}^{3}$?
-The goal of this paper is to give a negative answer to this question when $K$ has a certain knot type.
-As a notation, let $K_{1}\#K_{2}$ denote the connected sum of knots $K_{1}$ and $K_{2}$ in $\mathbb{R}^{3}$.
-Let $T_{(2,2m+1)}$ for $m\in\mathbb{Z}$ denote the $(2,2m+1)$-torus knot and let $4_{1}$ denote the figure-eight knot.
+Main results.
+The goal of this paper is to give a negative answer to Question 1.1 when $K$ has a certain knot type by using augmentation variety over $\mathbb{Q}$ and some arithmetic argument.
+Let $T_{(2,2m+1)}$ for $m\in\mathbb{Z}$ denote the $(2,2m+1)$-torus knot and $4_{1}$ denote the figure-eight knot.
 Theorem 1.1 (Theorem 3.14 and Theorem 3.17).
@@
 Then, there is no $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K})$ intersects $\mathbb{R}^{3}$ cleanly along the unknot in $\mathbb{R}^{3}$.
-For the proof, the geometric arguments in this paper are similar to those in [16], which are outlined as follows: Let $K_{0}$ and $K_{1}$ be knots in $\mathbb{R}^{3}$.
-For $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$, suppose that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ intersect cleanly along $K_{1}$.
-We use the results by Ekholm-Etnyre-Ng-Sullivan [6, 3] which show that the Chekanov-Eliashberg DGA of the unit conormal bundle of $K_{i}$ is stable tame isomorphic to the knot DGA of $K_{i}$ for $i=0,1$.
-The knot DGA, introduced by Ng [12], is defined combinatorially from a braid representation of a knot.
-Here, we refer to the one with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$ defined in [14, 3, 15].
-We then apply a DGA map associated to an exact Lagrangian cobordism between the unit conormal bundles of $K_{0}$ and $K_{1}$.
-This is a part of the symplectic field theory introduced by Eliashberg-Givental-Hofer [8].
-Consequently, we obtain a DGA map from the knot DGA of $K_{0}$ to that of $K_{1}$.
-From a simple description of the knot DGA of the unknot, we obtain the following result about augmentations of the knot DGA of $K_{0}$.
-Proposition 1.2 (Proposition 3.5).
-Let $\mathbb{k}$ be a field.
-Suppose that $K_{1}$ is the unknot.
-Then, for every $y,Z\in\mathbb{k}^{*}$ satisfying $y\notin\{Z^{n}\mid n\in\mathbb{Z}\}$, there exists an augmentation $\varepsilon$ of the knot DGA of $K_{0}$ to $\mathbb{k}$ such that $\varepsilon(\mu)=y$ and $\varepsilon(U)=Z$.
-There are two new points in algebraic arguments in this paper. Let us explain them by focusing on the knot $K^{\prime}\#T_{(2,2m+1)}$ for $m\in\mathbb{Z}_{\geq 1}$.
-The first point is that we find a constraint on augmentations of some knot DGAs in terms of roots of polynomials in $\mathbb{k}[T]$, where $\mathbb{k}$ is any field.
-Actually, we find a specific polynomial $P_{m}\in\mathbb{Z}[\mu,U][T]$ for every $m\in\mathbb{Z}_{\geq 1}$, whose degree in $T$ is $m+1$, with the following property (see Proposition 3.10): For any augmentation $\varepsilon$ of the knot DGA of $K^{\prime}\#T_{(2,2m+1)}$ to $\mathbb{k}$ such that $\varepsilon(\mu)=y$ and $\varepsilon(U)=Z$, the polynomial $\left.P_{m}\right|_{\mu=y,U=Z}\in\mathbb{k}[T]$ has a root in $\mathbb{k}$.
-About the derivation of $P_{m}$, see Remark 1.3 below.
-We will focus on the case $\mathbb{k}=\mathbb{Q}$.
-It is worth noting that this constraint on augmentations is trivial if $\mathbb{k}$ is algebraically closed.
-The second point is that we use a classical result in number theory.
-Indeed, Theorem 1.1 when $K=K^{\prime}\#T_{(2,2m+1)}$ for $m\in\mathbb{Z}_{\geq 1}$ is proved by a contradiction as follows: Suppose that $K_{0}=K^{\prime}\#T_{(2,2m+1)}$ and $K_{1}$ is the unknot.
-Then, Proposition 1.2 and the property of $P_{m}$ for $\mathbb{k}=\mathbb{Q}$ imply that $\left.P_{m}\right|_{\mu=2,U=Z}\in\mathbb{Q}[T]$ has a root in $\mathbb{Q}$ for every $Z\in\mathbb{Q}^{*}\setminus\{2^{\pm 1}\}$.
-On the other hand, by using Hilbert’s irreducibility theorem, we can show that there are infinitely many $Z\in\mathbb{Q}$ such that $\left.P_{m}\right|_{\mu=2,U=Z}\in\mathbb{Q}[T]$ does not have a root in $\mathbb{Q}$, which is a contradiction.
-(From the case of $m\geq 1$, we can deduce Theorem 1.1 when $K=K^{\prime}\#T_{(2,-2m-1)}$.)
-The proof of Theorem 1.1 when $K=K^{\prime}\#4_{1}$ goes in a similar way.
-We explicitly give a polynomial $P\in\mathbb{Z}[\mu^{\pm},U^{\pm}][T]$, whose degree in $T$ is $3$, with the following property (see Proposition 3.17): For any augmentation $\varepsilon$ of the knot DGA of $K^{\prime}\#4_{1}$ to $\mathbb{k}$ such that $\varepsilon(\mu)=y$ and $\varepsilon(U)=Z$, the polynomial $\left.P\right|_{\mu=y,U=Z}\in\mathbb{k}[T]$ has a root in $\mathbb{k}$.
-Using concrete calculations and the rational root theorem, we show that $\left.P\right|_{\mu=-1,U=-2}\in\mathbb{Q}[T]$ does not have a root in $\mathbb{Q}$.
-Combining with Proposition 1.2 for $\mathbb{k}=\mathbb{Q}$, Theorem 1.1 when $K=K^{\prime}\#4_{1}$ is deduced.
+For the proof, we use Theorem 1.2 below.
+For an oriented knot $K$ in $\mathbb{R}^{3}$, the knot DGA is defined combinatorially from a braid representation of a knot.
+In this paper, we refer to a definition in [13, 3, 14] with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$.
+Fix a field $\mathbb{k}$.
+From the set of augmentations of the knot DGA of $K$ to $\mathbb{k}$, following [14, Definition 5.1], we will define the augmentation variety $V_{\mathbb{k}}(K)\subset(\mathbb{k}^{*})^{3}$ in Definition 3.4.
+Let $(x,y,Z)$ be the coordinate of $\mathbb{k}^{3}$.
+Theorem 1.2 (Corollary 3.5).
+For $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ and oriented knots $K_{0}$ and $K_{1}$ in $\mathbb{R}^{3}$, suppose that $\varphi(L_{K_{0}})$ intersects $\mathbb{R}^{3}$ cleanly along $K_{1}$.
+Then, there exist $a\in\{\pm 1\}$ and integers $b,c$ and $d$ for which we have an injective map
+$V_{\mathbb{k}}(K_{1})\hookrightarrow V_{\mathbb{k}}(K_{0})\colon(x,y,Z)\mapsto(x^{a}y^{b}Z^{c},y^{a}Z^{d},Z).$
+This is a refinement of [15, Theorem 1.1], and the proof is parallel to [15] except an enhancement of the coefficient ring.
+We use the results by Ekholm-Etnyre-Ng-Sullivan [6, 3] which show that the Chekanov-Eliashberg DGA of the unit conormal bundle of $K_{i}$ is stable tame isomorphic to the knot DGA of $K_{i}$ for $i\in\{0,1\}$, and we compose the stable tame isomorphisms with a DGA map associated to an exact Lagrangian cobordism.
+For more on the coefficient ring, see Subsection 2.5.1.
 Remark 1.3.
-The polynomials $P_{m}$ for $K^{\prime}\#T_{(2,2m+1)}$ and $P$ for $K^{\prime}\#4_{1}$ are found from concrete computations on some part of (the homology of) the knot DGA which comes from the component $T_{(2,2m+1)}$ or $4_{1}$ of the connected sum.
-Our computation relies on braid representations of $T_{(2,2m+1)}$ and $4_{1}$, which are rather simple.
-It might be possible to find a polynomial with a similar property as $P_{m}$ for more complicated knots, such as the $(p,q)$-torus knot when $3\leq p<|q|$.
-We also provide a theorem which is a refinement of [16, Theorem 1.1].
-For an oriented knot $K$ in $\mathbb{R}^{3}$, let $V_{K}$ denote the augmentation variety defined by [15, Definition 5.1] as a subset of $(\mathbb{C}^{*})^{3}$.
-Let $(x,y,Z)$ be the coordinate of $\mathbb{C}^{3}$.
-Theorem 1.4 (Corollary 3.7).
-For oriented knots $K_{0}$ and $K_{1}$ in $\mathbb{R}^{3}$, suppose that there exists $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ intersects $\mathbb{R}^{3}$ cleanly along $K_{1}$.
-Then, there exist $a\in\{\pm 1\}$ and integers $b,c$ and $d$ for which we have an injective map
-$V_{K_{1}}\hookrightarrow V_{K_{0}}\colon(x,y,Z)\mapsto(x^{a}y^{b}Z^{c},y^{a}Z^{d},Z).$
-Remark 1.5.
-Identifying $\{(x,y,Z)\in(\mathbb{C}^{*})^{3}\mid Z=1\}$ with $(\mathbb{C}^{*})^{2}$, we obtain $V_{K}\cap\{Z=1\}\subset(\mathbb{C}^{*})^{2}$, which is a version of augmentation variety used in [16].
+Identifying $\{(x,y,Z)\in(\mathbb{k}^{*})^{3}\mid Z=1\}$ with $(\mathbb{k}^{*})^{2}$, we obtain $V_{\mathbb{k}}(K)\cap\{Z=1\}\subset(\mathbb{k}^{*})^{2}$, which is a version of augmentation variety used in [15].
 This corresponds to the reduction of the coefficient ring of the knot DGAs from $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$ to $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$ by setting $U=1$.
 It follows from Theorem 1.2 that there exists an injective map
-$V_{K_{1}}\cap\{Z=1\}\hookrightarrow V_{K_{0}}\cap\{Z=1\}\colon(x,y)\mapsto(x^{a}y^{b},y^{a})$
+$V_{\mathbb{k}}(K_{1})\cap\{Z=1\}\hookrightarrow V_{\mathbb{k}}(K_{0})\cap\{Z=1\}\colon(x,y)\mapsto(x^{a}y^{b},y^{a})$
 for some $a\in\{\pm 1\}$ and $b\in\mathbb{Z}$.
-This is a slight refinement of [16, Theorem 1.1], but it gives no constraint on $K_{0}$ if we suppose that $K_{1}$ is the unknot.
-The reason is that $V_{K_{1}}\cap\{Z=1\}=\{x=1\}\cup\{y=1\}$ if $K_{1}$ is the unknot (see [15, Example 3.13]) and $\{x=1\}\cup\{y=1\}$ is a subset of $V_{K_{0}}\cap\{Z=1\}$ for every knot $K_{0}$ (see [13, Proposition 5.6] and [15, Exercise 3.17]).
-Notations.
-In this paper, for any manifold $X$, we take $\mathbb{Z}$ as the coefficient ring of the singular homology group of $X$ and denote it by $H_{*}(X)$, unless otherwise specified.
-Likewise, the relative homology group over $\mathbb{Z}$ is denoted by $H_{*}(X,A)$ for any submanifold $A$ of $X$.
+This is a slight refinement of [15, Theorem 1.1], but it gives no constraint on $K_{0}$ if $K_{1}=O$, where $O$ is the unknot.
+The reason is that a relation $V_{\mathbb{k}}(O)\cap\{Z=1\}\subset V_{\mathbb{k}}(K)\cap\{Z=1\}$ automatically holds for any knot $K$.
+See [12, Section 5.2]
+Compared with [15], where $V_{\mathbb{C}}(K_{i})\cap\{Z=1\}\subset(\mathbb{C}^{*})^{2}$ for $i\in\{0,1\}$ were used, a novel point of this paper lies in an algebraic argument to deduce Theorem 1.1 from Theorem 1.2.
+Let us explain the idea in the case $K=K^{\prime}\#T_{(2,2m+1)}$ for $m\in\mathbb{Z}_{\geq 1}$.
+It suffices to show that for some field $\mathbb{k}$, $V_{\mathbb{k}}(O)$ cannot be embedded into $V_{\mathbb{k}}(K^{\prime}\#T_{(2,2m+1)})$ by the map of Theorem 1.2.
+From the knot DGA of $O$, which is quite simple, it is easy to compute that $V_{\mathbb{k}}(O)=\{Z-x-y+xy=0\}$.
+However, determining $V_{\mathbb{k}}(K^{\prime}\#T_{(2,2m+1)})$ is hard because the knot DGA is complicated in general.
+As a matter of fact, for the closure of a braid with $n$ strands, the 0-th homology of its knot DGA has n(n1)n(n-1) generators and $2n^{2}$ relations.
+See Subsection 3.2.1.
+To solve this problem, we use an algebraic constraint on $V_{\mathbb{k}}(K^{\prime}\#T_{(2,2m+1)})$.
+In Section 3.2, by focusing on specific two relations in the 0-th homology, we will extract a polynomial $P_{m}\in\mathbb{Z}[\mu^{\pm},U^{\pm}][T]$ defined in an explicit way with the following property:
+$V_{\mathbb{k}}(K^{\prime}\#T_{(2,2m+1)})\subset\{(x,y,Z)\mid\left.P_{m}\right|_{\mu=y,U=Z}\in\mathbb{k}[T]\text{ has a root in }\mathbb{k}\}.$
+Then, the proof of Theorem 1.1 is reduced to finding $(x,y,Z)\in V_{\mathbb{k}}(O)$ such that the $\mathbb{k}$-coefficient polynomial $\left.P_{m}\right|_{\mu=y^{a}Z^{d},U=Z}$ does not have a root in $\mathbb{k}$.
+It is worth noting that this approach does not work if $\mathbb{k}$ is algebraically closed.
+Actually, we take $\mathbb{k}=\mathbb{Q}$ and apply Hilbert’s irreducibility theorem in the proof of Theorem 3.14.
+In the case $K=K^{\prime}\#4_{1}$, we can explicitly find a polynomial (of degree $3$ in $T$) with a similar property.
+We take $\mathbb{k}=\mathbb{Q}$ and use the rational root theorem in the proof of Theorem 3.17.
 Organization of paper.
@@
 We then define a DGA map associated to an exact Lagrangian cobordism in $\mathbb{R}\times Y$.
-In Section 2.5, we will see how to apply this DGA map to a setting with clean Lagrangian intersection as in Question 1.1.
+In Section 2.5, we will see how to apply this DGA map to a setting with clean Lagrangian intersection.
 In Section 3, we combine the argument in Section 2.5 with the results of [3, 14] about the knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$.
-Proposition 1.2 and Theorem 1.4 are proved in Section 3.1.
+Theorem 1.2 is proved in Section 3.1.
 Theorem 1.1 for $K=K^{\prime}\#T_{(2,2m+1)}$ with $m\neq 0,-1$ (resp.
@@
 In Appendix A, by using Lagrangian Floer cohomology group over $\mathbb{F}_{2}[\lambda^{\pm}]$, we prove Proposition 2.7 which gives a homological constraint on $1$-dimensional clean intersections.
-This strengthen a result in [16].
+This strengthen a result in [15, Section 2].
 Nevertheless, the present proof is much simpler than that of [15].
+Notations. In this paper, we take $\mathbb{Z}$ as the coefficient ring of the singular homology groups, unless otherwise mentioned.
 Acknowledgements.
@@
 We fix a $2n$-dimensional simply connected manifold $P$ with a Liouville $1$-form $\lambda$ such that $P$ is the completion of a compact Liouville domain $(\bar{P},\left.\lambda\right|_{\bar{P}})$ satisfying $2c_{1}(T\bar{P})=0$.
-Let us consider
-$\begin{array}[]{ll}Y\coloneqq P\times\mathbb{R},&\alpha\coloneqq dz-\lambda,\end{array}$
-where $z$ is the coordinate of $\mathbb{R}$.
+Let us consider $Y\coloneqq P\times\mathbb{R}$ and $\alpha\coloneqq dz-\lambda$, where $z$ is the coordinate of $\mathbb{R}$.
 Then, $\alpha$ is a contact form on $Y$ and the Reeb vector field on $Y$ with respect to $\alpha$ agrees with $\partial_{z}$.
 Let $\pi_{P}\colon Y\to P$ denote the projection to $P$.
-The arguments from Section 2.1 to Section 2.4 are parallel to those in [16, Section 3], except the difference of the coefficient rings.
+The arguments from Section 2.1 to Section 2.4 are parallel to those in [15, Section 3], except the difference of the coefficient rings.
 In Section 2.5, we will discuss a particular case related to clean Lagrangian intersections.
@@
 Near each boundary puncture of the Riemann surface $(D_{m+1},j_{\kappa})$, we fix a strip coordinate by a biholomorphic map
-$\psi_{0}$ $\colon[0,\infty)\times[0,1]\to D_{m+1},$
-$\psi_{k}$ $\colon(-\infty,0]\times[0,1]\to D_{m+1}\text{ for }k=1,\dots,m,$
-with respect to $j_{0}$ and $j_{\kappa}$ such that $\lim_{s\to\infty}\psi_{0}(s,t)=p_{0}$ and $\lim_{s\to-\infty}\psi_{k}(s,t)=p_{k}$ uniformly on $t\in[0,1]$.
+$\begin{array}[]{cc}\psi_{0}\colon[0,\infty)\times[0,1]\to D_{m+1},&\psi_{k}\colon(-\infty,0]\times[0,1]\to D_{m+1}\end{array}$
+for $k\in\{1,\dots,m\}$ with respect to $j_{0}$ and $j_{\kappa}$ such that $\lim_{s\to\infty}\psi_{0}(s,t)=p_{0}$ and $\lim_{s\to-\infty}\psi_{k}(s,t)=p_{k}$ uniformly on $t\in[0,1]$.
 Let $\overline{D_{m+1}}$ be a compactification of $D_{m+1}$ obtained by gluing $0,\infty]\times[0,1]$ via $\psi_{0}$ and $\infty,0]\times[0,1]$ via $\psi_{k}$ for every $k=1,\dots,m$.
 In a natural way, $\overline{D_{m+1}}$ has a structure of a compact oriented surface with boundary and corners.
-Let $\Lambda$ be a compact Legendrian submanifold of $(Y,\alpha)$. We denote the set of Reeb chords of $\Lambda$ by
+Let $\Lambda$ be a compact Legendrian submanifold of $(Y,\alpha)$. Denote the set of its Reeb chords by
 ${\mathcal{R}(\Lambda)\coloneqq\{a\colon[0,T]\to Y\mid T>0,\ \frac{da}{dt}(t)=(\partial_{z})_{a(t)}\text{ for every }t\in[0,T]\text{ and }a(0),a(T)\in\Lambda\}.}$
 Hereafter, for any Reeb chord $(a\in[0,T]\to Y)\in\mathcal{R}(\Lambda)$, we write $T=\int a^{*}\alpha$ as $T_{a}$.
-For each $a\in\mathcal{R}(\Lambda)$, we fix a continuous map $w_{a}\colon\overline{D_{1}}\to Y$ such that $w_{a}(\partial D_{1})\subset\Lambda$ and
-$\lim_{s\to\infty}w_{a}\circ\psi_{0}(s,t)=a(T_{a}t)$
-for every $t\in[0,1]$.
+For each $a\in\mathcal{R}(\Lambda)$, we fix a continuous map $w_{a}\colon\overline{D_{1}}\to Y$ such that $w_{a}(\partial D_{1})\subset\Lambda$ and $\lim_{s\to\infty}w_{a}\circ\psi_{0}(s,t)=a(T_{a}t)$ for every $t\in[0,1]$.
 2.2 Moduli space of pseudo-holomorphic curves with boundary in an exact Lagrangian cobordism
@@
     and $L\cap([r_{-},r_{+}]\times Y)$ is compact.
-  • There exists a $C^{\infty}$ function $f\colon L\to\mathbb{R}$ such that $df=\left.(e^{r}\alpha)\right|_{L}$ and its restriction
-    $\left.f\right|_{(-\infty,r_{-}]\times\Lambda_{-}}\colon(-\infty,r_{-}]\times\Lambda_{-}\to\mathbb{R}$
-    is a constant function. Here, $r$ is the $\mathbb{R}$-coordinate of $\mathbb{R}\times Y$.
+  • There exists a $C^{\infty}$ function $f\colon L\to\mathbb{R}$ such that $df=\left.(e^{r}\alpha)\right|_{L}$ and $f$ is constant when restricted on $(-\infty,r_{-}]\times\Lambda_{-}$.
+Here, $r$ is the $\mathbb{R}$-coordinate of $\mathbb{R}\times Y$.
 In the rest of this paper, we only consider the case where $\Lambda_{-}$ is connected.
@@
 Associated to $L$, we use the following notations for maps between pairs of spaces:
-$\begin{split}i_{+}&\colon(Y,\Lambda_{+})\to(\mathbb{R}\times Y,L)\colon x\mapsto(r_{+},x),\\ i_{-}&\colon(Y,\Lambda_{-})\to(\mathbb{R}\times Y,L)\colon x\mapsto(r_{-},x),\end{split}$ (1)
+$\begin{array}[]{cc}i_{+}\colon(Y,\Lambda_{+})\to(\mathbb{R}\times Y,L)\colon x\mapsto(r_{+},x),&i_{-}\colon(Y,\Lambda_{-})\to(\mathbb{R}\times Y,L)\colon x\mapsto(r_{-},x),\end{array}$ (2)
 and
@@
 We associate to $J$ an almost complex structure $\widetilde{J}$ on $\mathbb{R}\times Y$ determined by $\tau_{s}^{*}\widetilde{J}=\widetilde{J}$ for every $s\in\mathbb{R}$, $\widetilde{J}(\operatorname{Ker}\alpha)=\operatorname{Ker}\alpha$, $(\pi_{P})_{*}(\widetilde{J}(v))=J((\pi_{P})_{*}(v))$ for every $v\in\operatorname{Ker}\alpha$ and $\widetilde{J}(\partial_{z})=\partial_{r}$.
-For any $a\in\mathcal{R}(\Lambda_{+})$, $b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda_{-})$ and $C\in H_{2}(\mathbb{R}\times Y,L)$, we define
-$\hat{\mathcal{M}}_{L,C}(a;b_{1},\dots,b_{m})$
-to be the space of pairs $(u,\kappa)$ of $\kappa\in\mathcal{C}_{m+1}$ and a $C^{\infty}$ map $u\colon D_{m+1}\to\mathbb{R}\times Y$ such that:
+For any $a\in\mathcal{R}(\Lambda_{+})$, $b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda_{-})$ and $C\in H_{2}(\mathbb{R}\times Y,L)$, we define $\hat{\mathcal{M}}_{L,C}(a;b_{1},\dots,b_{m})$ to be the space of pairs $(u,\kappa)$ of $\kappa\in\mathcal{C}_{m+1}$ and a $C^{\infty}$ map $u\colon D_{m+1}\to\mathbb{R}\times Y$ such that:
   • $du+\widetilde{J}\circ du\circ j_{\kappa}=0$.
@@
   • There exist $s_{0},s_{1},\dots,s_{m}\in\mathbb{R}$ such that
-    $\lim_{s\to\infty}\tau_{-s}\circ u\circ\psi_{0}$ $=(s_{0},a(T_{a}(1-t))),$
-    $\lim_{s\to-\infty}\tau_{s}\circ u\circ\psi_{k}$ $=(s_{k},b_{k}(T_{b_{k}}(1-t)))\text{ for }k=1,\dots,m,$
-    $C^{\infty}$-uniformly on $[0,1]$.
+    $\begin{array}[]{cc}\lim_{s\to\infty}\tau_{-s}\circ u\circ\psi_{0}=(s_{0},a(T_{a}(1-t))),&\lim_{s\to-\infty}\tau_{s}\circ u\circ\psi_{k}=(s_{k},b_{k}(T_{b_{k}}(1-t)))\end{array}$
+    for $k\in\{1,\dots,m\}$, $C^{\infty}$-uniformly on $[0,1]$.
   • Let $\bar{u}\colon\overline{D_{m+1}}\to L$ denote the continuous extension of the map $R_{L}\circ u\colon D_{m+1}\to L$. Then, the 2-chain
@@
 The second condition is satisfied for generic $\Lambda$.
-Under the second condition, the Conley-Zehnder index $\mu(a)\in\mathbb{Z}$ of $a$ is defined by using the capping path $\left.w_{a}\right|_{\partial D_{1}}$ for $a$ as in [16, Section 3.1].
-We then define
-$|a|\coloneqq\mu(a)-1.$
-We also assume that $\Lambda$ is admissible in the sense of the definition in [5, Section 2.3]. Let us fix a spin structure $\mathfrak{s}$ on $\Lambda$.
+Under the second condition, the Conley-Zehnder index $\mu(a)\in\mathbb{Z}$ of $a$ is defined by using the capping path $\left.w_{a}\right|_{\partial D_{1}}$ for $a$ as in [15, Section 3.1].
+We then define $|a|\coloneqq\mu(a)-1$.
+We also assume that $\Lambda$ is admissible in the sense of the definition in [5, Section 2.3].
+Let us fix a spin structure $\mathfrak{s}$ on $\Lambda$.
 We take an almost complex structure $J$ on $P$ as in [5, Proposition 2.3] such that the following hold for every $a,b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda)$ and $A\in H_{2}(Y,\Lambda)$ such that $(b_{1},\dots,b_{m})\neq(a)$: The moduli space $\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ is cut out transversely and it is a manifold of dimension $|a|-(|b_{1}|+\dots+|b_{m}|)-1$.
 Moreover, when $|b_{1}|+\dots+|b_{m}|=|a|-1$, $\coprod_{A\in H_{2}(Y,\Lambda)}\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ is a compact 0-dimensional manifold.
-(For the relation between $J$-holomorphic disks in $P$ defined as in [5, Section 2.3] and $\widetilde{J}-$holomorphic maps in $\mathbb{R}\times Y$ defined as in Section 2.2, see [2, Theorem 2.1] and [16, Appendix A.4].)
+(For the relation between $J$-holomorphic disks in $P$ defined as in [5, Section 2.3] and $\widetilde{J}-$holomorphic maps in $\mathbb{R}\times Y$ defined as in Section 2.2, see [2, Theorem 2.1] and [15, Appendix A.4].)
 In addition, $\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ has an orientation depending on $\mathfrak{s}$ and a choice of an orientation of the determinant line of the capping operator for each Reeb chord of $\Lambda$.
-See [16, Definition A.4] for details which follow the conventions in [4] related to orientations.
+See [15, Definition A.4] for details which follow the conventions in [4] related to orientations.
 Let $\mathcal{A}_{*}(\Lambda)$ be the graded $\mathbb{Z}[H_{2}(Y,\Lambda)]$-algebra freely generated by $\mathcal{R}(\Lambda)$ whose grading is given by $|\cdot|\colon\mathcal{R}(\Lambda)\to\mathbb{Z}$.
-A derivation over $\mathbb{Z}[H_{2}(Y,\Lambda)]$ of degree 1-1
-$\partial_{J,\mathfrak{s}}\colon\mathcal{A}_{*}(\Lambda)\to\mathcal{A}_{*-1}(\Lambda)$
-is defined for every $a\in\mathcal{R}(\Lambda)$ by
+A derivation over $\mathbb{Z}[H_{2}(Y,\Lambda)]$ $\partial_{J,\mathfrak{s}}\colon\mathcal{A}_{*}(\Lambda)\to\mathcal{A}_{*-1}(\Lambda)$ is defined for every $a\in\mathcal{R}(\Lambda)$ by
 $\partial_{J,\mathfrak{s}}(a)\coloneqq\sum_{|b_{1}|+\dots+|b_{m}|=|a|-1}\sum_{A\in H_{2}(Y,\Lambda)}(-1)^{(n-1)(|a|+1)}\left(\#_{\operatorname{sign}}\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})\right)e^{A}\cdot b_{1}\cdots b_{m}$
@@
 $\mathfrak{s}_{-}$) are induced.
-See [16, Remark 3.6] for their precise descriptions.
-Associated to $(\Lambda_{+},\mathfrak{s}_{+})$ and $(\Lambda_{-},\mathfrak{s}_{-})$, the Chekanov-Eliashberg DGAs
-$\begin{array}[]{cc}(\mathcal{A}_{*}(\Lambda_{+}),\partial_{J,\mathfrak{s}_{+}}),&(\mathcal{A}_{*}(\Lambda_{-}),\partial_{J,\mathfrak{s}_{-}})\end{array}$
-are defined as in Section 2.3. The maps $i_{+}$ and $i_{-}$ of (1) induce ring homomorphisms
+See [15, Remark 3.6] for their precise descriptions.
+Associated to $(\Lambda_{+},\mathfrak{s}_{+})$ and $(\Lambda_{-},\mathfrak{s}_{-})$, the Chekanov-Eliashberg DGAs $(\mathcal{A}_{*}(\Lambda_{+}),\partial_{J,\mathfrak{s}_{+}})$ and $(\mathcal{A}_{*}(\Lambda_{-}),\partial_{J,\mathfrak{s}_{-}})$ are defined as in Section 2.3.
+The maps $i_{+}$ and $i_{-}$ of (2) induce ring homomorphisms
 $\begin{array}[]{ll}(i_{+})_{*}\colon\mathbb{Z}[H_{2}(Y,\Lambda_{+})]\to\mathbb{Z}[H_{2}(\mathbb{R}\times Y,L)],&(i_{-})_{*}\colon\mathbb{Z}[H_{2}(Y,\Lambda_{-})]\to\mathbb{Z}[H_{2}(\mathbb{R}\times Y,L)].\end{array}$
@@
 In addition, it has an orientation depending on $\mathfrak{s}_{L}$ and a choice of an orientation of the determinant line of the capping operator for each Reeb chord of $\Lambda_{+}$ and $\Lambda_{-}$.
-See [16, Section A.3].
-We refer to the proof of [16, Theorem 3.7] which takes $\mathbb{Z}[H_{1}(\Lambda_{\pm})]$ as the coefficient rings.
-We modify the arguments in [16, Appendix A.6] to the present case where we take $\mathbb{Z}[H_{2}(Y,\Lambda_{\pm})]$ as the coefficient rings.
-(As referred in [16, Appendix A], see also [10, Theorem 2.5] about the DGA map over $\mathbb{Z}$, but note that some conventions in [10] related to orientations are different from those in [4, 16].) Then, we obtain the following theorem.
+See [15, Section A.3].
+We refer to the proof of [15, Theorem 3.7] which takes $\mathbb{Z}[H_{1}(\Lambda_{\pm})]$ as the coefficient rings.
+We modify the arguments in [15, Appendix A.6] to the present case where we take $\mathbb{Z}[H_{2}(Y,\Lambda_{\pm})]$ as the coefficient rings.
+(As referred in [15, Appendix A], see also [9, Theorem 2.5] about the DGA map over $\mathbb{Z}$, but note that some conventions in [9] related to orientations are different from those in [4, 15].) Then, we obtain the following theorem.
 Theorem 2.3.
@@
 The manifold $\mathbb{R}^{3}$ is equipped with a standard Riemannian metric $\langle\cdot,\cdot\rangle\coloneqq\sum_{i=1}^{3}dq_{i}\otimes dq_{i}$ and an orientation such that $dq_{1}\wedge dq_{2}\wedge dq_{3}$ is a positive volume form.
-A coordinate $((q_{1},q_{2},q_{3}),(p_{1},p_{2},p_{3}))$ of $T^{*}\mathbb{R}^{3}$ is defined via the diffeomorphism
-$\mathbb{R}^{3}\times\mathbb{R}^{3}\to T^{*}\mathbb{R}^{3}\colon((q_{1},q_{2},q_{3}),(p_{1},p_{2},p_{3}))\mapsto\sum_{i=1}^{3}p_{i}(dq_{i})_{(q_{1},q_{2},q_{3})}.$
+Let $(p_{1},\dots,p_{3})$ be the coordinate of the fiber.
 For any $r>0$, we denote $D^{*}_{r}\mathbb{R}^{3}\coloneqq\{(q,p)\in T^{*}\mathbb{R}^{3}\mid|p|\leq r\}$.
 The canonical Liouville form on $T^{*}\mathbb{R}^{3}$ is given by $\lambda_{\mathbb{R}^{3}}=\sum_{i=1}^{3}p_{i}dq_{i}$.
-We consider the unit cotangent bundle of $\mathbb{R}^{3}$
-$S^{*}\mathbb{R}^{3}\coloneqq\{(q,p)\in T^{*}\mathbb{R}^{3}\mid|p|=1\}.$
-It is endowed with a contact form $\alpha_{\mathbb{R}^{3}}\coloneqq\left.\lambda_{\mathbb{R}^{3}}\right|_{S^{*}\mathbb{R}^{3}}$. Note that there exists a diffeomorphism
+We consider the unit cotangent bundle $S^{*}\mathbb{R}^{3}\coloneqq\{(q,p)\in T^{*}\mathbb{R}^{3}\mid|p|=1\}$.
+It is endowed with a contact form $\alpha_{\mathbb{R}^{3}}\coloneqq\left.\lambda_{\mathbb{R}^{3}}\right|_{S^{*}\mathbb{R}^{3}}$.
+Note that there exists a diffeomorphism
 $S^{*}\mathbb{R}^{3}\to T^{*}S^{2}\times\mathbb{R}\colon(q,p)\mapsto((p,q-\langle q,p\rangle p),\langle q,p\rangle)$
@@
 Therefore, by taking $(T^{*}S^{2},\lambda_{S^{2}})$ as $(P,\lambda)$, we can define the Chekanov-Eliashberg DGA for Legendrian submanifolds of $S^{*}\mathbb{R}^{3}$ as in Section 2.3.
-We focus on a special kind of Legendrian submanifolds.
-For any knot $K$ in $\mathbb{R}^{3}$, recall its conormal bundle $L_{K}$ defined in Section 1.
-Then, the unit conormal bundle of $K$ is defined by
+We focus on a special kind of Legendrian submanifolds. For any knot $K$ in $\mathbb{R}^{3}$, let $L_{K}$ be its conormal bundle. Then, the unit conormal bundle of $K$ is defined by
 $\Lambda_{K}\coloneqq L_{K}\cap S^{*}\mathbb{R}^{3}=\{(q,p)\in S^{*}\mathbb{R}^{3}\mid q\in K,\ \langle p,v\rangle=0\text{ for every }v\in T_{q}K\}.$
@@
 We also fix a spin structure $\mathfrak{s}_{L_{K}}$ given by a trivial $\mathrm{Spin}(3)$-bundle on $L_{K}$.
-As the positive end of $L_{K}$, an orientation and a spin structure $\mathfrak{s}_{K}$ on $\Lambda_{K}$ are determined as in [16, Remark 3.6].
+As the positive end of $L_{K}$, an orientation and a spin structure $\mathfrak{s}_{K}$ on $\Lambda_{K}$ are determined as in [15, Remark 3.6].
 We fix an orientation of $K$.
@@
 Then, the homology classes $[l]$ and $[m]$ give a basis of $H_{1}(\Lambda_{K})$.
-We choose a non-vanishing vector field $V$ on $K$. Then, $\Lambda_{K}$ is bounded by a compact orientable manifold
-$\{(q,p)\in S^{*}\mathbb{R}^{3}\mid q\in K,\ p(V_{q})\geq 0\}.$
-Therefore, the homology class $\Lambda_{K}]\in H_{2}(S^{*}\mathbb{R}^{3})$ vanishes. For the pair $(S^{*}\mathbb{R}^{3},\Lambda_{K})$, we have a short exact sequence
+We choose a non-vanishing vector field $V$ on $K$.
+Then, $\Lambda_{K}$ is bounded by a compact orientable manifold $\{(q,p)\in S^{*}\mathbb{R}^{3}\mid q\in K,\ p(V_{q})\geq 0\}$.
+Therefore, the homology class $\Lambda_{K}]\in H_{2}(S^{*}\mathbb{R}^{3})$ vanishes.
+For the pair $(S^{*}\mathbb{R}^{3},\Lambda_{K})$, we have a short exact sequence
 ${H_{2}(\Lambda_{K})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{0}$${H_{2}(S^{*}\mathbb{R}^{3})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\partial}$${H_{1}(\Lambda_{K})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{1}(S^{*}\mathbb{R}^{3})=0.}$
@@
 For the proof of the main results in Section 3, it is not important how to specify the lifts $\widehat{l}$ and $\widehat{m}$ (see Theorem 3.2).
-We denote $\lambda\coloneqq e^{\widehat{l}}$, $\mu\coloneqq e^{\widehat{m}}$ and $U\coloneqq e^{\widehat{s}}$, then
-$\mathbb{Z}[H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K})]=\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}].$
+We denote $\lambda\coloneqq e^{\widehat{l}}$, $\mu\coloneqq e^{\widehat{m}}$ and $U\coloneqq e^{\widehat{s}}$, then $\mathbb{Z}[H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K})]=\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$.
 From the arguments in Section 2.3, the Chekanov-Eliashberg DGA $(\mathcal{A}_{*}(\Lambda_{K}),\partial_{J,\mathfrak{s}_{K}})$ for $(\Lambda_{K},\mathfrak{s}_{K})$ is defined with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$.
@@
 Definition 2.5.
-Let $L$ and $L^{\prime}$ be submanifolds of $T^{*}\mathbb{R}^{3}$. We say that $L$ and $L^{\prime}$ have a clean intersection along $L\cap L^{\prime}$ if the following hold:
-  • $L\cap L^{\prime}$ is a submanifold of both $L$ and $L^{\prime}$.
-  • For every $x\in L\cap L^{\prime}$, $T_{x}(L\cap L^{\prime})=T_{x}L\cap T_{x}L^{\prime}$.
-Let us make a setup to deal with Question 1.1.
+Let $L$ and $L^{\prime}$ be submanifolds of $T^{*}\mathbb{R}^{3}$.
+We say that $L$ and $L^{\prime}$ have a clean intersection along $L\cap L^{\prime}$ if $L\cap L^{\prime}$ is a submanifold of both $L$ and $L^{\prime}$, and $T_{x}(L\cap L^{\prime})=T_{x}L\cap T_{x}L^{\prime}$ for every $x\in L\cap L^{\prime}$.
+Let us make a setup.
 Let $K_{0}$ and $K_{1}$ be oriented knots in $\mathbb{R}^{3}$ and let $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$.
 Suppose that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along $K_{1}$.
-By moving $\varphi(L_{K_{0}})$ by a Hamiltonian isotopy supported in a neighborhood of $\varphi(L_{K_{0}})\cap\mathbb{R}^{3}=K_{1}$, we may assume that there exists $\varepsilon>0$ such that
-$\varphi(L_{K_{0}})\cap D_{\varepsilon}^{*}\mathbb{R}^{3}=L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}$
-For the proof, see [16, Lemma 2.3]. We consider a diffeomorphism
+By moving $\varphi(L_{K_{0}})$ by a Hamiltonian isotopy supported in a neighborhood of $\varphi(L_{K_{0}})\cap\mathbb{R}^{3}=K_{1}$, we may assume that there exists $\varepsilon>0$ such that $\varphi(L_{K_{0}})\cap D_{\varepsilon}^{*}\mathbb{R}^{3}=L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}$.
+For the proof, see [15, Lemma 2.3].
+We consider a diffeomorphism
 $F\colon\mathbb{R}\times S^{*}\mathbb{R}^{3}\to T^{*}\mathbb{R}^{3}\setminus\mathbb{R}^{3}\colon(r,(q,p))\mapsto(q,e^{r}p),$
-for which $F^{*}(\lambda_{\mathbb{R}^{3}})=e^{r}\alpha_{\mathbb{R}^{3}}$. Furthermore,
-$L_{\varphi}\coloneqq F^{-1}(\varphi(L_{K_{0}})\setminus K_{1})$
-is an exact Lagrangian cobordism from $\Lambda_{K_{1}}$ to $\Lambda_{K_{0}}$. We have the inclusion maps (1)
+for which $F^{*}(\lambda_{\mathbb{R}^{3}})=e^{r}\alpha_{\mathbb{R}^{3}}$.
+Furthermore, $L_{\varphi}\coloneqq F^{-1}(\varphi(L_{K_{0}})\setminus K_{1})$ is an exact Lagrangian cobordism from $\Lambda_{K_{1}}$ to $\Lambda_{K_{0}}$.
+We have the inclusion maps (2)
 $\begin{array}[]{cc}i_{+}\colon(S^{*}\mathbb{R}^{3},\Lambda_{K_{0}})\to(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi}),&i_{-}\colon(S^{*}\mathbb{R}^{3},\Lambda_{K_{1}})\to(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi}).\end{array}$
@@
 The proof is left to Appendix A.
-Note that the first assertion about orientations was shown in [16, Lemma 4.4] and a version of the second assertion with coefficients in $\mathbb{F}_{2}$ was shown by [16, Proposition 2.8, Lemma 4.5] using Floer cohomology group over $\mathbb{F}_{2}$ together with the triangle product.
-It should be possible to prove Proposition 2.7 in a way parallel to [16, Section 2.3] by lifting the coefficient ring from $\mathbb{F}_{2}$ to $\mathbb{Z}$.
+Note that the first assertion about orientations was shown in [15, Lemma 4.4] and a version of the second assertion with coefficients in $\mathbb{F}_{2}$ was shown by [15, Proposition 2.8, Lemma 4.5] using Floer cohomology group over $\mathbb{F}_{2}$ together with the triangle product.
+It should be possible to prove Proposition 2.7 in a way parallel to [15, Section 2.3] by lifting the coefficient ring from $\mathbb{F}_{2}$ to $\mathbb{Z}$.
 However, we will give a simpler proof in Appendix A by using Floer cohomology with coefficients in $\mathbb{F}_{2}[H_{1}(L_{K_{0}})]\cong\mathbb{F}_{2}[\lambda^{\pm}]$.
@@
 From Proposition 2.7, we take $a\in\{\pm 1\}$ such that $K_{1}]=a\cdot[\varphi(K_{0})]$ in $H_{1}(\varphi(L_{K_{0}}))$.
-Consider the map induced by $i_{\pm}$ on the first homology groups
-$\begin{array}[]{ll}(i_{+})_{*}\colon H_{1}(\Lambda_{K_{0}})\to H_{1}(L_{\varphi}),&(i_{-})_{*}\colon H_{1}(\Lambda_{K_{1}})\to H_{1}(L_{\varphi}).\end{array}$
-A basis of $H_{1}(L_{\varphi})$ is given by $\{l_{0},m_{0}\}$ for $l_{0}\coloneqq(i_{+})_{*}([l])$ and $m_{0}\coloneqq(i_{-})_{*}([m])$. It suffices to show that the equations
+Consider the map induced by $i_{\pm}$ on the first homology groups $(i_{+})_{*}\colon H_{1}(\Lambda_{K_{0}})\to H_{1}(L_{\varphi})$ and $(i_{-})_{*}\colon H_{1}(\Lambda_{K_{1}})\to H_{1}(L_{\varphi})$.
+A basis of $H_{1}(L_{\varphi})$ is given by $\{l_{0},m_{0}\}$ for $l_{0}\coloneqq(i_{+})_{*}([l])$ and $m_{0}\coloneqq(i_{-})_{*}([m])$.
+It suffices to show that the equations
 $\begin{array}[]{cc}(i_{-})_{*}([l])=a\cdot l_{0}+b\cdot m_{0},&(i_{+})_{*}([m])=a\cdot m_{0},\end{array}$ (6)
@@
 This shows that the first equation of (6) holds for some $b\in\mathbb{Z}$.
-Next, take an embedded disk $D\subset\varphi(L_{K_{0}})$ which bounds the circle $(F\circ i_{+}\circ m)(S^{1})$. $D$ is oriented so that
-[…]
-with respect to the diffeomorphism $\left.F\right|_{L_{\varphi}}\colon L_{\varphi}\to\varphi(L_{K_{0}})\setminus K_{1}$.
+Next, take an embedded disk $D\subset\varphi(L_{K_{0}})$ which bounds the circle $(F\circ i_{+}\circ m)(S^{1})$.
+$D$ is oriented so that $\partial D]=(\left.F\right|_{L_{\varphi}})_{*}((i_{+})_{*}([m]))$ in $H_{1}(\varphi(L_{K_{0}})\setminus K_{1})$ with respect to the diffeomorphism $\left.F\right|_{L_{\varphi}}\colon L_{\varphi}\to\varphi(L_{K_{0}})\setminus K_{1}$.
 We may assume that $D$ intersects $K_{1}$ transversely.
@@
 (Here, we need to care about the underlying orientations of spin manifolds.
-See the arguments in [16, Subsection 4.2.3].) Then, the existence of a DGA map $\Phi_{L_{\varphi}}$ follows from Theorem 2.3 and the description of the base change from $R_{-}$ to $R_{+}$ follows from Lemma 2.6.
+See the arguments in [15, Subsection 4.2.3].) Then, the existence of a DGA map $\Phi_{L_{\varphi}}$ follows from Theorem 2.3 and the description of the base change from $R_{-}$ to $R_{+}$ follows from Lemma 2.6.
 ∎
@@
 We will combine Proposition 2.8 with the results in [3, 13, 14] about the knot DGA.
-Some conventions about the knot DGA in these papers are slightly different from each other, as summarized in [15, Appendix A].
+Some conventions about the knot DGA in these papers are slightly different from each other, as summarized in [14, Appendix A].
 In this section, we mainly refer to [14] for the definition of the knot DGA.
 Let $K$ be an oriented knot in $\mathbb{R}^{3}$ given by the closure of a braid $B$.
-Then, a DGA over $R$, called the knot DGA of $K$, is constructed from $B$ as in [15, Definition 3.11].
+Then, a DGA over $R$, called the knot DGA of $K$, is constructed from $B$ as in [14, Definition 3.11].
 Its stable tame isomorphism class as a DGA over $R$ depends only on the isotopy class of $K$ as an oriented knot.
-See [14, Theorem 1.4] and [15, Theorem 3.1].
+See [13, Theorem 1.4] and [14, Theorem 3.1].
 Let us denote this DGA by $(\mathcal{A}^{\mathrm{knot}}_{*}(K),\partial_{K})$ and its homology $\operatorname{Ker}\partial_{K}/\operatorname{Im}\partial_{K}$ by $HC_{*}(K)$.
@@
 $\psi^{\prime}\colon R\to R\colon\lambda\mapsto-\lambda,\ \mu\mapsto-\mu,\ U\mapsto U.$
-The base change by $\psi^{\prime}$ is due to the difference of conventions between [3] and [15]. See [15, Appendix A]. It follows from this theorem that there are unital graded ring homomorphisms
+The base change by $\psi^{\prime}$ is due to the difference of conventions between [3] and [14]. See [14, Appendix A]. It follows from this theorem that there are unital graded ring homomorphisms
 $\begin{array}[]{cc}I_{K}\colon\mathcal{A}_{*}(\Lambda_{K})\to\mathcal{A}^{\mathrm{knot}}_{*}(K),&\bar{I}_{K}\colon\mathcal{A}^{\mathrm{knot}}_{*}(K)\to\mathcal{A}_{*}(\Lambda_{K}),\end{array}$ (9)
@@
 For $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$, suppose that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along $K_{1}$.
-Then we have the following theorem analogous to [16, Proposition 4.6].
+Then we have the following theorem analogous to [15, Proposition 4.6].
 Theorem 3.2.
@@
 then $\psi$ is described as in the theorem by using the integers $a,b,c$ and $d$ which determine (4). ∎
-We recall the definition of augmentations of a DGA. See, for instance, [15, Definition 2.6].
+We recall the definition of augmentations of a DGA. See, for instance, [14, Definition 2.6].
 Definition 3.3.
@@
 In the following, given an augmentation $\varepsilon$ as in the definition, we abbreviate $\varepsilon(r\cdot\mathbb{1})$ by $\varepsilon(r)$ for any $r\in R$.
-Let us prepare a lemma which is necessary to deduce Proposition 3.10 and Proposition 3.17.
-Lemma 3.4.
-Let $(\mathcal{A}_{*},\partial)$ be a unital non-commutative DGA over $R$ and let $\mathbb{k}$ be a field.
-Let $H^{\mathrm{ab}}_{0}(\mathcal{A}_{*},\partial)$ denote the commutative $R$-algebra obtained by the abelianization of the 0-th homology of $(\mathcal{A}_{*},\partial)$.
-Suppose that there exist a polynomial $P\in R[T]$ and a unital $R$-algebra map
-$J\colon R[T]/(P)\to H^{\mathrm{ab}}_{0}(\mathcal{A}_{*},\partial).$
-Then, for any augmentation $\varepsilon$ of $(\mathcal{A}_{*},\partial)$ to $\mathbb{k}$ such that $(\varepsilon(\lambda),\varepsilon(\mu),\varepsilon(U))=(x,y,Z)\in(\mathbb{k}^{*})^{3}$, the polynomial $\left.P\right|_{\lambda=x,\mu=y,U=Z}\in\mathbb{k}[T]$ has a root in $\mathbb{k}$.
+Let us also consider a knot invariant defined from $\mathbb{k}$-valued augmentations of the knot DGA.
+Definition 3.4 (Definition 5.1 of [14]).
+Fix a field $\mathbb{k}$. For an oriented knot $K$ in $\mathbb{R}^{3}$, its augmentation variety over $\mathbb{k}$ is a subset of $(\mathbb{k}^{*})^{3}$ defined by
+[…]
+From Theorem 3.2, we obtain a refinement of [15, Theorem 1.1].
+Corollary 3.5.
+There exist $a\in\{\pm 1\}$ and integers $b,c$ and $d$ for which we have an injective map
+$V_{\mathbb{k}}(K_{1})\hookrightarrow V_{\mathbb{k}}(K_{0})\colon(x,y,Z)\mapsto(x^{a}y^{b}Z^{c},y^{a}Z^{d},Z).$
 Proof.
-Any augmentation $\varepsilon$ of $(\mathcal{A}_{*},\partial)$ to $\mathbb{k}$ such that $(\varepsilon(\lambda),\varepsilon(\mu),\varepsilon(U))=(x,y,Z)$ induces a unital ring homomorphism $\bar{\varepsilon}\colon HC^{\mathrm{ab}}_{0}(\mathcal{A}_{*},\partial)\to\mathbb{k}$ such that $(\bar{\varepsilon}(\lambda),\bar{\varepsilon}(\mu),\bar{\varepsilon}(U))=(x,y,Z)$.
-We obtain a unital ring homomorphism $\bar{\varepsilon}\circ J\colon R[T]/(P)\to\mathbb{k}$, and $(\bar{\varepsilon}\circ J)(T)\in\mathbb{k}$ is a root of $\left.P\right|_{\lambda=x,\mu=y,U=Z}\in\mathbb{k}[T]$.
-∎
+Using $\Psi_{\varphi}$ of Theorem 3.2, it follows that for any augmentation $\varepsilon$ of $(\mathcal{A}_{*}^{\mathrm{knot}}(K_{1}),\partial_{K_{1}})$ to $\mathbb{k}$ such that $(\varepsilon(\lambda),\varepsilon(\mu),\varepsilon(U))=(x,y,Z)$, the composite map $\varepsilon\circ\Psi_{\varphi}\colon\mathcal{A}_{*}^{\mathrm{knot}}(K_{0})\to\mathbb{k}$ is an augmentation of $(\mathcal{A}_{*}^{\mathrm{knot}}(K_{0}),\partial_{K_{0}})$ to $\mathbb{k}$ such that
+$(\varepsilon\circ\Psi_{\varphi}(\lambda),\varepsilon\circ\Psi_{\varphi}(\mu),\varepsilon\circ\Psi_{\varphi}(U))=(x^{a}y^{b}Z^{c},y^{a}Z^{d},Z).$
+From the definition of $V_{\mathbb{k}}(K_{i})$ for $i\in\{0,1\}$, we obtain a map from $V_{\mathbb{k}}(K_{1})$ to $V_{\mathbb{k}}(K_{0})$ as in the assertion. ∎
 The next proposition is used to prove Theorem 3.14 and Theorem 3.17 by taking $\mathbb{k}=\mathbb{Q}$.
+Consider a projection $\pi_{y,Z}\colon(\mathbb{k}^{*})^{3}\to(\mathbb{k}^{*})^{2}\colon(x,y,Z)\mapsto(y,Z)$.
 Proposition 3.6.
-Suppose that $K_{1}$ is the unknot.
-Then, for any $y,Z\in\mathbb{k}^{*}$ satisfying $y\neq Z^{n}$ for every $n\in\mathbb{Z}$, there exists an augmentation $\varepsilon$ of $(\mathcal{A}^{\mathrm{knot}}_{*}(K_{0}),\partial_{K_{0}})$ to $\mathbb{k}$ such that $(\varepsilon(\mu),\varepsilon(U))=(y,Z)$.
+Suppose that $K_{1}$ is the unknot. Then,
+$\{(y,Z)\in(\mathbb{k}^{*})^{2}\mid y\neq Z^{n}\text{ for all }n\in\mathbb{Z}\}\subset\pi_{y,Z}(V_{\mathbb{k}}(K_{0})).$
 Proof.
-If $K_{1}$ is the unknot, from the trivial braid on a single strand, we can compute as in [15, Example 3.13] that $\mathcal{A}^{\mathrm{knot}}_{0}(K_{1})=R=\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$ and $\partial_{K_{1}}(\mathcal{A}_{1}^{\mathrm{knot}}(K_{1}))$ is the ideal generated by
-$U-\lambda-\mu+\lambda\mu.$
-Let us take $a\in\{\pm 1\}$ and $b,c,d\in\mathbb{Z}$ to be the integers in Theorem 3.2.
+If $K_{1}$ is the unknot, from the trivial braid on a single strand, we can compute as in [14, Example 3.13] that $\mathcal{A}^{\mathrm{knot}}_{0}(K_{1})=R=\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$ and $\partial_{K_{1}}(\mathcal{A}_{1}^{\mathrm{knot}}(K_{1}))$ is the ideal generated by $U-\lambda-\mu+\lambda\mu$.
+Let us take $a\in\{\pm 1\}$ and $b,c,d\in\mathbb{Z}$ to be the integers in Corollary 3.5.
 For any $y,Z\in\mathbb{k}^{*}$ satisfying $y\notin\{Z^{n}\mid n\in\mathbb{Z}\}$, we take an augmentation $\varepsilon_{1}\colon\mathcal{A}^{\mathrm{knot}}_{*}(K_{1})\to\mathbb{k}$ determined by
@@
 Note that $y\cdot Z^{-d}\neq 1$ and $y\cdot Z^{-d}\neq Z^{a}$ from the condition on $y$ and $Z$.
-Composing with the map $\Psi_{\varphi}$ of Theorem 3.2, we obtain $\varepsilon\coloneqq\varepsilon_{1}\circ\Psi_{\varphi}\colon\mathcal{A}^{\mathrm{knot}}_{*}(K_{0})\to\mathbb{k}$.
-It is an augmentation of the knot DGA of $K_{0}$ to $\mathbb{k}$ such that
-$\begin{array}[]{cc}\varepsilon(\mu)=\varepsilon_{1}(\mu^{a}U^{d})=y,&\varepsilon(U)=\varepsilon_{1}(U)=Z.\end{array}$
-Therefore, $\varepsilon$ satisfies the required condition. ∎
-Let us also consider a knot invariant defined from $\mathbb{C}$-valued augmentations of the knot DGA.
-Definition 3.6 (Definition 5.1 of [15]).
-For an oriented knot $K$ in $\mathbb{R}^{3}$, its augmentation variety is a subset of $(\mathbb{C}^{*})^{3}$ defined by
-[…]
-From Theorem 3.2, we obtain a refinement of [16, Theorem 1.1].
-Corollary 3.7.
-There exist $a\in\{\pm 1\}$ and integers $b,c$ and $d^{\prime}$ for which we have an injective map
-$V_{K_{1}}\hookrightarrow V_{K_{0}}\colon(x,y,Z)\mapsto(x^{a}y^{b}Z^{c},yZ^{d},Z).$
-Proof.
-Using $\Psi_{\varphi}$ of Theorem 3.2, it follows that for any augmentation $\varepsilon$ of $(\mathcal{A}_{*}^{\mathrm{knot}}(K_{1}),\partial_{K_{1}})$ to $\mathbb{C}$ such that $(\varepsilon(\lambda),\varepsilon(\mu),\varepsilon(U))=(x,y,Z)$, the composite map $\varepsilon\circ\Psi_{\varphi}\colon\mathcal{A}_{*}^{\mathrm{knot}}(K_{0})\to\mathbb{C}$ is an augmentation of $(\mathcal{A}_{*}^{\mathrm{knot}}(K_{0}),\partial_{K_{0}})$ to $\mathbb{C}$ such that
-$(\varepsilon\circ\Psi_{\varphi}(\lambda),\varepsilon\circ\Psi_{\varphi}(\mu),\varepsilon\circ\Psi_{\varphi}(U))=(x^{a}y^{b}Z^{c},y^{a}Z^{d},Z).$
-From the definition of $V_{K_{i}}$ for $i\in\{0,1\}$, we obtain a map from $V_{K_{1}}$ to $V_{K_{0}}$ as in the assertion. ∎
+By the map of Corollary 3.5, $(\varepsilon_{1}(\lambda),(y\cdot Z^{-d})^{a},Z)\in V_{\mathbb{k}}(K_{1})$ is sent to an element of $V_{\mathbb{k}}(K_{0})$ whose $(y,Z)$-component is $((y\cdot Z^{-d})\cdot Z^{d},Z)=(y,Z)$.
+∎
 3.2 Proof of Theorem 1.1 for $K^{\prime}\#T_{(2,2m+1)}$ ($m\neq 0,-1$)
@@
 First, let us review how to compute $HC_{0}(K)$ for an oriented knot $K$.
-We refer to [15, Section 3].
+We refer to [14, Section 3].
 Note that since $\mathcal{A}^{\mathrm{knot}}_{p}(K)=0$ for $p<0$, $HC_{0}(K)=\mathcal{A}^{\mathrm{knot}}_{0}(K)/\partial_{K}(\mathcal{A}^{\mathrm{knot}}_{1}(K))$.
@@
 Let $\mathcal{A}_{n}$ be the $R$-algebra freely generated by the set $\{a_{i,j}\mid i,j\in\{1,\dots,n\},\ i\neq j\}$.
-Then, a group homomorphism $B_{n}\to\mathrm{Aut}(\mathcal{A}_{n})\colon B\mapsto\phi_{B}$ is defined as in [15, Definition 3.3].
+Then, a group homomorphism $B_{n}\to\mathrm{Aut}(\mathcal{A}_{n})\colon B\mapsto\phi_{B}$ is defined as in [14, Definition 3.3].
 It is determined by $\phi_{\sigma_{k}}\colon\mathcal{A}_{n}\to\mathcal{A}_{n}$ for each generator $\sigma_{k}$, and $\phi_{\sigma_{k}}$ is given by
@@
 See also [6, Section 2.1]. We remark that $\phi_{B\cdot B^{\prime}}=\phi_{B}\circ\phi_{B^{\prime}}$ for any $B,B^{\prime}\in B_{n}$.
-The automorphism $\phi_{B}\in\mathrm{Aut}(\mathcal{A}_{n})$ for $B\in B_{n}$ is extended to an automorphism on $\mathcal{A}_{n+1}$ as follows: We consider the group homomorphism $B_{n+1}\to\mathrm{Aut}(\mathcal{A}_{n+1})$ defined as above and an injective map $B_{n}\to B_{n+1}$ defined by adding the extra $(n+1)$-th strand to any braid on $n$ strands (see [15, Remark 3.5]).
+The automorphism $\phi_{B}\in\mathrm{Aut}(\mathcal{A}_{n})$ for $B\in B_{n}$ is extended to an automorphism on $\mathcal{A}_{n+1}$ as follows: We consider the group homomorphism $B_{n+1}\to\mathrm{Aut}(\mathcal{A}_{n+1})$ defined as above and an injective map $B_{n}\to B_{n+1}$ defined by adding the extra $(n+1)$-th strand to any braid on $n$ strands (see [14, Remark 3.5]).
 Then, for any $B\in B_{n}$, $\phi_{B}\in\mathrm{Aut}(\mathcal{A}_{n+1})$ is defined via the composite map $B_{n}\hookrightarrow B_{n+1}\to\mathrm{Aut}(\mathcal{A}_{n+1})$.
@@
 Let $K$ be the oriented knot in $\mathbb{R}^{3}$ given by the closure of the braid $B$.
-By [15, Theorem 4.1, Remark 4.2], $HC_{0}(K)$ is the quotient of $\mathcal{A}_{n}$ by the two-sided ideal generated by all entries of the two matrices
+By [14, Theorem 4.1, Remark 4.2], $HC_{0}(K)$ is the quotient of $\mathcal{A}_{n}$ by the two-sided ideal generated by all entries of the two matrices
 $\begin{array}[]{cc}\mathbb{A}-\bm{\Lambda}\cdot\phi_{B}(\mathbb{A})\cdot\bm{\Lambda}^{-1},&\widehat{\mathbb{A}}-\bm{\Lambda}\cdot\bm{\Phi}^{L}_{B}\cdot\mathbb{A},\end{array}$ (13)
@@
 The isomorphism class of $HC_{0}(K)$ as a unital $R$-algebra depends only on the isotopy class of $K$ as an oriented knot.
-See [15, Theorem 3.15] and [14, Theorem 1.4 (2)].
+See [14, Theorem 3.15] and [13, Theorem 1.4 (2)].
 We consider a commutative $R$-algebra $HC_{0}^{\mathrm{ab}}(K)$ defined as the quotient of $HC_{0}(K)$ by the relations $\xi\eta=\eta\xi$ for all $\xi,\eta\in HC_{0}(K)$.
@@
 $B^{\#}_{m}\coloneqq\sigma_{n}\cdot(\sigma_{n+1})^{2m+1}\cdot B\in B_{n+2}$
-for every $m\in\mathbb{Z}_{\geq 0}$.
-Then, the closure of $B^{\#}_{m}$ gives the connected sum $K\#T_{(2,2m+1)}$ of $K$ and the $(2,2m+1)$-torus knot.
-See Figure 1 in the case of $m=1$.
-(When $m=0$, $T_{(2,1)}$ is the unknot.)
-[…]
-Figure. 1: The braid $B^{\#}_{m}\in B_{n+2}$ for $m=1$. The box with ‘$B$’ in the center represents the braid $B$ on $n$ strands.
+for every $m\in\mathbb{Z}_{\geq 0}$. Then, the closure of $B^{\#}_{m}$ gives the connected sum $K\#T_{(2,2m+1)}$ of $K$ and the $(2,2m+1)$-torus knot. (When $m=0$, $T_{(2,1)}$ is the unknot.)
 By using (11), let us compute $\phi_{B^{\#}_{m}}(a_{n+2,*})=\left(\phi_{\sigma_{n}}\circ\phi_{(\sigma_{n+1})^{2m+1}}\circ\phi_{B}\right)(a_{n+2,*})$.
@@
 Proposition 3.9.
-Let $\mathbb{k}$ be a field.
-For any augmentation $\varepsilon$ of $\mathcal{A}^{\mathrm{knot}}_{*}(K\#T_{(2,2m+1)})$ to $\mathbb{k}$ such that $(\varepsilon(\mu),\varepsilon(U))=(y,Z)\in(\mathbb{k}^{*})^{2}$, the polynomial $\left.P_{m}\right|_{\mu=y,U=Z}\in\mathbb{k}[T]$ has a root in $\mathbb{k}$.
+For any field $\mathbb{k}$,
+$\pi_{y,Z}(V_{\mathbb{k}}(K\#T_{(2,2m+1)}))\subset\{(y,Z)\in(\mathbb{k}^{*})^{2}\mid\left.P_{m}\right|_{\mu=y,U=Z}\text{ has a root in }\mathbb{k}\}.$
 Proof.
@@
 $J_{m}\colon R[T]/(P_{m})\to R[a_{i,j}\mid 1\leq i,j\leq n+2,\ i\neq j]/\mathcal{I}_{B^{\#}_{m}}=HC^{\mathrm{ab}}_{0}(K\#T_{(2,2m+1)})$
-which maps $T$ to $XY=a_{n,n+2}a_{n+2,n}$. Now the proposition follows from Lemma 3.4. ∎
+which maps $T$ to $XY=a_{n,n+2}a_{n+2,n}$.
+Given a $\mathbb{k}$-valued augmentation $\varepsilon$ of $\mathcal{A}^{\mathrm{knot}}_{*}(K\#T_{(2,2m+1)})$, it induces a unital ring homomorphism from $HC^{\mathrm{ab}}_{0}(K\#T_{(2,2m+1)})$ to $\mathbb{k}$.
+By precomposing $J_{m}$, we obtain a unital ring homomorphism $R[T]/(P_{m})\to\mathbb{k}$ which sends $\mu$ to $\varepsilon(\mu)$ and $U$ to $\varepsilon(U)$.
+This means that $\left.P_{m}\right|_{\mu=\varepsilon(\mu),U=\varepsilon(U)}\in\mathbb{k}[T]$ has a root in $\mathbb{k}$.
+Now, the proposition follows from the definition of $V_{\mathbb{k}}(K\#T_{(2,2m+1)})$.
+∎
 Example 3.10.
@@
 $=-\mu T^{2}+(3\mu-\mu^{2}-U)T+(U-\mu)(1-\mu).$
-From a braid representation $(\sigma_{1})^{3}\in B_{2}$ of $T_{(2,3)}$, it is not difficult to compute $HC_{0}(T_{(2,3)})$ as outlined in [15, Exercise 3.10].
+From a braid representation $(\sigma_{1})^{3}\in B_{2}$ of $T_{(2,3)}$, it is not difficult to compute $HC_{0}(T_{(2,3)})$ as outlined in [14, Exercise 3.10].
 $HC_{0}(T_{(2,3)})$ is isomorphic to a commutative $R$-algebra $R[X]/(G_{1},G_{2})$, where
 $\begin{array}[]{cc}G_{1}(X)\coloneqq UX^{2}-\mu UX+\lambda\mu^{3}(1-\mu),&G_{2}(X)\coloneqq UX^{2}+\lambda\mu^{2}X+\lambda\mu^{2}(\mu-U).\end{array}$
-See [15, Exercise 5.5]. We have identities
+See [14, Exercise 5.5]. We have identities
 $XG_{1}+\mu G_{2}$ $=UX^{3}+\lambda\mu^{3}(2-\mu)X+\lambda\mu^{3}(\mu-U),$
@@
 and
-$UX(XG_{1}+\mu G_{2})+\lambda\mu^{2}(\mu-U)G_{1}$
-$=\$$(UX^{2})^{2}+\lambda\mu^{2}(3\mu-\mu^{2}-U)(UX^{2})+\lambda^{2}\mu^{5}(1-\mu)(\mu-U)$
+$UX(XG_{1}+\mu G_{2})+\lambda\mu^{2}(\mu-U)G_{1}=\$$(UX^{2})^{2}+\lambda\mu^{2}(3\mu-\mu^{2}-U)(UX^{2})+\lambda^{2}\mu^{5}(1-\mu)(\mu-U)$
 […]
@@
 Consider the reduction of the coefficient ring from $R$ to $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$ by setting $U=1$.
+In this case, we obtain $V_{\mathbb{k}}\cap\{Z=1\}\subset\{(x,y)\mid\left.P_{m}\right|_{\mu=y,U=1}\text{ has a root in }\mathbb{k}\}$.
+However, this is a trivial.
 By the substitution $U=1$, we have $\left.P_{0}\right|_{U=1}=(1-\mu)^{2}+\mu T$, and if we define $h^{\prime}_{m},k^{\prime}_{m}\in\mathbb{Z}[\mu^{\pm}]$ by $h^{\prime}_{m}\coloneqq\left.h_{m}\right|_{T=-\mu^{-1}(1-\mu)^{2}}$ and $k^{\prime}_{m}\coloneqq\left.k_{m}\right|_{T=-\mu^{-1}(1-\mu)^{2}}$, then
@@
 $=\mu\left.P_{m}\right|_{U=1,T=-\mu^{-1}(1-\mu)^{2}}.$
-Here, the second equality follows from (20).
-By induction on $m=0,1,\dots$, we can see that the polynomial $\left.P_{m}\right|_{U=1}\in\mathbb{Z}[\mu^{\pm}][T]$ has a root $T=-\mu^{-1}(1-\mu)^{2}$.
-Hence, Proposition 3.10 is trivial for if an augmentation $\varepsilon$ satisfies $\varepsilon(U)=1$.
+Here, the second equality follows from (21). By induction on $m=0,1,\dots$, we can see that the polynomial $\left.P_{m}\right|_{U=1}\in\mathbb{Z}[\mu^{\pm}][T]$ has a root $T=-\mu^{-1}(1-\mu)^{2}$.
 3.2.3 Application of $\mathbb{Q}$-valued augmentations
@@
 ∎
-Combining Proposition 3.5 and Proposition 3.10 with $\mathbb{k}=\mathbb{Q}$, we obtain the following theorem.
+Now, we prove Theorem 1.1 for $K^{\prime}\#T_{(2,q)}$, where $q\neq\pm 1$ is an odd number.
 Theorem 3.14.
@@
 Proof.
-For a knot $K_{0}$, suppose that there exists $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection $K_{1}=\varphi(L_{K_{0}})\cap\mathbb{R}^{3}$ and $K_{1}$ is the unknot in $\mathbb{R}^{3}$.
-Assuming that $K_{0}=K\#T_{(2,2m+1)}$, we will deduce contradictions in two cases: the case $m\geq 1$ and the case $m\leq-2$.
-(The case $m\geq 1$) By Proposition 3.5, for every $Z_{0}\in\{r\in\mathbb{Q}^{*}\mid r^{n}\neq y_{0}\text{ for every }n\in\mathbb{Z}\}$, there exists an augmentation $\varepsilon$ of $(\mathcal{A}^{\mathrm{knot}}_{*}(K_{0}),\partial_{K_{0}})$ to $\mathbb{Q}$ such that $\varepsilon(\mu)=y_{0}$ and $\varepsilon(U)=Z_{0}$.
-Then, it follows from Proposition 3.10 that $\left.P_{m}\right|_{\mu=y_{0},U=Z_{0}}\in\mathbb{Q}[T]$ has a root in $\mathbb{Q}$.
-From the definition (23) of $P_{m}$, it is equivalent to that the equation for $T$
-$(Z_{0}-y_{0})\cdot\bar{h}_{m}(T)+y_{0}T\cdot\bar{k}_{m}(T)=0$ (24)
-has a solution in $\mathbb{Q}$.
-As a polynomial in $\mathbb{Q}(Z)[T]$, let us consider $(Z-y_{0})\cdot\bar{h}_{m}+y_{0}T\cdot\bar{k}_{m}$. Let $p_{1},\dots,p_{e}\in\mathbb{Q}(Z)[T]$ be irreducible elements such that
-$(Z-y_{0})\cdot\bar{h}_{m}+y_{0}T\cdot\bar{k}_{m}=p_{1}\cdots p_{e}.$
+Let us first consider the case $m\geq 1$. From Proposition 3.6 and Proposition 3.9 with $\mathbb{k}=\mathbb{Q}$, it suffices to show that in $(\mathbb{Q}^{*})^{2}$
+[…]
+For the fixed rational number $y_{0}\neq 1$, note that
+$\left.P_{m}\right|_{\mu=y_{0},U=Z}=(Z-y_{0})\cdot\bar{h}_{m}(T)+y_{0}T\cdot\bar{k}_{m}(T)$
+for any $Z\in(\mathbb{Q}^{*})^{2}$.
+Let us regard $\left.P_{m}\right|_{\mu=y_{0},U=Z}$ as a polynomial in $\mathbb{Q}(Z)[T]$.
+Take $p_{1},\dots,p_{e}\in\mathbb{Q}(Z)[T]$ to be irreducible elements such that $\left.P_{m}\right|_{\mu=y_{0},U=Z}=p_{1}\cdots p_{e}$.
 We claim that the degree of $p_{i}$ as a polynomial in $T$ is greater than or equal to $2$ for every $i\in\{1,\dots,e\}$.
@@
 By Hilbert’s irreduciblity theorem (see e.g.
-[11, Chapter 9, Corollary 2.4], which shows that $\mathbb{Q}$ is a Hilbertian field), there are infinitely many $Z_{0}\in\mathbb{Q}$ such that all of the specializations
+[10, Chapter 9, Corollary 2.4], which shows that $\mathbb{Q}$ is a Hilbertian field), there are infinitely many $Z_{0}\in\mathbb{Q}$ such that all of the specializations
 $\left.p_{1}\right|_{Z=Z_{0}},\dots,\left.p_{e}\right|_{Z=Z_{0}}\in\mathbb{Q}[T]$
@@
 In addition, for the fixed rational number $y_{0}\neq 1$, the set $\{r\in\mathbb{Q}^{*}\mid r^{n}=y_{0}\text{ for some }n\in\mathbb{Z}\}$ is also finite.
-As a consequence, there are infinitely many $Z_{0}\in\mathbb{Q}^{*}\setminus\{r\in\mathbb{Q}^{*}\mid r^{n}=y_{0}\text{ for some }n\in\mathbb{Z}\}$ such that $\left.p_{i}\right|_{Z=Z_{0}}\in\mathbb{Q}[T]$ is irreducible and $\deg(\left.p_{i}\right|_{Z=Z_{0}})\geq 2$ for every $i\in\{1,\dots,e\}$, which means that the equation (24) has no solution in $\mathbb{Q}$.
-This contradicts Proposition 3.5 and Proposition 3.10.
-(The case $m\leq-2$) Take a Hamiltonian $H\colon T^{*}\mathbb{R}^{3}\times[0,1]\to\mathbb{R}$ such that $\varphi=\varphi^{1}_{H}$. Consider an involution
+As a consequence, there are infinitely many $Z_{0}\in\mathbb{Q}^{*}\setminus\{r\in\mathbb{Q}^{*}\mid r^{n}=y_{0}\text{ for some }n\in\mathbb{Z}\}$ such that $\left.p_{i}\right|_{Z=Z_{0}}\in\mathbb{Q}[T]$ is irreducible and $\deg(\left.p_{i}\right|_{Z=Z_{0}})\geq 2$ for every $i\in\{1,\dots,e\}$, which means that $\left.P_{m}\right|_{\mu=y_{0},U=Z_{0}}$ has no root in $\mathbb{Q}$.
+It remains to show the case $m\geq-2$.
+Let $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ and suppose that $\varphi(L_{K_{0}})$ intersects $\mathbb{R}^{3}$ cleanly along the unknot $K_{1}$.
+Take a Hamiltonian $H\colon T^{*}\mathbb{R}^{3}\times[0,1]\to\mathbb{R}$ such that $\varphi=\varphi^{1}_{H}$.
+Consider an involution
 $\iota\colon T^{*}\mathbb{R}^{3}\to T^{*}\mathbb{R}^{3}\colon(q,p)=((q_{1},q_{2},q_{3}),(p_{1},p_{2},p_{3}))\mapsto((q_{1},q_{2},-q_{3}),(p_{1},p_{2},-p_{3}))$
@@
 $\varphi^{1}_{\overline{H}}(L_{\overline{K_{0}}})$ intersects $\mathbb{R}^{3}=\iota(\mathbb{R}^{3})$ cleanly along $\overline{K_{1}}$.
-Since $\overline{K_{1}}$ is the unknot and $\overline{K_{0}}$ is the connected sum of the mirror of $K$ and $T_{(2,-2m-1)}$, this contradicts to the result from the above case.
+Since $\overline{K_{1}}$ is the unknot and $\overline{K_{0}}$ is the connected sum of the mirror of $K$ and $T_{(2,-2m-1)}$, this contradicts to the result from the previous case.
 ∎
-Remark 3.16.
-To the author’s knowledge, it seems difficult to show Theorem 3.15, even when $K_{0}=T_{(2,3)}$, by only using the knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$.
-Indeed, for the trefoil knot $T_{(2,3)}$ and the unknot $O$,
-$\left.HC_{0}(T_{(2,3)})\right|_{U=1}$ $\cong\mathbb{Z}[\lambda^{\pm},\mu^{\pm}][X]/(X^{2}-\mu X+\lambda\mu^{3}(1-\mu),\ X^{2}+\lambda\mu^{2}X+\lambda\mu^{2}(\mu-1)),$
-$\left.HC_{0}(O)\right|_{U=1}$ $\cong\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]/((1-\lambda)(1-\mu)),$
-and we can find a unital $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra map
-$\left.HC_{0}(T_{(2,3)})\right|_{U=1}\to\left.HC_{0}(O)\right|_{U=1}\colon X\mapsto\mu-\mu^{2}.$
-Therefore, we cannot deduce a contradiction from the existence of a unital $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra map between the 0-th homology which is induced from the DGA map as in Theorem 3.2.
 3.3 Proof of Theorem 1.1 for $K^{\prime}\#4_{1}$
@@
 $\begin{split}P(T)\coloneqq&-(1-\mu)(U-\mu)+\left(-2+2\mu+\mu^{2}+(\mu^{-1}-1-\mu)U-\mu^{3}U^{-1}\right)T\\ &+(-2\mu^{2}+\mu U)T^{2}-\mu UT^{3}.\end{split}$ (25)
-Let $\mathbb{k}$ be a field.
-Then, for any augmentation $\varepsilon$ of $\mathcal{A}^{\mathrm{knot}}_{*}(K\#4_{1})$ to $\mathbb{k}$ such that $(\varepsilon(\mu),\varepsilon(U))=(y,Z)$, the polynomial $\left.P\right|_{\mu=y,U=Z}\in\mathbb{k}[T]$ has a root in $\mathbb{k}$.
+Then, for any field $\mathbb{k}$,
+$\pi_{y,Z}(V_{\mathbb{k}}(K\#4_{1}))\subset\{(y,Z)\in(\mathbb{k}^{*})^{2}\mid\left.P\right|_{\mu=y,U=Z}\text{ has a root in }\mathbb{k}\}.$
 Proof.
@@
 $+(-2\mu^{2}+\mu U)(XY)^{2}-\mu U(XY)^{3}.$
-If we define $P\in\mathbb{Z}[\mu^{\pm},U][T]$ by (25), then
-$(U-\mu)\cdot F+\mu Y\cdot G=P(XY),$
-which implies that $P(XY)$ is contained in $\mathcal{I}_{B^{\#}}$. Therefore, we have a well-defined unital $R$-algebra map
+If we define $P\in\mathbb{Z}[\mu^{\pm},U][T]$ by (25), then $(U-\mu)\cdot F+\mu Y\cdot G=P(XY)$, which implies that $P(XY)$ is contained in $\mathcal{I}_{B^{\#}}$.
+Therefore, we have a well-defined unital $R$-algebra map
 $J\colon R[T]/(P)\to R[a_{i,j}\mid 1\leq i,j\leq n+3,\ i\neq j]/\mathcal{I}_{B^{\#}}=HC^{\mathrm{ab}}_{0}(K\#4_{1})$
-which maps $T$ to $XY=a_{n,n+3}a_{n+3,n}$. Now, the proposition follows from Lemma 3.4. ∎
+which maps $T$ to $XY=a_{n,n+3}a_{n+3,n}$.
+The rest of argument is parallel to Proposition 3.9.
+Given a $\mathbb{k}$-valued augmentation $\varepsilon$ of $\mathcal{A}^{\mathrm{knot}}_{*}(K\#4_{1})$, it induces a unital ring homomorphism from $HC^{\mathrm{ab}}_{0}(K\#4_{1})$ to $\mathbb{k}$.
+By precomposing $J$, we obtain a unital ring homomorphism $R[T]/(P)\to\mathbb{k}$ which sends $\mu$ to $\varepsilon(\mu)$ and $U$ to $\varepsilon(U)$.
+This means that $\left.P\right|_{\mu=\varepsilon(\mu),U=\varepsilon(U)}\in\mathbb{k}[T]$ has a root in $\mathbb{k}$.
+Now, the proposition follows from the definition of $V_{\mathbb{k}}(K\#4_{1})$.
+∎
 Remark 3.16.
@@
 Proof.
-For a knot $K_{0}$, suppose that there exists $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection $K_{1}=\varphi(L_{K_{0}})\cap\mathbb{R}^{3}$ and $K_{1}$ is the unknot in $\mathbb{R}^{3}$.
-Assuming that $K_{0}=K\#4_{1}$, we deduce a contradiction as follows: By Proposition 3.5, there exists an augmentation $\varepsilon$ of $(\mathcal{A}^{\mathrm{knot}}_{*}(K_{0}),\partial_{K_{0}})$ to $\mathbb{Q}$ such that $\varepsilon(\mu)=-1$ and $\varepsilon(U)=-2$.
-Then, it follows from Proposition 3.17 that the polynomial $\bar{P}\coloneqq\left.P\right|_{\mu=-1,U=-2}\in\mathbb{Q}[T]$ has a root in $\mathbb{Q}$.
-On the other hand, from the definition of $P$,
-$\bar{P}(T)=2-\frac{3}{2}T-2T^{3}$
-Multiplying by $2$, we obtain a polynomial with integer coefficients
-$4-3T-4T^{3}.$
+Let us abbreviate $\bar{P}=\left.P\right|_{\mu=-1,U=-2}\in\mathbb{Q}[T]$.
+Note that $(-1,-2)\in\{(y,Z)\mid y\neq Z^{n}\text{ for all }n\in\mathbb{Z}\}$.
+If we can check that $\bar{P}$ has no root in $\mathbb{Q}$, the theorem immediately follows from Proposition 3.6 and Proposition 3.15.
+From the definition of $P$, $\bar{P}=2-\frac{3}{2}T-2T^{3}$.
+Multiplying by $2$, we obtain a polynomial with integer coefficients $4-3T-4T^{3}$.
 By the rational root theorem, the only possible rational roots of this polynomial are of the form $\frac{p}{q}$, where $p\in\mathbb{Z}$ is a divisor of $4$ and $q\in\mathbb{Z}$ is a divisor of 4-4.
-The set of such rational numbers consists of the integers
-$\pm 1,\pm 2,\pm 4,$
-and those which are not integers
-$\pm\frac{1}{2},\pm\frac{1}{4}.$
+The set of such rational numbers is $\{\pm 1,\pm 2,\pm 4\}\cup\{\pm\frac{1}{2},\pm\frac{1}{4}\}$.
 We note that $2\cdot\bar{P}(T)=4-T(4T^{2}+3)$.
-For every $n\in\{\pm 1,\pm 2,\pm 4\}$, the integer $4n^{2}+3$ is not a divisor of $4$, which implies that $2\cdot\bar{P}(n)\neq 0$.
+For any $n\in\{\pm 1,\pm 2,\pm 4\}$, the integer $4n^{2}+3$ is not a divisor of $4$, which implies that $2\cdot\bar{P}(n)\neq 0$.
 One can also check that
 $\begin{array}[]{llll}\bar{P}(\frac{1}{2})=1,&\bar{P}(-\frac{1}{2})=3,&\bar{P}(\frac{1}{4})=\frac{51}{32},&\bar{P}(-\frac{1}{4})=\frac{77}{32}.\end{array}$
-Therefore, we can conclude that $\bar{P}$ does not have a root in $\mathbb{Q}$, which contradicts Proposition 3.5 and Proposition 3.17. ∎
+Therefore, we can conclude that $\bar{P}$ does not have a root in $\mathbb{Q}$. ∎
 A Proof of Proposition 2.7
@@
 The Lagrangian submanifolds $\mathbb{R}^{3}$ and $L_{K_{0}}$ of $T^{*}\mathbb{R}^{3}$ are oriented as in Subsection 2.5.1.
-For any $\psi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$, we introduce a $\mathbb{Z}/2$-graded Floer cochain complex
+For any $\psi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$, we will introduce a $\mathbb{Z}/2$-graded Floer cochain complex
 $(CF^{*}(\psi(\mathbb{R}^{3}),L_{K_{0}};H),d_{J})$
-with coefficients in the Laurent polynomial ring $\mathbb{F}_{2}[\lambda^{\pm}]$. It is a slight enhancement of the Floer cochain complex over $\mathbb{F}_{2}$ constructed as in [19, Section 8].
+with coefficients in the Laurent polynomial ring $\mathbb{F}_{2}[\lambda^{\pm}]$. It is a slight enhancement of the Floer cochain complex over $\mathbb{F}_{2}$ constructed as in [18, Section 8].
 We choose a Hamiltonian $H\colon T^{*}\mathbb{R}^{3}\times[0,1]\to\mathbb{R}$ with compact support such that the Lagrangian submanifold $P_{H}\coloneqq(\varphi^{1}_{H}\circ\psi)(\mathbb{R}^{3})$ intersects $L_{K_{0}}$ transversely.
@@
 From the orientations of $P_{H}$ and $L_{K_{0}}$, a $\mathbb{Z}/2$-grading $|x|\in\mathbb{Z}/2$ is defined for every $x\in P_{H}\cap L_{K_{0}}$.
-See [18, Example 2.7, Section 2.d].
+See [17, Example 2.7, Section 2.d].
 Note that the grading is shifted by $1$ if we use the opposite orientation of $P_{H}$.
 We also choose an almost complex structure $J$ such that $d\lambda_{\mathbb{R}^{3}}(\cdot,J\cdot)$ is a Riemannian metric on $T^{*}\mathbb{R}^{3}$ and $J(\partial_{p_{i}})=\partial_{q_{i}}$ for $i\in\{1,2,3\}$ outside a compact subset of $T^{*}\mathbb{R}^{3}$.
-[…]
-Figure.
-2: The shaded region is the image of $u$.
-$c_{x}$ and $c_{y}$ are paths in $L_{K_{0}}$ and $u(\mathbb{R}\times\{1\})$ is contained in $L_{K_{0}}$.
-$u(\mathbb{R}\times\{0\})$ is contained in $\psi(\mathbb{R}^{3})$.
 Fix a base point $z_{0}\in L_{K_{0}}$.
@@
   • $A\in H_{1}(L_{K_{0}})$ is represented by a loop based at $z_{0}$ obtained by concatenating the three paths in $L_{K_{0}}$: $c_{x}$, the path $\infty,\infty]\to L_{K_{0}}\colon s\mapsto u(s,1)$ from $x$ to $y$, and $c_{y}^{-1}$.
-Figure 2 depicts $u(\mathbb{R}\times[0,1])\subset T^{*}\mathbb{R}^{3}$ and the three paths in $L_{K_{0}}$.
+Figure 1 below illustrates $u(\mathbb{R}\times[0,1])\subset T^{*}\mathbb{R}^{3}$ and the three paths in $L_{K_{0}}$.
 The group $\mathbb{R}$ acts on $\mathcal{M}_{(H,J),A}(x,y)$ by $(u\cdot\sigma)(s,t)=u(s+\sigma,t)$ for every $\sigma\in\mathbb{R}$.
-To define the $\mathbb{Z}/2$-graded Floer cochain complex, we refer to [18, Section 2.f].
+[…]
+Figure.
+1: The shaded region is the image of $u$.
+$c_{x}$ and $c_{y}$ are paths in $L_{K_{0}}$ and $u(\mathbb{R}\times\{1\})$ is contained in $L_{K_{0}}$.
+$u(\mathbb{R}\times\{0\})$ is contained in $\psi(\mathbb{R}^{3})$.
+To define the $\mathbb{Z}/2$-graded Floer cochain complex, we refer to [17, Section 2.f].
 We take the pair $(H,J)$ to be regular so that the moduli space $\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ is cut out transversely for every $x,y\in P_{H}\cap L_{K_{0}}$ such that $x\neq y$.
@@
 Then, the claim of the proposition is equivalent to that $a\in\{1,-1\}$ and $\sigma\equiv 0\mod 2$.
-Since $\sigma\equiv 0\mod 2$ is already shown in [16, Lemma 4.4], we only give a proof for $a\in\{1,-1\}$ (see also Remark A.2 below).
+Since $\sigma\equiv 0\mod 2$ is already shown in [15, Lemma 4.4], we only give a proof for $a\in\{1,-1\}$ (see also Remark A.2 below).
 We choose a Mose function $f\colon\psi(K_{1})\to\mathbb{R}$ which has only two critical points $x_{0},x_{1}\in\psi(K_{1})$ such that $f$ takes its minimum at $x_{0}$ and its maximum at $x_{1}$.
@@
   • $\lim_{s\to-\infty}\gamma(s)=x_{0}$ and $\lim_{s\to\infty}\gamma(s)=x_{1}$.
-Since $\psi(\mathbb{R}^{3})$ and $L_{K_{0}}$ intersects cleanly along $\psi(K_{1})$ and $\psi(K_{1})$ is connected, we can apply [17, Theorem 3.4.11, Lemma 3.4.12] to get the following results: From the pair $(f,g)$ with $f||_{C^{1}}$ sufficiently small, we can construct a Hamiltonian $H_{f}\colon T^{*}\mathbb{R}^{3}\to\mathbb{R}$ supported in a neighborhood of $\psi(K_{1})$ and an almost complex structure $J_{g}$ on $T^{*}\mathbb{R}^{3}$ such that the pair $(H_{f},J_{g})$ is regular and the following hold:
-  • $P_{H_{f}}\cap L_{K_{0}}=\{x_{0},x_{1}\}$.
-  • There is a bijection
+Since $\psi(\mathbb{R}^{3})$ and $L_{K_{0}}$ intersects cleanly along $\psi(K_{1})$ and $\psi(K_{1})$ is connected, we can apply [16, Theorem 3.4.11, Lemma 3.4.12] to get the following results: From the pair $(f,g)$ with $f||_{C^{1}}$ sufficiently small, we can construct a Hamiltonian $H_{f}\colon T^{*}\mathbb{R}^{3}\to\mathbb{R}$ supported in a neighborhood of $\psi(K_{1})$ and an almost complex structure $J_{g}$ on $T^{*}\mathbb{R}^{3}$ such that the pair $(H_{f},J_{g})$ is regular and the following hold: $P_{H_{f}}\cap L_{K_{0}}=\{x_{0},x_{1}\}$ and there is a bijection
 $\Phi\colon\coprod_{A\in H_{1}(L_{K_{0}})}\mathcal{M}_{(H_{f},J_{g}),A}(x_{0},x_{1})\to\mathcal{T}_{(f,g)}(x_{0},x_{1})\colon u\mapsto u(\cdot,1).$
@@
 since $\psi(K_{1})]\in H_{1}(L_{K_{0}})$ is represented by a loop obtained by concatenating the two paths: $\infty,\infty]\to L_{K_{0}}\colon s\mapsto\gamma_{1}(-s)$ and $\infty,\infty]\to L_{K_{0}}\colon s\mapsto\gamma_{2}(s)$.
-From this observation, the differential $d_{J_{g}}\colon\mathbb{F}_{2}[\lambda^{\pm}]\cdot x_{1}\to\mathbb{F}_{2}[\lambda^{\pm}]\cdot x_{0}$ is given by
-$d_{J_{g}}(x_{1})=(\lambda^{k}+\lambda^{k+a})x_{0}.$
+From this observation, the differential $d_{J_{g}}\colon\mathbb{F}_{2}[\lambda^{\pm}]\cdot x_{1}\to\mathbb{F}_{2}[\lambda^{\pm}]\cdot x_{0}$ is given by $d_{J_{g}}(x_{1})=(\lambda^{k}+\lambda^{k+a})x_{0}$.
 Therefore, $HF^{*}(\psi(\mathbb{R}^{3}),L_{K_{0}};(H_{f},J_{g}))$ is computed as follow (we omit writing $(H_{f},J_{g})$):
@@
 Remark A.2.
-The fact $\sigma\equiv 0\mod 2$ was proved in [16, Lemma 4.4] by introducing $\mathbb{Z}$-gradings of Lagrangian submanifolds which are lifts of orientations of Lagrangian submanifolds.
-Analogous to [16, Remark 2.9, Lemma 4.4], by an argument on $\mathbb{Z}/2$-grading in the present case, we can deduce $\sigma\equiv 0\mod 2$ from $\sigma_{\psi}\equiv\sigma_{\operatorname{id}}\mod 2$.
+The fact $\sigma\equiv 0\mod 2$ was proved in [15, Lemma 4.4] by introducing $\mathbb{Z}$-gradings of Lagrangian submanifolds which are lifts of orientations of Lagrangian submanifolds.
+Analogous to [15, Remark 2.9, Lemma 4.4], by an argument on $\mathbb{Z}/2$-grading in the present case, we can deduce $\sigma\equiv 0\mod 2$ from $\sigma_{\psi}\equiv\sigma_{\operatorname{id}}\mod 2$.